A Complete Bibliography of Publications in Applied Mathematics and Computation: 1995–1999

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

20 December 2023 Version 1.05

Title word cross-reference

(k, n − k − 2) [AW99b]. 1 [Ada99, Ram97a, RGL97a]. 2 [LS98a, Van95]. 3 [Cha95a, SK96]. ^{(1)} [Chu97]. ^{ti} [DF96]. _{0} [Sol99a]. _{1} [KL99a]. _{2} [Chu97]. A [Che96a]. A_{T,S}^{(2)} [Che96a]. C^{0} [Lin98]. C^{∞} [Sas95a]. L_{2} [Yür99]. D [Zha99b]. E(3) [Beh95]. H [DK96, CW97]. L [GS95]. L_{p} [DEV95, JMS98]. LC^{1} [Sun98]. M [ABT98]. M/M/1 [Jac95]. N [JC97b, AP95, Kan97, Nar99a, YK97a]. ∇^{2}V (x_{j}) = 0 [Yan99]. p [Kwa96]. P (Z < Y) [Ami99]. P 1 [YK97b]. Q [GBS99, JC97b]. S [MBS98]. T [AdO98]. θ [Ram98a, Ram98b]. V [XL98, XL99]. Y_{n}(z) [ZB97]. z [Tag98]. ζ(2) [CRS99]. ζ(2n + 1) [Sri98].

-adaptive [CW97]. -approximation [DEV95]. -ary [GBS99]. -class [Sas95a]. -cycle [Kwa96, XL98, XL99]. -D [RGL97a, Cha95a, Ram97a, SK96]. -dimensional [Kan97, Nar99a]. -equation [DK96]. -matrices [ABT98]. -methods [Ram98a, Ram98b]. -nonconforming [KL99a]. -norm [JMS98]. -orientation [Zha99b]. -periodic [AdO98]. -simplexes [JC97b]. -stability [Sol99a]. -symmetric [ST96]. -system [MBS98]. -systems [GS95]. -transforms [Yür99]. -triangulated [JC97b].

1 [Arg99d]. 1994 [Cab95].

2-tori [Tak98]. 2nd [VD97].

63 [Cab95].

88 [Mus98].

abandonments [Mil95]. Ablowitz [TPGV97]. abrupt [Abu97]. ABS [ABT98]. absolute [SBH96, Waz97c]. accelerating [Boy95, SL99]. Acceleration [Zha96b, Zha99a, Zha98b]. account [HZ95, LC97]. accumulation [WZ98]. Accuracy [HY99b, HM99, Zen99]. Accurate [MRJ95, BS98, Hos99, HY99a]. action [MS97]. active [IOAB95]. Adaptive [ADS96, HAST98, MBCvdV95, RRB95, Sug95, XS95, CW97, KKH95, SU95, TSK98, WZZ98, ZS98]. Addendum [JS98a]. additive [XL98, YK97b]. admissible [AH99b]. Admission [Mil95]. Adomian [Waz98a, Waz99d]. adsorbers [RBPK96]. advanced [JS98b]. Advances [FpDS99]. advection [Waz95b]. advective [Hos99, HY99a]. advective-dispersive [HY99a]. aerial [AKK99]. aesthetic [Mur96]. affects [FHK97]. affinity [Bay98a]. AGE [KR95a, IKP97, KP98]. age-structured [IKP97, KP98]. agents [JL97]. aggregation [AB95b]. aging [PLPW96]. agriculture [SAF^{+}97]. albedo [KUO95]. algebra [Ada98, Bay98b, Bay98a, BY99, GSI96, TG97]. algebraic [Eis98, JN95, Lam99, LP97, MBS98]. algorithm [Abu99b, Die96, EMEH96, GS99, GMGCL99, HT98, JL97, Jin95, KKL99, Mor95, PS96, RG98, SS98a, SSY99, WZZ98, XZW96, XZL98, ZZ98, xZ98, ZL98c, dSYS97, ONOT98]. algorithms [CLC97, JBV98, JCC98, Lia97, MZV96, MFGW97, Per96, SR97, XZL97]. allied [CRS99]. allocation [BDS96]. almost [Kin97, OM99]. alternate [UK95]. alternating [SL96, Szi97, KR95a]. alternative [Ham98a, MSZ96]. alternatives [MS99, YB99]. Ambiguities [Kin97]. American [Váz98]. among [ITN97]. amphiphilic [Kub97]. amplitude [KSZ^{+}95]. Analysis [DML98, FH95, GL99, JC97a, RAL97, Sci97, Tuc96, WK95, ZA96, ABT98, AAES99, ALZ99, AFM97, BDS96, Gra97, JFW96, Kwa96, LS95, MZ98, MD98, Moh99b, Mor98, PZZy99, Ram99a, TH96, TJ95, Van97, VR99f, VR99a, VR99e, Won98, YL98a, Zha98b]. Analytic [ZL98a, ARM97, CSY98, DS99]. Analytical [Waz99a, SA96, SK95, TJ95]. Analyzing [Jac95, Bay97, PN97b]. anisotropy [SR97]. AOR [Bai99a, YJ99]. Appearance [TyG99, Waz97a]. Appl [Mus98]. Application [Bay98b, Bay98a, CH96, KA96a, Lia98, SKG97, AW99a, AK96, KKH95, Khu98b, Lin96, MNK99, MZV96, RRB95, SZ97, UR97, Waz95b, WW99a, WW99b, Yil96, ZL98b, AM98, HS99]. applications [ABT98, Arg96a, Arg99a, Arg99b, Arg99d, CS97, CSY98, Chu97, LF99, MA95, Par95, Yür99, Zha98a, dRS98]. Applied [Cab95, Bel95, DK96, DGU96, Udw95]. approach [Ano97i, Arg96a, Arg99d, BS98, Cha95a, CBYK99, EGH98, FHG96b, Ham97, Ham98a, JS98b, Khu98b, Lei95, Mus97a, Mus98, SEG96, SST95, SK98a, Tag99, Uen95, Váz98, Waz95b]. approaches [MSZ96, SK98b]. approximants [Waz99a, Waz99c]. Approximate [YST97, AD98, JPF95, Waz95a, XZL98, xZ98]. Approximating [MZ98, SUJ98]. approximation [BPZ96, Beh96b, CyGK99, DEV95, GL99, IKP97, JMS98, MS97, MvL98, MS95b, Pap96, Pav95, SK95, VD97, WR96, ZW95a]. approximations [AM98, Hal98, HH99, Let97, Li98, NR99b, Waz99a, ZB97]. aquaculture [PN97b]. arbitrary [ESR99]. architectures [KKPL96]. arcs [Mit99]. area [KUO95]. argument [AY98, Ino99, ZB97]. arguments [LG99, LF99, Pei99, WA98]. arising [ALZ99, DH99a, KW96, KW97, Khu98c, NR99a, RR98, Sas95b, Waz97b]. Arithmetic [GR97, Jay96]. arrival [Mil95, MS95b]. arrivals [Mil95]. Artificial [MZ98, PZZy99, JS98b, YL98b, ZSL98, ZPY99, Sug98]. ary [GBS99]. Aspects [KH96]. assay [MMMK96]. assessment [AS99]. assignment [Mou96]. associated [CSZ98, DS99, GS97, MS95a, MR99, PN97a, WLTS99]. associative [GKT96]. Asymptotic [FW98, GL95, Kar98, Koł95, Ram99a, WA98, Kim96, Kur99, NR99b, Sas95a, VR99f, VR99e, VR99f, VR99a, VR99e]. asymptotically [WA96]. Asynchronous [Bai98, Bai99a]. atmospheric [ÖGDB98]. atoms [RR98]. attainability [HK96]. attended [Mou96]. attenuation [LR95]. attracting [Sol99b]. attraction [Kur99]. attractivity [ZAES96]. attractors [DVD95]. attribute [MS99]. Author [Ano98d, Ano98e, Ano98f, Ano98a, Ano98b, Ano98c, Ano99c, Ano99a, Ano99d, Ano99b]. autocorrelation [BR99b]. automata [Ada99]. automaton [Ada97]. autonomous [HAST98]. autoregressive [AAAN98, LSKH96]. average [MF99]. averaged [Boy95]. axisymmetric [Khu98a, ZL98a].

backtracking [Ye99]. balance [YH97]. balls [WW99a]. Banach [Arg96a, Arg96b, Arg99d, Arg99e, Guo99, Ibr96]. Bargaining [Szi98, SL96, SS98b]. baroclinic [Dua99a]. barrier [AS99]. base [RRB95]. base-isolated [RRB95]. based [AG99, Cic99, Cro95, JS98b, Kha96, LR95, MFGW97, RGL97a, TSK98, VR99b, VR99c, VR99d, XS95, ZSL98]. bases [ZL98b]. Basic [CY99]. basin [BDS96, Nar99b]. basis [FS98, HH99, KR95b, Li98, PJ95]. Bayes [ADS96]. Bayesian [AB95b, Won98]. beam [HZ96]. behavior [Har97, Kar98, PGHX99, TPL95]. behaviors [IFU98]. behaviour [WA98]. beliefs [Ada98]. Bell [ESR99]. Bellman [EORA99]. bending [KWS95, XL98].

Bernstein [CS96, Jia99]. Bessel [ZB97]. bet [Fin95]. between [Bur96, Cha95b, GSI96, JY97, Van97, Waz98a, Yam95, ZK98]. Bi [MPD97]. Bi-objective [MPD97]. Bifurcation [Gra97, DVD95, GLL98, Le98, LLL97, Lui99]. Bifurcations [FM96, Tak98]. bilayer [Kub97]. binary [Ada98]. binding [Bay98a]. binomial [AAAN98]. bioactive [RBPK96]. biomolecular [ZS98]. Biophoton [BKS98]. Biorthogonality [Khu98a]. biped [IOAB95]. biphasic [HLXZ97]. black [Zha96b]. Block [BT97, Bai97, Bai99b, Die96, Jin95]. block-triangular [Die96]. blocks [Ada99]. Blow [CZ99]. Blow-up [CZ99]. blurred [FG95]. Board [Ano95a, Ano95b, Ano95c, Ano95d, Ano95e, Ano95f, Ano95g, Ano96a, Ano96b, Ano96c, Ano96d, Ano96e, Ano96f, Ano96g, Ano97a, Ano97b, Ano97c, Ano97d, Ano97e, Ano97f, Ano97g, Ano97h, Ano98g, Ano98h, Ano98i, Ano98j, Ano98k, Ano98l, Ano98m, Ano98n]. body [AA99b, Tuc95]. Boltzmann [MBS97]. Boolean [GSI96]. Booster [NR99a]. Bootstrapping [GSI96]. bordered [Bai97, Bai99b]. bound [Wei99a]. boundaries [Nar99b]. Boundary [BT97, Guo99, KA95, Pei96, AW99b, AD98, AG99, AC97, AY98, BKM95, BDJR98, Cab94, Cab95, CZ99, Cha96a, CW97, GG96, Gol96, Gup98, Had95, HG99, Jia97, KR95a, KBS96, Mit97, MR99, NR98, Pei95, QA98, RGL98, Ram98d, SZ97, Sev99b, VD97, Van95, VS96, VR99f, Wan99a, Wan99b, Waz97b, Waz99b, WCG97, WA99, YK97a, HH96b, HH97]. boundary-layer [Had95, Waz97b]. boundary-value [VR99f]. Bounded [JC97b, Che96b, JPF95, Sor99]. boundedness [FL99]. bounds [CLC97, Eis98, QA98]. Boussinesq [DH99a]. box [FpDS99]. box-counting [FpDS99]. brain [Kje99]. break [Szi98]. break-down [Szi98]. breakthrough [Mus97b]. brittle [LK95]. Brouwer [YL96]. buoyancy [SK96, ZW95a, ZW95b]. Burgers [DVD95, HM98]. Burridge [Ado96a].

CA [BR99a]. cable [BD95]. calculating [EMEH96]. calculation [YBR97, Zen99]. calculus [AK96, CS97, CSY98, dRS98]. California [Fin95]. Camassa [Boy97b]. cancer [DGU96]. Capacity [MZ98]. capturing [Dav98]. carbon [RBPK96]. Carlo [QC98]. Cartesian [CS99]. case [Rin98, SS98b]. cases [MS95a]. Cauchy [Beh96b, TETH98]. caused [zHhT99]. cavity [AAAaZ99b, SK96]. Cayley [MSZ96]. CDC [XZL97]. cell [KKL99, MMMK96, XZW96, XZL97]. cellular [Ada97, Ada99, GMS96]. Center [SST95, Lia98]. centered [KKL99]. centers [PK98]. cerebrospinal [STD98]. certain [KR95b, MS95a, PGHX99, Wol98, WLTS99]. chain [JA99]. Chandrasekhar [DK96, Sug95, SO95]. change [Gho96, MBG95, WC99]. channels [HF95, HFR98, Ram98d]. chaos [LG99, WL97]. Chaotic [Son96, GKT96, MMC98]. Characteristic [Gu99, KP98]. characteristics [MW98, QC98, SAL96]. Characterization [Szi99]. Chebyshev [BPE99, DVD95, nHyG97, HS98, yHyG99, ZB97]. chemical [Lar97, NR99a, Sel95]. chemostat [BR98]. choice [EFS95]. Cholesky [Zha98d]. chromatic [Kor97]. circuits [MNK99]. circulant [Jin97, JS98a, RG98]. Circular [Car97, AA95, SL99]. circulation [ÖGDB98]. claims [BS98]. class [AG99, AK96, AB95b, BS97, Bai99a, Bai99b, Che95, CSY98, Eis98, Gar98, Ham98a, KBS96, KR95b, LLX96, RC99, Rya96, Sas95a, Wan99a, Wan99b, Wol98, WLTS99, YL96, Zha98c]. Classes [DS99, HHT98]. classical [CC99a, Sai99]. classification [ÖGDB98]. Classifier [IFU98]. climate [MBG95]. Closed [NFJ94, Lin96]. clusters [BI98]. codes [GMGCL99, GBS99]. coefficient [BB96, HZL97, TG97]. coefficients [FBSS96, WJO97]. Coexistence [BR98]. collisions [TPS98]. collocated [Kum95]. collocation [BPE99, Bla96, FS98, GL95]. collocation-type [GL95]. combination [He98, Jia99, QC98]. combinatorial [GS97]. Combined [Lu98, EM98, LYF97, ZYB96]. combustion [LS98b]. commonality [ITN98a]. communication [WW99a, WW99b]. compact [Ram98a]. Compactly [Wu97]. comparative [KA96b, Waz99b]. comparison [Beh95, KSZ^{+}95, MBS98, Waz98a, Zha95]. comparisons [JK98]. competing [MFGW97]. Competition [ZK98, FW98, LG97, SSW99]. competition-diffusion [FW98]. competitive [AdO98]. complementarity [Bai98, Lin96, Zha99b]. Complete [ITN98a, Koz95]. completeness [Kor97]. Complex [SY98, Tak98, TSK98, Waz97c, ZB97]. componentwise [SBH96]. composite [AAAaZ99a, ENAAA95, FH98]. compositions [Ada98]. compound [Ram99a, Rol98]. compressible [HRA98]. Comput [Cab95, Mus98]. Computation [BJTW98, WSL99, Dég96, GS95, LLL97, LLX96, LUL97, RR99, Wan96]. Computational [Cos99, HRA98, GSZS98, Ham98b, JR95, MFGW97, NR98, Sel95, Sev99a]. computations [HSA97]. Computed [PZ99]. Computer [DGU96, GMGCL99, BR99a, Bay98b, Bay98a, BY99, KKPL96, TG97]. computing [Abu99b, Bil96, Che96a, JCC98, SY98, XZL98]. concave [Van95]. concavity [HS99]. concept [SS98b]. condition [Khu98a, Lin96, Wu97]. conditions [AD98, Arg97a, Arg98, Arg99b, Cab94, Cab95, CZ99, GG96, ITN98b, Ino99, JPF95, KBS96, RGL98, Ram98d, SZ97, VD97, Van95, Waz97a]. conduction [CL98a, Ona97]. conductivity [CBYK99, KBÖ99, YS98b]. Conference [Ano99e, Ano99f]. confidence [WLB97]. conflict [BDS96, HKFP97, ITN97, RPT97]. conforming [Lee98, XL98]. conjugate [Mou95, PZ99]. conjunction [ZL98b]. connecting [BD95]. connectivity [CS99]. Conservative [HSA97, FPGV95]. Consistency [Beh96a, JK98]. consistency-driven [JK98]. constant [BNN95, LG99]. Constitutive [LK95, BNN95]. Constrained [KUX95, Let97, Fab96, RL96, Sun98, UK95]. constraints [Gzy97, IKU96]. constructing [Arg96a, Arg98, Arg99d]. Construction [Boy97a, NLC95, EGH97, EGH98, Jin97, JS98a, JPF95, TJ95, BK99]. Constructions [Sas95a]. consumer [SMY95]. contact [Abd99]. contaminant [HY98, HY99b, LS98a]. contemporary [MA95]. contents [Ano97u, Ano95o]. Context [SAL96]. Context-free [SAL96]. contingent [BS98]. continuation [ARM97, McG98]. Continuous [SO95, CT96, Jan95, MD98, Pei99, ZK98, Zha99b]. Continuous-time [SO95]. contraction [Han98]. Control [Che96b, Chu97, AM98, BD95, Cha95b, Che95, Chu96, EM98, Fab96, GL99, Hab95, Kum95, Lam98, LR95, Lei95, Lia97, Liu95, Mil95, MZV96, Pap95, Per96, PJ95, RL96, RRB95, RB95, SST95, SWL98, SMY95, UTL^{+}95, UHL96, VLP95, WSL99, ZW95a, ZW95b]. controllability [Chu98]. Controllable [Sha97]. controlled [TH95]. controller [KH95]. Controlling [KP96]. controls [IOAB95]. convection [Had95, YS98a]. Convergence [Arg99a, CL99, HW98, JN95, MLB99, YJ99, Arg97b, Arg99c, BS97, CN97, Cha96a, EGH97, EGH98, FHG96a, FT97, Gao98, GB98, KBS96, XL99, Yan98, ZS99]. convergent [HZL97, SS98a]. convex [Cro95, LYF97, PS96, Sor94, Sor97a, XZL98, Zha97c]. convexity [HS98, HS99]. conveying [ST96]. convolution [GS97, WP97]. coordinate [Ham97, FHG96a]. corner [HHZ97]. correction [JNF95, JNF96, PGHX99, VR99e]. corrections [VR99f]. correlated [Ami99, Kha96, Mil95]. corrosion [BDJR98]. coshoidal [Boy97b]. Couette [Uen95]. counterions [Kub97]. counting [FpDS99]. Coupled [Ado96b, FH98, Ham97, RRB95, UR97]. couples [Rin98]. Crank [HM99]. criteria [AFM97, FL99, HM99, HH96b, HH96a, HH97, MS99, Pei99, Sol98, Wan99b]. Criterion [HKL98]. Critical [Yam95, HF95, TJ95]. cubic [Beh95, EB99, Khu98b]. cubical [EGH97, EGH98]. culture [ZK98]. curved [Ham97]. curvilinear [HFR98]. cutting [PS96]. cycle [Kwa96, XL98, XL99]. cycles [Alw97, ST96]. cyclic [WK95]. cyclically [MS95b]. cylinder [AA95, AA96, AA99a, ENAAA95]. Cylindrical [Yan99, MS95a]. cylindrically [AAAaZ99a].

D [RGL97a, Ada99, Cha95a, LS98a, Ram97a, SK96, Van95]. D-domain [Van95]. damage [LK95]. damped [BKS98, JC97a, Kar98]. damping [Bar95]. Darboux [CZ99]. data [GSC95, KUO95, Kha96, KL99b, MMC98]. data-fitting [MMC98]. Dawson [Let97]. deadlock [ITN98b]. decaying [CG98]. decentralized [HAST98]. deception [ITN97]. decision [HKFP97, PN97b, RPT97, YH97]. decomposable [MS99]. Decomposing [HT98]. decomposition [ARM97, BS97, DK96, Gol99, He96, Khu98b, LB96, MLB99, Waz95a, Waz95b, Waz97a, Waz98a, Waz99c, Waz99d, XS95, Ye98b, Ye98a]. decomposition-type [BS97]. defined [GSI96, Ibr96]. Defining [PLPW96]. definite [LLX96]. deformed [HW98]. degenerate [CY98, CC99a]. degree [EB99, Kor97]. delay [Chu97, Chu98, HG99, HH96b, HH96a, HH97, LL96, She96, Wan99a, Wan99b, YHXX95, ZA96, ZAES96]. delay-differential [HH96a, HH97]. delayed [WZ95]. delays [AR98b, FW98, Kum95, Lu98]. denominator [KKH95]. dependent [AAES99, AKK99, Szi98, ZA96]. Deployment [BD95]. derivative [GSC95]. derived [Beh96a]. described [MvL98]. describing [TETH98, TET99]. Design [HZ95, Che95, HZ96, MZ98]. designing [YL98b]. destabilization [OMN95]. detecting [JY97]. Detection [Gho96, KK95, SO95, WC99]. detector [MFGW97]. determinants [CS98, SC96]. determination [BPE99]. determine [SP95]. determining [XZW96]. developments [Gol96]. deviating [LF99, Pei99, WA98]. deviation [Cha96b]. diagram [HT98]. diamond [Kan97, Kan99]. dice [SM95]. difference [AW99a, Asa98, BT98, BS98, Hos99, HY99a, JNF95, JPF95, JNF96, KTZK96, KBS96, KBÖ99, KKL99, Mor98, MR99, NFJ94, Pei95, Pei96, RGL97a, RBPK96, Sim99b, SS97, TPL95, VC98, VR99a, VR99b, VR99c, VR99d, WA98, Zha95]. differences [JY97, Yil96]. differencing [FHG96a]. different [CL99, KKPL96]. differentiability [Arg99b, Ibr96]. Differential [CL98a, KTZK96, Ado97a, ASK98, AY98, BKM95, BD96, Béc98, CG96, CH96, CH99, Chu97, DPTT98, ESI95, FS98, FL99, GG96, GGS97, Gra95, GSZS98, Guo99, Gup98, HZL97, HG99, HH96a, HH97, JC97a, KA95, KW96, KW97, LF99, MD98, MRJ95, NS96, NJC99, PF99, Pei95, Pei96, PGHX99, Ram98b, Ram98c, Ram99c, RAL95, SJ96, SKG97, She96, Sol98, SK98b, SIA96, TET99, WZZ98, WA96, XYY96, Yan97, YHXX95, ZAES96, Zha98c]. differential-difference [Pei95, Pei96]. differintegral [AK96]. diffusion [Ado97d, AR98b, FW98, Gar98, JV98, KP98, Ram97a, Ram97b, Ram98a, Ram99b, Sai99, SZ97, Wan98, ZL98b]. diffusion-reaction [Gar98]. diffusive [Ram99d]. Digamma [WLTS99]. dilation [TyG99]. dilution [MMMK96]. dim [KK95]. Dimension [Sol99b, FpDS99]. dimensional [Ano97i, Asa97b, CH99, Dua99a, nHyG97, HK96, yHyG99, JBV98, Kan97, KKH95, KD95b, Mus97a, Mus97b, Mus98, Nar99a, NS97, OU99, ST96, Waz98b, Zha97c, ZMV99]. dimensions [GL99]. direct [Ona97, VR99e]. direct-correction [VR99e]. directions [CK99]. Dirichlet [Van95, Won98]. disaster [zHhT99]. discontinuous [RC99, WJO97]. Discrete [Fri97, JMS98, AP95, CT97, EFS95, FW98, Gol95, GB98, He96, HSA97, HH96b, Kum95, LS99, Lia97, MS99, MB95, Sug95, SMY95, SS98b, WP97, Zen99, ZYW97]. discrete-delay [HH96b]. discrete-numerical [HSA97]. Discrete-time [Fri97, Lia97, Sug95]. discretization [LY98, Ona97]. discretizations [Arg97a, Ye99]. discretized [XL98]. disease [Alw97, MAS95]. disparity [TETH98]. dispersive [Hos99, HY99a]. displacement [YBR97]. displacements [dBL95]. dissipation [Dua99b, OMN95]. Dissipation-induced [OMN95]. dissipative [WW98]. distance [RG98]. distortion [Abu99a]. distributed [Ada98, Bil96, Che96b, LF99, RAL97, WA98]. distribution [BB96, Fin95, Moh99a, Won98]. distributions [Bur96, GBS99, JY97, MMT99, Rol98]. diversion [RPT97]. Domain [He96, Ye98b, Ye98a, Jac95, KR95b, LB96, MLB99, Van95, ZL98a]. Domains [Kur99, Guo99, Ham97, HHZ97, Jia97, JPF95]. double [RAL97]. doubly [MS97]. doubly-stochastic [MS97]. down [Szi98]. Drazin [Wei98]. driven [JK98, Mil95, SK96]. duality [Cha95b]. due [AA95, AA99b]. Duffing [WL97]. during [Bar95, Tuc96, YBR97]. Dusty [HS96, AH99b]. Dynamic [LSY98, AA96, AM98, Bil96, BGY97, CT96, CT97, DF96, Hab95, Kum95, MPD97, SMY95]. dynamical [Fri97, GKT96, LR95, Lei95, LS99, OMN95, RR98, Sol99b, SAF^{+}97, Son96, Tuc95, Unb96]. Dynamics [Hal98, UR97, Ado97d, BK99, Chu96, Dua99a, HSA97, IKP97, Kaz99, Lam98, Rin98, UTL^{+}95, UHL96, Uen95, ZA96, ZAES96].

early [JY97]. ecological [UR97]. economic [Chu96, Chu97, Chu98, Lin96, SAF^{+}97, TETH98, Van97]. ecosphere [SAF^{+}97]. ecotoxicology [AB95a]. Editorial [Ano95a, Ano95b, Ano95c, Ano95d, Ano95e, Ano95f, Ano95g, Ano96a, Ano96b, Ano96c, Ano96d, Ano96e, Ano96f, Ano96g, Ano97a, Ano97b, Ano97c, Ano97d, Ano97e, Ano97f, Ano97g, Ano97h, Ano98g, Ano98h, Ano98i, Ano98j, Ano98k, Ano98l, Ano98m, Ano98n]. effect [AA99a, BR99b, Kub97, Raj97, SEG96]. effectivity [HD96]. Effects [FHG96a, Kum95, CY98]. Efficient [JBV98, HM98, Lia97, RC99, SK98a]. eigensystems [AAES99]. eigenvalue [AAES99, ABW99, CH96, VD97, Van95]. Eigenvalues [WA99, GG96, Lui99]. eigenvectors [LLX96]. Elastic [KD95a, KD95b, AAAaZ99b, AAA99, AA99c, Che96b, Le98, LK95]. elasticity [Lee98]. elastodynamics [LM95]. electric [JS98b]. electrical [KBÖ99, Sev99b]. electromagnetic [Ado96b]. electronic [MNK99]. electronics [SP95]. element [Cha96a, CW97, CL98b, DZ99, Gol96, Gu99, nHyG97, HY98, yHyG99, KA95, KWS95, KL99a, KP98, Lee98, LC97, Lin98, Moh99a, Ohm98, VD97, Van95, Yan98, Ye98b, Ye98a]. elementary [HZ95, LS99, Sas95b]. Elements [BR99a, Gri98, HZ95, HZ96, XL98, YK97b]. elliptic [AD98, AZK96, CG96, Cen97, GL99, HHZ97, Jia97, KA95, SKG97, Sri95, VD97, WJO97, Yan98, Ye98a]. elliptic-type [AZK96, Cen97, Sri95]. embedding [LFS95]. emission [BKS98]. enclosures [Wol98]. enforcement [FHK97]. engineered [AS99]. entire [KR95b]. Entropy [FT97, Din98, FT98, Tag99, ZYW97]. Entropy-convergence [FT97]. entry [Tuc96]. environmental [FHK97, HKFP97, Van97, YH97]. environments [LG97]. enzyme [Bay97, Bay98b, BY99]. epidemic [BGY97, DML98, Kim96, KP98, MAS95]. epidemiology [IKP97]. Epstein [AZK96, Cen97, Sri95]. equation [Abd99, Ado96c, Ado97b, AR98a, Asa98, AY98, BK99, Boy97b, BT98, DVD95, DH99a, DK96, DZ99, nHyG97, HM98, yHyG99, JBV98, KW96, KW97, Khu98b, Khu98c, LM95, Ona97, Pei95, Pei96, Ram98b, Ram98c, RC99, Sai96, Sai99, Sim99a, Sim99b, TET99, Tak98, TG97, UK95, VR99b, VR99c, VR99d, Wan99a, Waz97b, Waz98b, Waz99c, WL97, WCG97, Yan99, Zha96b, Asa97a]. Equations [Mur96, Ado96b, Ado97a, Ado97c, AW99a, AR98b, ASK98, AK96, AC97, Bai99a, BKM95, BD96, BNN95, Bay98a, Béc98, Bla96, CY98, CZ99, Cha95b, Cha95a, Cha96b, Che97, CH99, CC99b, Chu97, DEV95, Din98, DPTT98, ESI95, ES99, FS98, FHG96b, FL99, GG96, GGS97, Gol99, Gra95, GSZS98, Gu99, Guo99, GL95, Gup98, HZL97, HG99, JV98, JN95, KA95, KTZK96, Kar98, LL95, LY98, LP97, MRJ95, Mor95, MBS98, NS96, NJC99, NS97, NLC95, OM99, Pap95, PF99, Pei99, PGHX99, Ram97a, Ram97b, RGL97a, Ram98a, Ram99b, Ram99c, RAL95, SJ96, SKG97, She96, SS98a, SK98b, SIA96, SS97, TETH98, TPL95, VR99f, VR99a, VR99b, VR99c, VR99d, VR99e, Wan98, Wan99b, WZZ98, WK96a, WK96b, WA96, WA98, XYY96, XZW96, Yan97, Ye98b, Ye99, YHXX95, ZAES96]. equations [Zha95, Zha98c, ZW95a, ZW95b, ZL98b, ZMV99]. equicorrelated [KB98]. equicorrelation [BB96]. Equilibrium [AD98]. Equivalence [Zha97b]. equivalent [VR99b, VR99c]. ER-materials [LR95]. Erdélyi [ASK98]. Erratum [Ano97i, Cab95, JNF95, Mus98, Sor97a]. Error [CLC97, GL95, Hua99, QA98, Tuc96]. errors [BBB97, Kha96]. estimate [CLC97, Sas95b]. estimates [GKT96, Hua99]. Estimating [AAAN98, Fin95]. Estimation [Ami97, Ami99, WJO97, BY99, CyGK99, Cha95b, CG98, Eis98, FBSS96, FpDS99, GSC95, HD96, KA96b, MFGW97, SU95, Whi97]. estimators [ADS96]. etch [QC98]. Euclidean [RG98]. Euler [Boy95, CZ99, CSZ98, VR99a, VR99b, VR99c, VR99d]. European [CY99, Váz98]. evaluate [BR99b]. Evaluating [FHG96b, ZS98]. evaluation [CS98, CS96, Gol96, GMGCL99, SC96]. evaluations [CRS99]. evidential [KK95]. evolution [BKS98, Cha95b, HAST98, Kje99, Pap95, YL98b]. evolutionary [ES99, IFU98]. Evolving [JL97]. Exact [Jan95, RR99, Ham98a, JPF95, WP97]. exactly [LMSS97]. example [Chu98]. Exceptional [Zha97c]. excitation [EB99]. exclusion [XZW96, XZL97]. exhibiting [NR98]. Existence [AW99b, CC99a, Gar98, Kim96, Sor99, CY98, JS96, Lin96, Sor96, JPF95, MR99]. exoatmospheric [CO96]. expansion [GL95]. experimentally [Abu99b]. experiments [Jac95]. explanatory [Ami97]. Explicit [Ado97a, ASK98, KR95a, Sim99a, TS98, Hos99, PF99]. explosion [CK95]. exponent [WL97]. exponential [Ami97, CMM96, Gho96, JMS98]. exponentially [Ram99d, Sim99a]. exponentials [CG98]. exponents [Kaz99]. Extensions [GS95]. external [Raj97, Van95]. Extinction [AdO98]. Extrapolation [HHZ97, KKPL96]. Extremum [Sor94, Sor97a]. eye [EM98].

factorization [Ram98a, Ram99b, Zha98d]. factors [HI97]. Falkner [Asa97a, Asa98]. families [Gho96, Mus97b, Zha97c]. family [Ano97i, BT98, HZ98, MW96, Mus97a, Mus98]. farms [HYL97]. Fast [ZL98c, BPE99, Jay96, Jin95, MOP^{+}97, Tuc95]. faster [Arg98]. feasibility [Cro95]. feasible [CK99]. feedback [WSL99]. feedforward [WSL99, MZ98]. Fériet [MS95a]. Fermi [Waz99c]. field [SEG96]. fields [SK95]. film [KH96]. Filter [CO96, Cic99, KK95, Sug95]. filtering [SU95, DH99a]. filters [KKH95]. filtration [SK95]. Finding [XYY96, Die96, HZ98, XZL97, ZS99]. Finite [Moh99a, VD97, Asa98, BT98, BS98, Cha96a, DZ99, FG95, Gri98, Gu99, nHyG97, HK96, HY98, Hos99, HY99a, yHyG99, HZ95, HZ96, KL99a, KBS96, KP98, KBÖ99, KKL99, Lee98, LC97, Lin98, LS98a, MS97, Ohm98, RGL97a, RR99, RBPK96, Sim99b, Van95, XL98, Yan98, Ye98b, Ye98a, Yil96, YK97b, Zha97c]. finite-difference [Asa98, BT98]. finite-dimensional [HK96, Zha97c]. first [BKM95, Cab94, Cab95, CN97, GL99, HG99, MR99, VS96, Yan98, YL98a, YBR97]. first-order [Cab94, Cab95, GL99, VS96, Yan98, YL98a]. fitted [Ram99d, Sim99a]. fitting [MMC98]. five [Zha96b]. five-point [Zha96b]. Fix [Wu97]. fixed [BR99b, Wan96, YL96]. fixed-effect [BR99b]. fixed-point [YL96]. flares [zHhT99]. flat [Mit97]. flatness [JC97b]. flexible [LFS95]. flood [zHhT99]. flow [AH99a, AH99b, FH98, Had95, HF95, HS96, Ham97, HFR98, KW96, KW97, Khu98a, KL99b, Sel95, Uen95, WZ98, HF95]. flows [BJTW98, Gre97, Ham98a, HRA98, Ohm98, PN97a, Raj97, Ram98d, TSK98]. fluctuating [Raj97]. fluid [AH99b, Ham97, HRA98, KH96, KD95a, KD95b, KW96, Kub97, Ohm98, PN97a, STD98, ST96, Waz97b, WZ98, ZW95a, ZW95b]. fluids [Had95]. fluorescence [HRA98]. flux [Cos99]. FM [Abu99a]. Focal [NS96]. Fokker [ZMV99]. follower [OMN95]. food [JA99]. Forced [Wan99a]. forces [Che96b, OMN95, ZS98]. forcing [Dua99b]. forecast [zHhT99, ZPY99]. forecasting [CT96, CT97]. forest [KUO95]. form [FH95, LY98, NFJ94, Ye99]. formal [FHG96b]. formation [Gre97]. forms [Die96]. formula [CSZ98, Zen99]. formulation [XZL97]. formulations [KSZ^{+}95, Lam98]. forward [VR99a, VR99b, VR99c, VR99d]. Fourier [Abu99b, Gzy95, Gzy97, Kan99, KWS95, RAL95, RAL97, Zen99]. fourth [Pei95]. fourth-order [Pei95]. fractal [FpDS99]. Fractional [CS97, ES99, Nar99a, ASK98, AK96, CSY98, ESI95, ESR99, KVT98, dRS98]. Fractional-order [ES99]. framework [PMMA97]. Fredholm [GGS97]. free [BJTW98, CS96, Kor97, SAL96, Sai96]. freedom [EB99]. freezing [QA98]. frequency [BKS98, Fin95, Jac95, Reg96, SEG96, SK95]. Frobenius [Din98, DZ99, MS97]. frontier [LSY98]. Function [ZSL98, CSZ98, DS99, HH99, JS96, JMS98, LY98, Li98, LP97, Pap96, PJ95, Sor94, Sor97a, Ye99, ZB97]. functional [Cha95b, Gra95, HKL98, KTZK96, TPL95, Yan97, YHXX95]. functions [AAES99, Ano97i, Beh96a, Boy97a, CS95, CS97, CSY98, CMM96, DH99b, DS99, FS98, Ibr96, KR95b, Koł95, Kur99, MZ98, MS99, MW96, MS95a, MvL98, Mus97a, Mus98, Ona97, RC99, Sas95a, Wu97, WLTS99, XZL97, Zha99b]. Further [Car97, MS95a, Sri98]. Fuzzy [zHhT99, KK95, ÖGDB98, WC99, IFU98].

Galerkin [Asa97b, HY99b, HM99, HY99c, MOP^{+}97, Wan98]. gamble [DH99b]. gas [AH99a, HS96, HFR98, Uen95]. gas-particulate [AH99a, HFR98]. gasdynamic [LS98b]. gases [MBS97]. Gauss [GMGCL99, KL99b, Zha96b]. Gaussian [CMM96, MMT99, Mil98]. gel [Kub97]. gel-fluid [Kub97]. Gel’fand [AR98a, McG98]. general [AK96, AB95b, MW96, NFJ94, SA96, Sun98, XZL97, ZYB96]. generalization [Boy95, CK99]. generalizations [AZK96]. Generalized [MS99, VS96, Won98, AA99b, Arg96b, Arg97a, Arg99e, Cen97, Che96a, DS99, Gup98, HZL97, He96, HZ98, JS96, Sri95, Sta99, TyG99, XS95, YJ99, Zha98d]. generated [Arg97b, Arg99c, KVT98, KTZK96]. Generating [BI98, CS95]. generation [SR97]. generators [Ada99]. genetic [CLC97, JL97, RG98, ONOT98]. geometric [AB95a, Sas95b]. geometries [AH99b]. geometry [Mit97]. geostrophic [Dua99a, Kaz99]. Ginzburg [Tak98]. Global [GLL98, LG99, CY98, Chu96, Le98, MBG95, MvL98, Wol96, Yam95, ZAES96]. globally [SS98a]. GMCR [HKFP97]. GMRES [Kwa97]. Gompertz [ADS96]. Gordon [Ado97b, BT98, WCG97]. Goursat [Waz95a]. gradient [BNN95, Mou95, PZ99, WSL99]. grain [Gre97]. Grande [BDS96]. granular [Ahm99]. graphs [BR99a, CS99, Kor97]. gravity [Ahm99, TJ95]. Gray [GBS99]. Greece [MBG95]. grid [LS98a, SR97, TSK98, Ye99, Zha98b]. ground [AKK99]. ground-level [AKK99]. groundwater [HY98, HY99b, HM99, HY99c, HY99a, LS98a, YCZ98]. Group [KR95a, JK98, Wei99b, YB99]. groups [Bil96, WW99a]. Growall [EORA99]. growth [Chu96, JA99, PN97a, TET99, ZA96, ZSL98]. Grünwald [He98].

Half [Reg96, AA99c, Abu99a, dBL95]. half-space [AA99c, dBL95]. Half-wave [Reg96, Abu99a]. Halley [HS99]. Hamburger [FT97]. Hamilton [YST97]. Hamiltonian [BT97]. Hammerstein [GL95]. Hankel [KA96a]. Harmonic [Abu97, EB99]. Hartley [Jay96]. Hausdorff [Tag99]. having [Sas95a]. Hawaii [RPT97]. head [EM98, HKL98]. heat [AA95, AA99b, CL98a, Ona97, ŞKB95]. heat-conduction [Ona97]. heavy [AKK99, TPS98]. heavy-ion [TPS98]. Helmholtz [JBV98]. hemopoietic [TH96]. hereditary [CY99, Chu96, HHT98, LS95]. Heteroclinic [ST96]. heterogeneous [LZ96]. HIC [HKL98]. hierarchical [PMMA97, TyG99]. High [LF99, Zen99, SP95, TS98, ZS99]. High-accuracy [Zen99]. Higher [GG96, AW99a, JNF95, JNF96, NFJ94, WA98]. Higher-order [GG96]. Hilbert [Cha95b, Mor95]. history [FHG96a]. HIV [HD96]. hollow [AA99a, ZL98a]. Holm [Boy97b]. holon [HAST98]. Homfly [GMGCL99]. homogeneous [AAAaZ99a, AAA99, ENAAA95, Kur99]. Homotopy [YL96, LYF97, XYY96, ZYB96]. Hopf [DVD95, GLL98, LLL97, NLC95]. Hopf-bifurcation [GLL98]. horizontal [Had95]. Hubbell [AZK96, Cen97, Sri95]. hull [BJTW98]. hulls [XZL98]. human [EM98, TET99]. hybrid [Bai97, Bai99b, ONOT98, RBPK96]. hydropower [MPD97]. hyper [Ano97i, Mus97a, Mus98, TyG99]. hyper-dilation [TyG99]. hyperbolic [BKM95, Car97, FBSS96, KTZK96, Pei99, Sev99a, Wan99a, Wan99b]. Hypergeometric [MSZ96, CS95, CRS99, DS99]. hyperquadrilaterals [Yeh97].

ideas [WK98]. Identification

[Eis98, RAL95, WW98, CBYK99, HHT98, SJ96, Sci97, Unb96, XS95]. identifier [YB99]. identify [PMMA97]. identities [GS97, MSZ96, Yür99]. II [Wan99b, Arg96b, Arg99e, Cos99, EGH98, GS95, GSZS98, HF95, KD95b, VR99b, Zha98a]. III [VR99c]. ill [DH99a]. ill-posed [DH99a]. images [KKH95]. imaging [HRA98]. imbedding [NS97]. imbricate [Boy97a]. Impact [MBG95]. implementation [FpDS99, GXQ99]. Implicit [Ram98a, DPTT98, JNF95, JNF96, NFJ94, Ram98b, VR99e]. Impossibility [ITN97]. impressing [Abd99]. improve [HYL97]. Improved [MF99, QA98]. Improvement [NR99b]. impulses [KTZK96, She96]. Impulsive [CY98, Liu95, BKM95, FL99, Guo99, HG99]. impulsively [Kar98]. inclusion [KD95a, KD95b]. inclusions [CG96]. incomplete [MS95a]. incompressible [Waz97b]. inconsistent [GSI96]. incorporating [ONOT98]. incorporation [NR99b]. increasing [Yil96]. incremental [GP97]. indefinite [YK97b]. independence [Arg97a]. independent [AP95]. Indeterminate [Har95]. Index [Ano99i, Ano99g, Ano99h, Ano99j, Wei98, Ano95h, Ano96h, Ano97k, Ano98d, Ano98e, Ano98f, Ano98a, Ano98b, Ano98c, Ano99c, Ano99a, Ano99d, Ano99b, Kor97]. indexed [MS95b]. Indices [HS99]. indirect [BPZ96]. induced [OMN95]. inelastic [BNN95]. inequalities [AP95, EORA99, KTZK96, LF99, PZ97, PK98, ZZ98, Zha97c]. Inequality [IKU96, YST97]. inexact [Arg97a, Arg99a]. infection [HD96]. inference [KB98]. infinite [AAA99, ENAAA95, Sas95b, Waz98b, WLTS99, YHXX95, dRS98]. Influence [Ahm99, BR99b, BBB97, DML98].

Information [Van98, IL95, Par95]. inhibiting [JA99]. inhomogeneous [AA96, Sai99]. initial [AA99a, AAA99, BBB97, CN97, ENS96, GKT96, NJC99, RGL97b, TS98, WCG97]. initial-boundary [WCG97]. initial-value [RGL97b, TS98]. Injury [HKL98]. input [ABT98, AA96, Moh99b]. input-output [ABT98]. instabilities [HSA97]. instantaneous [AA95]. integer [xZ98]. integrable [BK99]. Integral [Abd99, AG99, ASK98, AZK96, Bla96, Cen97, DEV95, EORA99, ES99, Gol95, KVT98, Khu98c, LM95, Let97, OM99, Sri95, WR96, WK96a, WK96b, ZL98c]. Integral-based [AG99]. Integrals [CSZ98, Beh96b, Koz95, MS95a, Waz97c, Wol98, Yeh97]. Integration [BNN95, ARM97, LC97, LM98, MF99, MRJ95, RAL95, TS98, WZZ98]. integro [FL99, GGS97, Guo99, TET99]. integro-differential [FL99, GGS97, Guo99]. integro-partial [TET99]. Integrodifferential [NS97, Sol99a]. Intelligent [SWL98, MFGW97]. interaction [AA99b, Yam95]. interactions [JFW96, WZ95]. interconnections [SP95]. interdependent [ITN97]. interdimensional [Ano97i, Mus97a, Mus98]. intergrid [KL99a]. interior [LYF97, ZYB96]. Intermediate [RGL98]. intermittently [Kar98]. intermodulation [Abu97]. internal [PJ95]. interpolating [Bur96]. interpolation [Dég96, GSC95, Jia99, MvL98]. Interval [TH96, Wol96, Wol98, Eis98, LL96, Sha97, ZYW97]. intervals [WLB97]. intervention [BGY97]. Intrinsically [Béc98]. Introduction [SU95, Sug98, Udw95, UTL^{+}95, UHL96]. invariant [LS99, NS97, Tak98]. Inverse [Lam99, Tag98, HZL97, Lin98, MB95, Sev99a, Sev99b, SY98, SP95, Wei98, Wei99b, YS98b]. inverses [Che96a, Sta99]. inversion [Hai97]. investment [OU99]. inviscid [PN97a]. involving [ASK98, CS97, CSY98, CSZ98, Rol98, Waz97c, Yür99, dRS98]. iodine [HRA98]. ion [TPS98]. irrigation [MPD97]. isolated [RRB95]. isoparametric [LC97]. isotropic [AA95, AA96]. issue [GSZS98, SU95, Sug98, Udw95, UTL^{+}95, UHL96]. Italian [Chu98]. Iterated [ZSL98]. iterates [Nar99a]. iteration [Bai97, Bai99b, SL99]. iterations [CG96, ZS99]. Iterative [Che96a, Che97, Pei95, CN97, EGH97, EGH98, FBSS96, JN95, LS95, Lu98, Zha97a, Zha98c]. IV [VR99d]. IVPs [KKPL96].

J. [Di 98]. Jacobi [CS95, YST97]. jet

[JFW96]. jets [Ram99a]. Jones [EMEH96]. July [Cab95]. jumping [IKU96]. Junction [RB95, SEG96]. Jupiter [Tuc97].

Kadomtsev [Ado96c, TG97]. Kakutani [Wan96]. Kampé [MS95a]. KdV [Ado97c]. Kernel [Kan99]. kernels [Jin97, JS98a, NS97]. kind [WK96a]. kinds [ESL98]. kinetic [Bay97, BY99, RR98]. kinetics [Bay98b]. Klein [Ado97b, WCG97]. Knopoff [Ado96a]. knot [Beh95]. Kober [ASK98]. Kobori [LS97]. Korteweg [RC99]. Kramer [Sai96]. Kutta [PF99, Sim97, SIA96].

lacunary [FT98]. Ladik [TPGV97]. lag [Boy95, Sim99b]. lag-averaged [Boy95]. Lagrangian [Lam98]. land [AFM97, PMMA97, RPT97]. land-water [RPT97]. Landau [Tak98]. Laplace [Gzy95, Gzy97, Hai97, MNK99, Yan99, Yür99]. Large [Cha96b, Bai98, BNN95, KWS95, KSZ^{+}95, ÖGDB98]. largest [AAES99]. lattices [DH99a]. laws [Chu96]. layer [AAA99, Had95, Waz97b]. layered [ENS96, dBL95]. layers [FH98, NR98]. lead [YHXX95]. learning [ESL98, IFU98, Per96]. Least [AB95a, Cha96a, SUJ98, SSY99, Whi97, Yan98, YL98a, Ye98b, Ye98a]. least-square [Ye98a]. Least-squares [Cha96a, Ye98b]. length [HF95]. level [AKK99, LL95, LY98, YK97b, Zha97a, Zha98c]. Levitan [AR98a]. Liapunov [Sol98]. libration [Tuc97]. Lidstone [WA99]. Life [Sug98, JS98b]. Lifelike [ZSL98]. Lighthill [Boy97a]. like [Arg96a, Arg96b, Arg97a, Arg99a, Arg99d, Arg99e, Beh96a, GP97, Hua99, Yan99]. Limit [Alw97, Sta99]. limiting [MMMK96, ZK98]. line [AA99b, Hai97, HT98, dBL95]. linear [Abu99a, Bai97, Bai99b, Béc98, BT97, Cab94, Cab95, Cha95b, Cha95a, Che97, DPTT98, Eis98, GS99, GKT96, He98, HI97, JY97, Jia99, KB98, LMSS97, Lee98, LM98, Mor95, MR99, NS96, Pap96, PN97a, PK98, QA98, RAL97, Reg96, Rin98, Sai99, Sug95, Wei98, Wei99a, YK97a]. linearities [EB99]. Linearization [Ram97a, Ram97b, RGL97a, RGL98]. Linearized [Ram99b, Ram99c, RGL97b, Ram98a, Ram98b, Ram98c, Uen95]. Liouville [ABW99]. lipid [Kub97]. Lipschitz [Arg97a]. liquid [Ram99a]. load [dBL95]. Local [GSC95, Kaz99, CW97, Sol99b, Yam95]. locally [MvL98]. location [Van98]. logistic [JS96, Ram98b, Ram98c, TET99]. lokahi [YH97]. long [ZS98]. long-range [ZS98]. Lotka [AdO98, LL96, MW98, SA96]. lottery [Fin95]. Love [AAA99, Rin98]. low [Ram99a, SEG96]. lowest [CL98b]. lowest-order [CL98b]. LTDRM [SZ97, ZL98b]. lubrication [KH96]. lunar [TJ95]. Lyapunov [Kaz99, WL97].

M [Di 98]. Maclaurin [CSZ98]. macroeconomic [Moh99c]. magnetized [AC97]. making [PN97b, YH97]. management [HKFP97, YH97]. manifold [Lia98, SST95]. manifolds [Sol99b, SK98b]. many [Dég96, HI97]. many-variable [Dég96]. mapping [GF96, Sor97b]. mappings [Sor99]. Maps [Ram98b, Nar99a, UR97]. Marcel [Kan99]. Marchenko [AR98a]. market [SMY95]. Markov [Har95, MS97]. Mars [Tuc96]. Massively [Dég96, HI97]. Master [Ada98]. Master-slave [Ada98]. material [AA99c, ENAAA95, HSA97, SZ97]. materials [LR95]. Math [Cab95, Mus98]. Mathematical [STD98, HD96, KSZ^{+}95, Mur96, ZPY99]. matrices [ABT98, JCC98, KA96a, LLX96, RG98]. matrix [Che96a, FHG96a, JN95, LP97, MS97, RAL95, SY98]. matrix-sign-function [LP97]. Maxentropic [Gzy95, Gzy97]. Maximum [FT98, Din98, Kor97, Tag99, ZYW97]. Maxwell [Ado96b]. Mean [GR97, AB95a, Ibr96]. measles [DML98]. measure [IL95, Van98]. measurement [Rol98, TyG99]. measures [FL99]. measuring [PLPW96, Tuc95]. mechanical [RRB95, UK95, ZL98a]. mechanics [Bel95, Udw95]. mechanism [IOAB95]. media [Ado97e, AH99a, AH99b, ENS96, HS96, LZ96, YBR97]. medium [AAAaZ99b, AAA99, Ahm99, KD95a, KD95b, Waz98b, WZ98]. meeting [ITN98b, Ino99]. memories [GKT96]. memory [ALZ99]. mesh [Arg97a]. meshes [KKL99]. method [AD98, Arg99b, Asa97a, Asa97b, Asa98, BPZ96, Bay97, Boy95, Cab94, Cab95, CW97, CT96, CLC97, CT97, CL98b, CK99, CH99, CC99b, Cro95, DEV95, DK96, Din98, DZ99, EFS95, FpDS99, GXQ99, Gol96, Gol99, GL95, HW98, He96, nHyG97, HS98, HS99, HLXZ97, HAST98, HY98, yHyG99, JR95, JV98, KR95a, KA95, KL99a, KL99b, Kwa96, Kwa97, Lee98, LYF97, LLL97, Lin98, LP97, LS98a, LM98, MW98, MOP^{+}97, MLB99, MS95b, Mou95, MBS98, NR98, NR99a, Ohm98, ONOT98, PZ99, Ram98c, RC99, SJ96, SUJ98, Sel95, Sev99a, SKG97, SK96, Sol98, Sun98, SP95, SL96, Szi97, Tuc95, Van95, VR99f, VR99a, VR99b, VR99c, VR99d, VR99e, WSL99, Waz95a, Waz95b, Waz98a, Waz99c, Waz99d, XYY96, XL98, Yan98, YL98a, Ye98a, Yil96, YBR97]. method [YL96, YJ99, Zha96a, Zha97a, Zha98d, ZYB96, ZL98b, ZYW97, GR97]. Methods [BT97, PK98, ABT98, AC97, Arg96a, Arg96b, Arg97b, Arg97a, Arg98, Arg99a, Arg99c, Arg99d, Arg99e, BS97, Bai97, Bai98, Bai99a, Bai99b, BPE99, Bla96, BT98, CL99, Che96a, Che97, Dav98, EGH97, EGH98, GG96, GGS97, GB98, GP97, Gu99, Han98, HZ98, Hua99, IKP97, KKPL96, KBS96, KP98, KSZ^{+}95, KBÖ99, LL95, LC97, Lu98, LB96, Ona97, PF99, Pei95, QC98, Ram97a, Ram97b, RGL97b, RGL98, Ram98a, Ram98b, Ram99d, Ram99c, Sim97, Sim99a, Son96, Sta99, SIA96, SBH96, TPGV97, TS98, TJ95, Wan98, WK96b, Wol96, XL99, Ye98b, YK97a, Zha98c]. microbial [ZK98]. midpoint [VR99e]. migration [TETH98, TET99]. mild [Arg99b]. Minesweeper [Ada97]. Minimal [Zha97a, Ada99, Gao98, Sim99b, Zha98b, Zha99a]. Minimax [Pap95, Let97, ZYW97]. minimize [PMMA97]. minmax [SW96]. Mises [FH95]. mismatched [Che95, RL96]. mistune [WK95]. mistuned [WK95]. Mixed [Had95, CL98b, Gra95, KTZK96, Lee98, Pei96]. mixture [HLXZ97]. mobile [SWL98]. mode [Jan95]. Model [ŞKB95, Ado96a, AKK99, Alw97, ADS96, BR99b, CK95, CT96, EM98, GQS98, GMS96, HLXZ97, HY99b, HM99, Hos99, HY99a, HD96, KD95a, KD95b, Kim96, LG97, LG99, MAS95, MW98, MBS97, OA97, PN97b, PZZy99, PJ95, Rol98, SSW99, SAF^{+}97, TPGV97, TH96, TyG99, TJ95, Váz98, Waz99a, WZ98, ZA96, ZPY99]. Modeling [HY98, Hos99, RBPK96, BDJR98, LSKH96, LS98a, ÖGDB98, UR97, Unb96]. modelling [Moh99a]. models [Ami97, Ami99, AH99a, AB95b, BBB97, Gra97, HF95, HY99c, IKP97, KB98, KP98, KA96b, Mil95, Moh99c, Raj97, STD98, WK98]. modes [WK95]. modification [Waz99d]. Modified [MMMK96, Sim97, KBÖ99, LL95, MW98, VR99f, VR99a, VR99b, VR99c, VR99d, VR99e, WSL99, Waz99c, ONOT98]. modulated [Reg96]. moment [FT97, FT98, Tag98, Tag99]. monomial [MBS98]. monotone [Cab94, Cab95, CG96, Fio95, WA96]. Monte [QC98]. most [Cro95]. motion [Dua99b, EM98, KUX95, ST96, UK95]. motivated [TETH98]. mountainous [KUO95]. moving [BDJR98, SUJ98, dBL95]. moving-boundary [BDJR98]. MSS [KUO95]. multi [Gup98, JL97, ÖGDB98, Zha97a, Zha98c]. multi-agents [JL97]. multi-level [Zha97a, Zha98c]. multi-objective [ÖGDB98]. multi-point [Gup98]. Multiblock [SR97]. multibody [SK98a]. Multicriteria [Van97, PMMA97]. Multicriterion [RPT97, BDS96]. multidimensional [Ram98a, Ram99b, XZL98, Ram97b]. Multigrid [KKL99, YO98, KL99a, Kwa96, SR97, XL98, XL99, ZS98, Zha96b, Zha97b, Zha98a, Zha98b, Zha99a]. multilayered [KD95a, KD95b]. multilevel [XL98]. multiobjective [EFS95]. Multiple [AFM97, Lui99, YH97, AS99, JFW96, LM95, MS99, Wol98, Yeh97]. Multiplication [CS96]. Multiplication-free [CS96]. Multiprecision [Jay96]. multiprocessors [MBCvdV95]. Multiquadric [HLXZ97, SKG97, ZL98b]. Multirational [Pap96]. multisplitting [Bai98, Bai99a, CL99]. multistep [SBH96]. Multivalued [ESI95]. multivariate [BB96]. mutual [Par95].

N [Jia99, Ham97]. n-phase [Ham97]. Nash [Szi99]. Natural [Ham97, Ada99, IL95].

Navier [Cha96b, GP97, Ham98a, nHyG97, yHyG99, LL95, LY98, Ye99]. navigation [Tuc96]. near [Let97]. near-minimax [Let97]. Necessary [Waz97a]. negotiable [YB99]. negotiation [LSY98, YB99]. netlet [GLL98]. network [LSKH96, SY98]. networks [HAST98, Li98, Mou95, PZZy99, PJ95, WZ95, YST97, YL98b, ZPY99, MZ98]. Neural [MZ98, GLL98, Li98, PZZy99, SY98, WZ95, YST97, YL98b, ZPY99]. neurons [ESL98]. neutral [Gra95, HHT98, HH96a, HH97, Yan97]. Newton [Arg96a, Arg96b, Arg97a, Arg98, Arg99a, Arg99b, Arg99d, Arg99e, GP97, HW98, Hua99, Sun98, SL99, SSY99]. Newton-like [Arg96a, Arg96b, Arg97a, Arg99a, Arg99d, Arg99e, Hua99]. Newtonian [Had95]. Nicolson [HM99]. nine [Mur96, YO98]. nine-point [YO98]. noise [KKH95, SEG96, Waz97a, WL97]. Non [Ado97b, AH99a, MMT99, SA96, VC98, AAAaZ99a, AAA99, AW99a, AD98, DPTT98, EB99, ENAAA95, GSI96, Had95, MR99, Sai99, SSY99, Szi99]. non-Boolean [GSI96]. non-classical [Sai99]. Non-Gaussian [MMT99]. non-homogeneous [AAAaZ99a, AAA99, ENAAA95]. non-linear [DPTT98, MR99, Sai99]. non-linearities [EB99]. non-Newtonian [Had95]. Non-perturbative [Ado97b, SA96]. non-positive [AW99a]. Non-reactive [AH99a]. non-smooth [SSY99]. Non-standard [VC98, AD98]. non-symmetric [Szi99]. nonautonomous [AdO98]. noncollocated [Kum95]. noncolocated [LFS95]. Nonconforming [LY98, CL98b, KL99a, XL99, YK97b]. Nonconstructible [Ada99]. nonconvex [YL96]. nonhomogeneous [Koł95, NFJ94]. Nonlinear [Abu99a, Ado96d, BGY97, MZV96, ONOT98, Ado97a, Bai98, Bai99a, Cab94, Cab95, Che95, CL98a, Chu96, Chu97, Chu98, DH99a, Eis98, FPGV95, Gu99, Gzy97, Hab95, He96, JC97a, Jia97, JBV98, Kar98, KA96b, LS99, Lia98, Lin96, LYF97, LF99, MOP^{+}97, Mor98, Pap95, PN97b, Pei99, PGHX99, RAL95, Rya96, SZ97, Sol99a, SS97, TET99, TH95, UR97, Unb96, Wan98, Waz95b, WCG97, WC99, XZW96, ZZ98, Zha99b, ZYB96, xZ98, ZL98b]. nonlinearities [Abu97]. nonlocal [CZ99, Jia97, VD97]. nonoscillatory [She96, TPL95]. nonsmooth [dSYS97]. nonsquare [LFS95]. Nonstandard [RGL97a]. nonsymmetric [YK97b]. norm [CK99, JMS98]. norm-relaxed [CK99]. normal [BB96, Kha96, Rol98, WR96]. not-a-knot [Beh95]. note [Ano97j, Gao98, Gol95, Gol99, Ham98b, Jin97, JS98a, Sri95, Szi97, XL99, Zha96a]. Notes [Wan96]. novel [Bay97, CSY98]. NP [Kor97]. NP-completeness [Kor97]. NR [KL99b]. number [Pav95, Sor97b]. numbers [Fin95, Ram99a]. Numerical [ARM97, ALZ99, AC97, DF96, Fab96, FPGV95, Gre97, HFR98, JFW96, McG98, Moh99b, MBS97, Ohm98, SK96, SK98b, TPS98, ZW95a, ZMV99, AR98a, Asa97a, CBYK99, DH99a, DEV95, DPTT98, Gol96, Hal98, HSA97, HLXZ97, HM98, HHZ97, IKP97, KSZ^{+}95, LC97, LLL97, LM98, Lu98, LS98b, MF99, MRJ95, NR99b, ONOT98, Pav95, RC99, RBPK96, ŞKB95, SKG97, Sim97, Sim99a, Sim99b, SK98a, TS98, TJ95, Váz98, VR99b, VR99c, VR99d]. nutrient [JA99]. Nyström [PF99].

O [ST96]. objective

[MPD97, ÖGDB98, YH97]. objectives [Van97]. objects [KK95]. oblique [Wei99b]. observational [Fri97]. observations [BBB97]. observers [Cic99]. obtain [MOP^{+}97]. obtained [Abu99b, Mur96]. Obtaining [GKT96]. occupancy [WW99a, WW99b]. ocean [Kaz99]. octonions [Car97]. ODE [VR99f]. ODEs [RGL97a, Sim97, SBH96]. off [MPD97, Van97]. offer [SL96, Szi97, Szi98]. oil [BR99b, YBR97]. One [Lei95, AAAN98, Asa97b, Gho96, JBV98, Waz98b]. one-dimensional [Asa97b, JBV98, Waz98b]. one-parameter [Gho96]. Op [DF96]. open [Lin96]. operated [KP96]. operational [RAL95]. operator [BD96, Din98, DZ99, Gol99, He98, Kan97, KL99a, MS97, NLC95, RGL98]. operator-differential [BD96]. operator-splitting [RGL98]. operators [ASK98, Arg97b, Arg99c, CS97, CSY98, Gol95, MA95]. Optical [Ado97e, Tuc96]. Optimal [AM98, GB98, Mil98, Mou96, ZW95b, BB96, Cha96a, Fab96, GL99, JS96, Lia97, RB95, WSL99, ZW95a]. optimally [KM95]. optimization [DF96, HZ95, Liu95, MBCvdV95, ONOT98, Sun98, Wol96, dSYS97]. Optimized [KKPL96, HY98]. Optimum [BBB97, HZ96]. option [CY99, Váz98]. Options [PMMA97]. orbit [BPE99, Gri98]. orbits [FG95]. order [AW99a, AG99, AAAN98, BKM95, Cab94, Cab95, CN97, Cha96a, CL98b, ES99, GG96, GGS97, Gri98, GL99, Gup98, HG99, JK98, JNF95, JPF95, JNF96, KM95, KVT98, LF99, MR99, NFJ94, NS96, Pei95, Pei96, PGHX99, SJ96, SBH96, TPL95, TS98, VD97, Van95, VS96, WA96, WA98, Yan97, Yan98, YL98a, Ye98a, YK97a, ZS99]. ordered [JY97]. ordering [KL99b]. orders [ESR99]. ordinary [GG96, Gup98, NJC99, Ram99c, SJ96, SIA96, XYY96]. organizing [JL97]. orientation [Zha99b]. Orr [MSZ96]. orthogonal [KVT98]. Orthogonality [CS95]. orthotropic [AAAaZ99b, AAAaZ99a, AAA99, AA99c, ENAAA95]. orthotropy [AA99a]. oscillating [BD96, Sim97]. Oscillation [SS97, Wan99a, WA96]. Oscillations [Gra95, Yan97, ZAES96, KSZ^{+}95, VLP95, WZ95]. oscillator [BKS98, Kar98, MOP^{+}97]. oscillators [MF99]. Oscillatory [Pei99, TPL95, Wan99b, PGHX99]. other [CRS99]. output [ABT98, Moh99b, Whi97]. Overcoming [SR97]. overhead [SW96]. overlapping [MLB99]. overlying [AAA99].

P [KL99a]. packing

[Ano97i, Mus97a, Mus97b, Mus98]. Padé [Dav98, Waz99a, Waz99c]. pairwise [JK98]. paper [Di 98]. parabolic [CY98, CC99a, Lu98]. paradigm [ESL98]. Parallel [Bai97, GMS96, BS97, Bai98, Bai99b, Béc98, CL99, Cro95, Dég96, FH98, GS99, HI97, HZ98, KKPL96, KL99b, Koł95, Kur99, SW96]. parallel-series [Koł95]. parallelizing [YCZ98]. Parameter [BY99, CyGK99, CG98, HHT98, Gho96, GKT96, KA95, MB95, Mur96, RAL97, SJ96, Sci97]. parameters [AAAN98, ADS96, Fab96, JS96, KA96b, PLPW96]. parametric [BT98]. parametrization [DF96]. parametrizations [HK96]. parasitics [SP95]. Pareto [LSY98]. part [HF95, Arg96a, Arg96b, Arg99d, Arg99e]. partial [Ado97a, BKM95, Béc98, CH99, FS98, KW96, KW97, LL96, SKG97, TET99, WZZ98, WA98, Zha98c]. Partially [DPTT98]. particular [Gol96]. particulate [AH99a, HFR98]. Partitioning [SIA96]. Pascal [Beh96a]. Pascal-like [Beh96a]. Passage [MFGW97]. Passage-detector-based [MFGW97]. pattern [Gre97]. Patterns [WZ95, Ano97i, Mus97a, Mus98, ÖGDB98]. PDEs [RGL97a]. Peakons [Boy97b]. penalty [FHK97]. pendulum [KSZ^{+}95]. perceptions [ITN98a]. percolation [BI98]. Perfect [AB95b]. Performance [AS99, Abu97, Cic99]. period [GF96]. period-to-period [GF96]. Periodic [BKM95, HG99, OM99, YHXX95, AdO98, Dua99b, GF96, KBS96, LG97, LZ96, MOP^{+}97, PF99, TS98, VS96, XYY96]. periodically [Fio95]. periodicity [FW98]. periods [WC99]. permanence [FW98].

Perron [Din98, DZ99, MS97]. persistence [FM96, LL96]. Perturbation [Wei99a, Cha96b, JR95, LO99, NR99a, NR99b, RR98, SEG96, VC98, Waz99b, Wei99b]. Perturbations [ABT98, PN97a, Sor99]. perturbative [Ado97b, SA96]. perturbed [Arg96a, Arg99d, KR95a, KM95, LM98, MF99, NR98, Rol98, WL97, Arg96b, Arg99e]. Petersburg [DH99b]. Petrov [Wan98]. Petviashvili [Ado96c, TG97]. Phase [Tuc97, HF95, Ham97, Kub97, Sim99b, YBR97]. phase-lag [Sim99b]. photons [RR98]. Picard [LL95]. Piecewise [Cha95a, RGL97b, LG99, RGL97a, Ram98c]. Piecewise-linearized [RGL97b]. pinning [Cos99]. pipes [BDJR98]. placement [Mil98]. planar [EM98, Rya96]. Planck [ZMV99]. plane [AH99b, PS96, KD95a]. planning [Van97]. plants [OA97]. plasma [AC97]. plate [Had95, KWS95, XL98]. plates [Le98]. plays [Ada97]. plume [JFW96]. point [Arg97b, Arg99c, Gho96, Gup98, KBS96, Kum95, LYF97, MR99, NR98, Tuc97, VR99f, YO98, YL96, Zha96b, Zha98d, ZYB96]. point-to-point [Arg97b, Arg99c]. points [Arg99a, NS96, Sor94, Sor96, Sor97a, Wan96, Waz99b]. Poisson [CZ99, Zha96b, ZL98c]. polar [FH95, KH96]. policing [Mil95]. policy [Mou96, PN97b]. pollutants [AKK99]. pollution [GMS96, YCZ98]. polygonal [KWS95]. polyhedral [Zha97c]. polynomial [Dég96, He98, Jia99, KA96a, MZ98, MS99, ZS99]. polynomials [CS95, CS96, EMEH96, ESR99, GMGCL99, HI97, HZ98, KVT98]. population [GQS98, IKP97, LG99, MW98, TET99, WK98, Waz99a, ZA96, ZAES96]. populations [KP98, ZK98]. poro [ZL98a]. poro-mechanical [ZL98a]. poroelastic [Bar95]. porous [AH99a, AH99b, FH98, HF95, HS96, HFR98, LZ96, Ram98d, WZ98, YBR97]. posed [DH99a]. position [KP96]. positive [AW99a, LLX96]. Positivity [LFS95]. postbifurcation [ST96]. potential [TJ95]. potentials [Kin97]. Powell [ONOT98]. power [CS97, JS98b, MZV96, SP95]. Practical [Zha98a]. precipitation [MBG95, ÖGDB98]. precisely [Mor95]. preciseness [Yil96]. precision [Van98]. Preconditioned [GP97, Kwa97]. preconditioners [Jin97, JS98a, NLC95, XL98, YK97b]. Preconditioning [CL98b, LC97]. predator [Gra97]. predator-prey [Gra97]. prediction [Kha96]. Predictive [Hab95, KB98]. preference [ITN97]. pressure [STD98]. prey [Gra97]. price [CY99, SMY95]. pricing [BS98, Váz98]. principal [Beh96b]. principle [Arg97a, LO99]. principles [Mur96]. Probabilistic [Uen95]. probabilities [JCC98, Szi98]. probability [Bur96]. problem [Abd99, Asa97b, AY98, BKM95, BPZ96, Cha96a, FT97, FT98, GP97, Gup98, HZL97, He96, HG99, JMS98, Lee98, Lin98, McG98, MB95, OU99, Pei95, Pei96, Sas95b, SJ96, Sci97, Sel95, Sev99a, Sev99b, SM95, SK96, TETH98, Tag98, Tag99, Van95, VS96, Wan99a, Wan99b, Waz95a, Waz95b, Waz99b, WW99a, WW99b, WCG97, XL98, Ye98a, YS98a, YK97a, YS98b, ZYW97]. problems [AW99b, ABW99, AD98, AG99, Bai98, Bel95, Cab94, Cab95, CN97, CC99a, CH96, CL98a, Cro95, DF96, Eis98, EFS95, Fab96, Gar98, GL99, Guo99, HHZ97, JR95, Jia97, KR95a, KM95, KTZK96, KBS96, KR95b, LMSS97, LM95, LO99, Lia97, Lin96, LM98, Lu98, Moh99a, MR99, NR98, NR99a, NR99b, NJC99, PZ99, PS96, QA98, Ram97a, Ram97b, RGL97b, ŞKB95, SZ97, SK95, SSY99, TS98, VD97, VC98, VR99f, WA99, YK97b, YL96, ZL98a, Zha98d, Zha99b, ZYB96]. procedure [EGH97, EGH98, FBSS96, MMC98, Sel95]. procedures [Moh99b, Whi97]. process [AAAN98, IKU96, SS98b, TyG99, YBR97]. processes [Hab95, LS95, Rol98]. processing [LSKH96, SW96]. producer [SMY95]. producer-consumer [SMY95]. product [CS99, Sas95b]. production [Rol98]. products [CMM96]. profile [QC98]. profitable [PMMA97]. programming [BGY97, EFS95, Han98, LMSS97, LYF97, MPD97, Pap96, PN97b, PS96, PK98, ZYB96, xZ98]. Projection [Han98, CL98b, Cro95, GB98, PS96, Wei99b]. proof [UK95, Zha97b]. Propagation [AAA99, AA99c, AA99a, Ado97e, Ahm99, ENS96]. proper [Sor99]. properties [Fio95, Sas95a, SZ97, Zha99a]. proposal [MNK99]. prospection [Sev99b]. proton [zHhT99]. pseudo [Hai97, SY98]. pseudo-inverse [SY98]. pseudo-inversion [Hai97]. pseudospectral [yHyG99, SK96]. pseudospectral-finite [yHyG99]. Publisher [Ano97j]. pulse [HZL97]. pulse-spectrum [HZL97]. pure [Lee98].

quadratic [Wan95]. Quadrature [MW96, CMM96, Mil98]. quality [HYL97]. quantitatively [Tuc95]. Quasi [SSY99, Dua99a, Gao98, Kaz99, Raj97, Sun98, SL99]. quasi-geostrophic [Dua99a, Kaz99]. quasi-minimal [Gao98]. Quasi-Newton [SSY99, Sun98, SL99]. quasi-steady [Raj97]. quasigeostrophic [Dua99b]. quasilinear [JR95, WA96]. quasilinearization [JV98, VS96]. quaternions [Car97]. queue [Jac95, MS95b, MFGW97]. queueing [Mil95, Mou95]. queues [Mou96]. Quick [SC96, CS98]. Quickly [BD96]. quotient [Har97].

Radial [HH99, PJ95, FS98, Li98, Wu97]. radiation [DGU96, SK95]. radiative [MA95]. radio [WW99a, WW99b]. railway [JS98b]. random [Ado96d, Ado97e, BR99a, BI98, Gho96, WP97]. range [ZS98]. rank [JY97]. ranking [JK98, MS99]. Rapid [CN97]. rarefied [Uen95]. rate [LK95]. rates [Arg99a, Cha96a, GB98, JN95, MS95b]. rating [MS99]. rational [Let97]. Raviart [CL98b]. Rayleigh [AA99c]. reach [ITN98b, Ino99]. reacting [MBS97]. reaction [Ado97d, Gar98, Hos99, JV98, Ram97a, Ram97b, Ram98a, Ram99b, Wan98]. reaction-diffusion [Ado97d, Ram97a, Ram97b, Ram98a, Wan98]. reactive [AH99a]. reactor [NR99a, Sel95]. Real [LLX96, LUL97, Hai97, MvL98]. Real-time [LLX96, LUL97]. reallocation [RPT97]. reanalysis [CW97]. recombination [SW96]. Reconstruction [FG95, Gzy95, Gzy97]. rectification [Abu99a, Reg96]. recurrences [SW96]. recurrent [Ino99, SY98]. Recursive [Hai97, WW99b]. red [Zha96b]. red-black [Zha96b]. reduced [WSL99]. reducible [MS95a]. reduction [KKH95, LO99, Lia98, ŞKB95]. Refined [HM99]. refinement [CW97]. Reflection [MA95]. reflectionless [Kin97]. region [Khu98a, SSY99, dSYS97]. regional [TETH98]. Regression [Kha96, AB95a, BBB97, KB98, KA96a]. Regression-based [Kha96]. regular [JC97b]. regularization [DH99a, PZ99]. regularized [Whi97]. regulations [FHK97]. related [CS95, Sta99]. relation [GSI96]. relations [Beh96a, CS95, SAL96]. relativistic [TPS98]. relaxation [BS97, Bai98, Son96, Zha96a]. relaxed [CL99, CK99]. reliability [Ami97, AB95b, Koł95, Kur99]. reliable [WK96a, Waz98b, Waz99d]. remark [Dua99a]. Remarks [Cab95]. remote [Cro95]. remotely [KP96]. repeated [RAL95]. representation [Gol95, CS96]. representations [Sri98, Sta99]. research [DGU96]. Residual [Zha97b, Zha98a, Gao98, RBPK96, Zha97a, Zha98b, Zha99a]. residual-finite [RBPK96]. residuals [CLC97]. residue [Cen97]. Resolution [Whi97, FG95, GP97]. resource [ZK98]. Response [EB99, AA96, JC97a]. restricted [Che97]. result [LS99]. resulting [WW99a, WW99b]. Results [Arg96a, Arg96b, Arg99d, Arg99e, Bla96, Car97, Gar98, Le98]. retarded [AY98, PF99]. retrieval [KUO95]. Reversible [OMN95]. revisited [KUX95, VC98]. revolution [BK99]. Reynolds [Ram99a]. Riccati [JN95, LP97, MRJ95]. Riemann [LS98b]. Riesz [Kan99]. Rio [BDS96]. rise [PMMA97]. robot [IFU98]. Robotics [Sug98]. Robust [Che95, LL96, RL96, KH95]. root [HZ98]. root-finding [HZ98]. rotation [AAAaZ99a, BNN95, ENAAA95, Pav95]. rule [Yeh97]. rules [DH99b]. Runge [PF99, Sim97, SIA96]. rural [Van97].

S. [Jia99]. saddle [Zha98d]. salinization [PMMA97]. satellite [KUO95, SST95, VLP95]. satisfied [Mur96]. saturated [Bar95]. saturation [YBR97]. scalar [JBV98, KM95]. scale [ÖGDB98]. scaled [KM95]. scales [ABW99]. scaling [Zha97b, Zha98a]. scan [HT98]. scattered [GSC95]. scattering [Ado96b, KD95a, KD95b, LM95]. scheme [CW97, FPGV95, HM98, JS98b, LR95, Sim99b]. schemes [CL99, DPTT98, FHG96a, MW96, VC98]. Schrödinger [BK99, FPGV95, He96, Khu98b, Sim99a, Sim99b]. Schrödinger-type [He96]. Schwarz [YK97b, ZZ98]. screening [Kim96]. search [HI97]. Second [JPF95, AG99, Cha96a, GG96, GGS97, Gup98, KM95, NS96, PGHX99, SJ96, Sol98, TPL95, Van95, VR99c, WK96a, WA96, Yan97, Ye98a]. second-kind [WK96a]. Second-order [JPF95, GG96, GGS97, KM95, NS96, TPL95, Van95, WA96]. sedenions [Car97]. Seidel [KL99b, Zha96b]. self [JL97, WZZ98]. self-adaptive [WZZ98]. self-organizing [JL97]. selfadjoint [NR99b]. selftuning [KH95]. semi [AAA99, He96]. Semi-discrete [He96]. semi-infinite [AAA99]. semiconductor [ALZ99]. semidefinite [Han98]. semilinear [CY98, CZ99, CC99a]. sensitivities [Tuc97]. Sensitivity [AAES99, CO96, TH96]. sensor [Mil98]. Separable [KKH95]. Separable-denominator [KKH95]. separated [TSK98]. separation [HF95]. sequences [Fio95, Zha99b]. sequential [Cro95, BR99a]. series [Ami99, Boy95, Boy97a, CS97, CRS99, KWS95, Koł95, Kur99, Lam99, LSKH96, MMC98, RAL95, RAL97, Sri98, Waz97a, Waz98a, WC99, ZB97, dRS98]. series-parallel [Koł95, Kur99]. server [MS95b, Mou96]. service [MS95b]. servomechanism [Rya96]. Set [PZ97, Cro95, Ibr96, Lin96, Sha97]. set-valued [Ibr96]. sets [HK96, Zha97c]. several [Hua99]. sewer [BDJR98]. sexually [MAS95]. shading [Koz95]. shape [Koz95]. shapes [Jan95]. share [PZZy99]. shares [ZPY99]. Sharp [AP95]. Sharpe [MW98]. ship [BD95]. Shock [Dav98]. shooting [JBV98, Sel95]. Shortest [WLB97]. shuffle [Lam99]. sided [MNK99, Tag98]. sigmoid [WR96]. sign [LP97, RR99]. signal [Abu99a]. signals [Abu99b]. similarity [Sai96]. Simple [KH95, Abu99b, HH96a, JA99, Lui99, WZ98]. simplexes [JC97b]. Simplicial [JCC98]. Simulating [HY99a]. Simulation [Raj97, TH95, BR99a, Cos99, DGU96, FPGV95, Gre97, GMS96, HFR98, JS98b, QC98, SK98a, TPS98]. simulations [ALZ99, GSZS98, MAS95, MBS97, ZS98]. simulator [YCZ98]. simultaneous [Li98, ZS99]. Sinc [LB96, MLB99]. sine [BT98]. sine-Gordon [BT98]. Single [HF95, MS95b, Mou96, ZK98]. Single-phase [HF95]. Singular [RR98, Arg99a, DEV95, JR95, LO99, LUL97, NR99a, NR99b, RB95, VC98, WK96a, Waz99b, Wei98, Wei99a]. singularity [AC97]. singularly [KR95a, KM95, NR98]. sinusoid [Reg96]. size [WK98]. size-structured [WK98]. Skan [Asa97a, Asa98]. skein [GMGCL99]. skein-template [GMGCL99]. skill [ESL98]. slave [Ada98]. small [Cha96b, KA95]. smooth [Sol99b, SSY99]. smoothers [SO95]. Smoothing [CG96, Gao98, HH99, HSA97, YO98, Zha97a, Zha98b, Zha99a]. Sobolev [Gu99]. social [DML98]. software [DF96]. soils [Bar95]. solar [zHhT99]. solids [LK95]. Solution [MW98, OU99, Ado97b, AR98a, AG99, AK96, Asa97a, Béc98, BT98, DEV95, DF96, Fab96, HLXZ97, JK98, KKPL96, KW96, KW97, Khu98c, KSZ^{+}95, NR99b, NFJ94, Ohm98, Sai99, SKG97, Sha97, SA96, Sim97, Sim99a, Sim99b, SK96, SK98b, SS98b, Szi99, TSK98, Waz95a, Waz97a, WW99b, YST97, YK97a, ZZ98, ZL98a, ZMV99]. solution-adaptive [TSK98]. Solutions [SW96, TG97, Ado97a, AW99b, ASK98, BD96, Boy97b, CY98, CZ99, CC99a, DVD95, DPTT98, Gol96, Ham98a, HHZ97, JPF95, JN95, Kan97, Kar98, Kim96, Lu98, LS98b, MOP^{+}97, NJC99, OM99, PF99, PGHX99, QA98, Sai96, She96, Sim97, Waz98a, WA96, WA98, XYY96, XZW96, Yan99, YHXX95]. solve [Mor95, ZL98b]. Solving [CH99, FS98, LMSS97, BPZ96, Che97, Din98, GS99, JR95, JBV98, LP97, MBS98, NR98, RC99, SJ96, Sel95, SS98a, SBH96, WZZ98, WK96a, WK96b, WK98, Waz98b, Waz99c, Zha98c]. Some [AH99b, Bel95, BB96, CSY98, CRS99, GS97, HH96a, Le98, Lia97, Par95, Unb96, ZS99, dRS98, Abu97, DPTT98, EORA99, GB98, IKP97, KVT98, Koł95, Kur99, Per96, SS97, MSZ96]. Soriano [Di 98]. source [AA95, AA99b, ZZ98]. sources [AKK99]. southern [PN97b]. space [AA99c, Arg96a, Arg99d, Cic99, Guo99, Ibr96, SK95, Tuc97, dBL95]. space-based [Cic99]. spaces [Arg96b, Arg99e, Cha95b]. sparse [Bai98, Gol95]. spatial [BR99b, SK95]. Special [GSZS98, Ham98a, SS98a, SU95, Sug98, Udw95, UTL^{+}95, UHL96]. specialist [PZZy99]. specializing [XZL97]. species [SA96]. specified [Sas95a]. spectral [CyGK99, CC99b, DVD95, nHyG97]. spectral-finite [nHyG97]. spectrophotometry [SK95]. spectrum [Abu99b, HZL97]. speed [SP95]. SPH [HSA97]. sphere [Ano97i, Mus97a, Mus97b, Mus98]. sphere-packing [Mus97b]. spherical [AAAaZ99b]. spherically [AAAaZ99b]. spills [BR99b]. spline [AM98, Beh96a, Beh96b, Ona97, Wan95]. spline-on-spline [Beh96b]. splines [Beh95, HH99]. Splitting [IKP97, NLC95, RGL98, Wei98]. SQP [Sun98]. square [Ye98a]. squares [AB95a, Cha96a, SUJ98, SSY99, Whi97, Yan98, YL98a, Ye98b]. squeezed [BKS98]. St. [DH99b]. Stability [AR98b, Bil96, Bla96, GQS98, GF96, HH96b, HH97, LS95, Mor98, SSW99, Sol98, VLP95, Bel95, FL99, HH96a, ITN98a, JNF95, JNF96, Kim96, Kum95, LS99, LL96, LG99, PN97a, Sol99a, SBH96, Tuc95, Yan98]. stabilization [Lia98, Zha99a]. stabilizes [HSA97]. Stable [Lar97, BKS98, GGS97]. stage [Bai99b, Kha96, SSW99, ZA96, GR97]. stage-structured [ZA96]. stamps [Abd99]. standard [AD98, BB96, VC98, WR96, Yan98]. state [BKS98, Bar95, GP97, Kim96, Yam95, ZA96]. state-dependent [ZA96]. static [Sha97]. stationary [JCC98, Sug95]. statistical [WC99]. statistics [JY97]. steady [AH99b, Bar95, BJTW98, CL98a, GP97, Kim96, Raj97]. steady-state [Bar95, GP97]. Stefan [Asa97b, YS98a]. stencils [YO98]. step [Arg97b, Arg99c, Asa97b]. stepping [Moh99c]. Stieltjes [FT97, FT98]. stiff [MRJ95]. stochastic [Ami99, Cha95b, Cha96b, JCC98, MBCvdV95, MS97, OU99, Sug95, SS98b, CS96]. Stokes [Cha95a, Cha96b, CC99b, GP97, Ham98a, nHyG97, yHyG99, KW97, Khu98a, LL95, LY98, Ye98b, Ye99]. Stokes-like [GP97]. Stopping [DH99b]. strain [KD95a]. Strang [Wu97]. strange [DVD95]. strategy [SWL98]. stream [LY98, RPT97, Ye99]. strength [Ami97]. stress [AA95, AA99a, AAA99, Ami97]. stress-strength [Ami97]. Stresses [dBL95, AAAaZ99b, AAAaZ99a, ENS96]. string [KSZ^{+}95]. strongly [JC97a]. structural [Kum95]. structure [DML98, FHG96a, Jin95, LK95, SSW99, TyG99, Wan95]. structured [IKP97, KP98, TSK98, WK98, ZA96]. structures [Ano97i, CY99, Mus97a, Mus98, RRB95, Sol99b]. studies [AB95a]. study [Ano97i, AH99b, Cen97, DH99a, GF96, KA96b, LS98b, MFGW97, Mus97a, Mus98, RR98, SM95, Waz97b, Waz97c, Waz99b]. Sturm [ABW99, Zha95]. Subdomain [KR95b]. Subject [Ano95h, Ano96h, Ano97k]. subjects [ITN97]. submerged [KD95a, KD95b]. subsystem [Mur96]. successive [VR99f, VR99e]. successive-correction [VR99e]. successive-corrections [VR99f]. Sufficient [Arg98]. sum [Gri98, RR99]. summation [CS97, CSZ98, dRS98]. sums [WLTS99]. Sun [Tuc97]. superconductors [Cos99]. superposition [CG98]. support [HKFP97, RPT97, YB99]. supported [Wu97]. suppression [RL96]. surface [BJTW98, KUO95]. Surfaces [BK99, HT98, SUJ98]. survey [Kha96]. survival [ADS96]. sustainable [SAF^{+}97]. sustained [WZ95]. SWAGMAN [PMMA97]. Symbolic [GS95]. symmetric [BB96, Szi99, YO98, ST96]. symmetry [Yan99]. symmetry-like [Yan99]. Symplectic [TPGV97]. System [Har97, Abd99, Ado97d, AdO98, BY99, BDJR98, CT96, CT97, EB99, FW98, FBSS96, GSI96, GF96, Ham97, HKFP97, JC97a, JV98, Lia98, Lin96, LS98a, MZV96, MBS98, MR99, PZZy99, RR98, SA96, SST95, TETH98, Tuc95, VLP95, WW98, Wei98, XS95, XZW96, Yil96, IFU98, ZSL98]. systems [Ada99, AM98, ALZ99, AS99, Bai97, Bai99a, Bai99b, Bay97, Béc98, Bil96, BT97, Che95, Che96b, Chu98, FPGV95, Fri97, FHG96b, GS99, GS95, GKT96, GL99, Har95, Har97, HHT98, He96, HH96b, HH96a, HH97, Jia97, Jin95, JS98b, JNF95, JPF95, JNF96, KM95, Koł95, Kum95, Kur99, Lar97, LFS95, LR95, Lei95, LS99, LL96, MMMK96, MD98, Mor98, Mor95, MBS98, NFJ94, OMN95, RAL97, RL96, RRB95, Rya96, Sha97, Sol98, Sol99a, Sol99b, Son96, SK98a, Sug95, TH95, UK95, Unb96, Wan98, WK95, Wei99a, Yan98, YL98a, MFGW97].

table [PMMA97]. tailed [WLB97]. taking [LC97]. target [CO96, Cic99]. Targeting [HYL97]. Taylor [HY99b, Waz98a]. TDM [WW99a, WW99b]. technique [CN97, CL98a, HZL97, JC97a, Khu98b, LM95, RBPK96, TSK98, WK96a, Waz98b, WC99, ZS98, Zha98b, Zha99a]. techniques [Bay98b, Bay98a, BY99, DH99a, RGL98, Ram99d, Ram99b, RR98, VR99f, VR99b, VR99c, VR99d, VR99e, Zha97b, Zha98a]. temperature [MBG95]. template [GMGCL99]. tendency [ZPY99]. tensor [Jin95]. terms [FL99, Waz97a, ZZ98]. tests [BB96]. tether [Jan95]. tethered [SST95, VLP95]. Thailand [PN97b]. their [ABT98, Arg96a, Arg97a, Arg99d, CS97, CSY98, Kim96, SAL96, Yür99]. theorem [Ibr96, LS97, Zha95]. theorems [SS97]. theoretic [EM98]. theory [BR99a, Cen97, CY99, FM96, KW96, Khu98c, LS99, LK95, NR99a, RR98]. therapies [HD96]. therapy [DGU96]. Thermal [AA95, AAAaZ99b, AAAaZ99a, CK95]. thermistor [KBÖ99]. thermoelastic [AA96, AA99b]. thin [KH96]. thin-film [KH96]. third [Pei96]. third-order [Pei96]. Thomas [CL98b, Waz99c]. three [Dua99a, EB99, GL99, nHyG97, yHyG99, SA96, ST96, Tuc95, ZK98, KD95b]. three-body [Tuc95]. three-degree-of-freedom [EB99]. three-dimensional [nHyG97, yHyG99, ST96, KD95b]. three-species [SA96]. Time [Dua99b, LZ96, Moh99c, ABW99, AR98b, AKK99, Ami99, Asa97b, CT96, Fri97, Kum95, LSKH96, LS99, LL96, Lia97, LSY98, Lu98, LLX96, LUL97, MMC98, Ram98c, Sug95, SO95, WZZ98, WC99, ZA96]. time-delay [LL96]. time-invariant [LS99]. time-linearized [Ram98c]. Time-periodic [Dua99b, LZ96]. time-varying [LSY98]. times [RB95]. Toeplitz [GS99, Jin95]. Tomography [PZ99]. tool [TH96]. tools [Van97]. topological [AW99a]. tori [Tak98]. toroidal [Khu98a]. total [JMS98, SUJ98]. tracking [CO96, Cic99, KK95, SO95]. traction [Lee98]. trade [Van97]. trade-off [Van97]. trading [MPD97]. traffic [Mil95, MFGW97]. trajectory [SMY95]. transfer [KL99a, MA95, ŞKB95]. transform [CH99, Hai97, Kan99, Tag98]. transformation [CH96, CL98a, FH95, FHG96a, JC97a, MD98]. transforms [Gzy95, Gzy97, Jay96, KVT98, MNK99, Yür99, Zen99]. Transient [AAAaZ99b, AAAaZ99a, AA96, SZ97]. transistors [SEG96]. transition [Kub97, NS97, VD97]. transmission [Alw97]. transmitted [MAS95]. transport [AKK99, HY98, HY99b, HM99, HY99c, Hos99, HY99a, LS98a, LZ96, WW98]. Transportation [MFGW97]. transversal [VLP95]. transversality [AW99a]. transversely [AA95, AA96]. Trapezoidal [Yeh97]. trapping [Cos99]. Traveling [Boy97b]. treatment [BR99b, PF99]. trees [ZSL98]. trends [Unb96]. triangle [Beh96a, Kor97]. triangular [CL98b, Die96, KKL99, Tuc97]. triangulated [JC97b]. triangulations [Wan95]. tridiagonal [MS97]. trigonometric [Jia99]. trimmed [HT98]. Trust [dSYS97, SSY99]. tubes [AAAaZ99a, ST96]. turbulent [Raj97]. turning [NR98]. twin [NR98]. Two [LL95, MR99, WK96b, Ye99, YK97a, YK97b, Zha98b, Arg97b, Arg99c, Bai99b, CH99, CS99, FL99, Gra97, JY97, KBS96, KKH95, Kha96, LY98, MNK99, Mou96, NS97, Ohm98, OU99, SW96, SSW99, Tag98, VR99f, Waz99b, WW99a, WLB97, ZMV99, GR97]. two-dimensional [CH99, KKH95, NS97, OU99]. two-fluid [Ohm98]. Two-grid [Zha98b]. Two-level [LL95, YK97b, LY98]. Two-point [MR99, KBS96, VR99f]. two-sided [MNK99, Tag98]. two-stage [Bai99b, Kha96, SSW99, GR97]. two-step [Arg97b, Arg99c]. two-tailed [WLB97]. type [Ado97c, AZK96, BS97, Cen97, EORA99, GL95, HHT98, He96, Pei96, Sri95, WK96a, Cos99]. Type-II [Cos99]. types [MSZ96]. typological [Ano97i, Mus97a, Mus98].

ultra [TPS98]. ultra-relativistic [TPS98]. unbounded [AA99b, Guo99, Jia97, JPF95]. uncertain [Che95, Lei95, RRB95, TH95]. Uncertainty [EFS95, Fri97, Ham98b, IL95, MB95, SL96]. Unconditionally [GGS97]. unconstrained [Lia97]. underwater [BD95, KP96]. unified [Arg96a, Arg99d, Tag99, YL98a]. Uniform [FL99]. uniqueness [Koz95, MR99, Sev99b]. unitary [Boy97a]. units [LSKH96]. unity [Boy97a]. Universal [Chu96, Rya96]. unknowns [GP97]. unsteady [BJTW98, nHyG97, yHyG99, JFW96, Raj97]. unstirred [BR98]. unstructured [LS98a]. updating [AB95b]. Upper [BDS96]. Upstream [Ram98d]. uptake [OA97]. upwind [Váz98]. Urn [MAS95]. Use [HD96, WC99, AFM97, Beh96b, ONOT98]. used [BR99b, TJ95]. useful [Par95]. uses [PMMA97]. using [BY99, Cen97, CW97, CLC97, Cic99, DF96, ESL98, FS98, GS95, GMGCL99, GP97, GKT96, HT98, HZ96, JC97a, Jay96, JY97, JL97, KL99b, LS98a, MBCvdV95, Ohm98, Ona97, PF99, Raj97, Sel95, SIA96, WK95, ZPY99]. utility [DH99b, MS99]. utilization [SS98a].

V [VR99e]. Validated [NJC99]. Value [BT97, AW99b, AG99, AY98, BKM95, Beh96b, CN97, Cha96a, Guo99, Gup98, HG99, Ibr96, Jia97, KR95a, KBS96, MR99, NJC99, Pei95, Pei96, QA98, RGL97b, Sas95b, Sev99b, TS98, VS96, VR99f, Wan99a, Wan99b, WCG97, WA99, YK97a]. valued [Ibr96, MvL98]. values [Gri98, Waz97c]. Vandermonde [GS97]. variable [Asa97b, Dég96, FBSS96, SW96, TG97]. variable-coefficient [TG97]. variables [AP95, Ami97, HI97, WP97]. variant [AY98]. variational [PZ97, PK98, ZZ98, Zha97c]. variety [KBÖ99]. various [ESL98, KSZ^{+}95]. Varying [SBH96, LSY98]. vector [XZL97]. vectors [LUL97]. vehicle [BD95, CO96, KP96, SWL98]. velocity [BNN95, Raj97]. vertex [Kor97]. very [KWS95]. VI [VR99f]. via [ABT98, BK99, CS96, HZ95, MF99, MW98, Mil98, Mus97b, Pap96, RAL95, RAL97, Sol98, TG97, XYY96]. Vibration [MD98, Ado96d, LR95, MB95, RL96]. vibrations [Bar95]. violin [Mit97, Mit99]. virtual [Zha98c]. viscous [HF95]. visualization [BK99, Kar98]. VML [Zha98c]. Volterra [AdO98, AR98b, Bla96, LL96, SA96, Sol99a, WK96a, Waz99a]. Volterra-type [WK96a]. Volume [Ano95o, Ano95i, Ano95k, Ano95m, Ano95j, Ano95l, Ano95n, Ano96o, Ano96i, Ano96k, Ano96m, Ano96j, Ano96l, Ano96n, Ano97u, Ano97l, Ano97m, Ano97p, Ano97r, Ano97n, Ano97o, Ano97q, Ano97s, Ano97t, Ano98o, Ano98p, Ano98d, Ano98e, Ano98f, Ano98a, Ano98b, Ano98c, Ano99c, Ano99a, Ano99b, LM95, LS98a, STD98]. volumes [Ano95h, Ano96h, Ano97k, Ano99d]. Voronoı [HT98]. vortex [Cos99]. Vries [RC99]. vs [AB95a].

W [Chu97]. water [BDS96, DH99a, HYL97, Khu98c, OA97, PMMA97, RPT97, YBR97]. Wave [ENS96, Abu99a, Boy97b, KD95a, KD95b, Reg96, Waz98b]. waveform [Son96, Zha96a]. wavelet [XS95, ZL98c]. waves [AA99a, AAA99, AA99c, Ahm99, Boy97b, DH99a, Khu98c]. weak [Abu97, JK98]. weakly [Bai99a, WK96a]. weight [GBS99]. weighted [CLC97, RBPK96, YL98a]. weighting [CL99]. white [WL97]. widely [HZL97]. Wiener [NLC95]. Wigley [BJTW98]. windowing [Zha96a].

Zakharov [Ado97b]. zero [Sor96]. zeros [Sor97b, Sor99, XZL97, ZS99].

References

Abd-Alla:1995:TST

[AA95] A. M. Abd-Alla. Thermal stress in a transversely isotropic circular cylinder due to an instantaneous heat source. Applied Mathematics and Computation, 68(2–3):113–124, March 15, 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
009630039400085I.

Abd-Alla:1996:TRI

[AA96] A. M. Abd-Alla. Transient response of inhomogeneous transversely isotropic thermoelastic cylinder to a dynamic input. Applied Mathematics and Computation, 74(1):1–13, January 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0096300394002460.

Abd-Alla:1999:EIS

[AA99a] A. M. Abd-Alla. The effect of initial stress and orthotropy on the propagation waves in a hollow cylinder. Applied Mathematics and Computation, 106(2–3):237–244, December 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300398101261.

Abd-Alla:1999:GTI

[AA99b] A. M. Abd-Alla. On a generalized thermoelastic interaction in an unbounded body due to a line heat source. Applied Mathematics and Computation, 100(2–3):285–295, May 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.

com/science/article/pii/
S0096300398000320.

Abd-Alla:1999:PRW

[AA99c]  A. M. Abd-Alla.  Propa-

gation of Rayleigh waves in an elastic half-space of orthotropic material. Applied Mathematics and Computation, 99(1):61–69, March 1, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300397101709.

Abd-Alla:1999:PLW

[AAA99] A. M. Abd-Alla and S. M. Ahmed. Propagation of Love waves in a non-homogeneous orthotropic elastic layer under initial stress overlying semiinfinite medium. Applied Mathematics and Computation, 106(2–3):265–275, December 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300398101285.

Abd-All:1999:TTSb

[AAAaZ99a] A. M. Abd-All, A. N. Abdalla, and N. A. Zeidan. Transient thermal stresses in a rotation non-homogeneous cylindrically orthotropic composite tubes. Applied Mathematics and Computation, 105(2–3): 253–269, November/December 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300398101042.

Abd-All:1999:TTSa

[AAAaZ99b] A. M. Abd-All, A. N. Abdalla, and N. A. Zeidan. Transient thermal stresses in a spherically orthotropic elastic medium with spherical cavity. Applied Mathematics and Computation, 105(2–3): 231–252, November/December 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300398100930.

Alwasel:1998:EPB

[AAAN98] I. Alwasel, A. Alzaid, and H. Al-Nachawati. Estimating the parameters of the binomial autoregressive process of order one. Applied Mathematics and Computation, 95 (2–3):193–204, September 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300397101023.

Abdel-Aziz:1999:SAL

[AAES99] Mohammedi R. Abdel-Aziz and Salah M. El-Sayed. Sensitivity analysis of the largest dependent eigenvalue functions of eigensystems. Applied Mathematics and Computation, 100(2–3):103–110, May 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300398000319.

Angleton:1995:LSR

[AB95a] George M. Angleton and Charles D. Bonham. Least squares regression vs. geometric mean regression for ecotoxicology studies. Applied Mathematics and Computation, 72(1):21–32, September 15, 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        009630039400161V.

Azaiez:1995:PAC

[AB95b] M. Naceur Azaiez and Vicki M. Bier. Perfect aggregation for a class of general reliability models with Bayesian updating. Applied Mathematics and Computation, 73(2–3):281–302, December 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0096300395000682.

Abdou:1999:IEC

[Abd99]  M. A. Abdou.  Integral

equation and contact problem for a system of impressing stamps. Applied Mathematics and Computation, 106(2–3):141–148, December 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S009630039810108X.

Abaffy:1998:PMA

[ABT98] Jozsef Abaffy, Marida Bertocchi, and Anna Torriero. Perturbations of M-matrices via ABS methods and their applications to input-output analysis. Applied Mathematics and Computation, 94 (2–3):145–170, August 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300397100856.

Abuelmaatti:1997:HIP

[Abu97] Muhammad Taher Abuelma’atti.

Harmonic and intermodulation performance of some abrupt and weak nonlinearities. Applied Mathematics and Computation, 83(1):87–96, April 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300396000495.

Abuelmaatti:1999:NDF

[Abu99a] Muhammad Taher Abuelma’atti.

Nonlinear distortion of an FM signal by half-wave linear rectification. Applied Mathematics and Computation, 99(1):47–59, March 1, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300397101795.

Abuelmaatti:1999:SAC

[Abu99b] Muhammad Taher Abuelma’atti. A simple algorithm for computing the Fourier spectrum of experimentally obtained signals. Applied Mathematics and Computation, 98(2–3):229– 239, February 1999. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic). URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/S0096300397101606.

Agarwal:1999:SLE

[ABW99] Ravi P. Agarwal, Martin Bohner, and Patricia J. Y. Wong. Sturm–Liouville eigenvalue problems on time scales. Applied Mathematics and Computation, 99(2–3):153–166, March 15, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300398000046.

Anghel:1997:NMM

[AC97] V. N. P. Anghel and W.- H. Choe. Numerical methods for magnetized plasma equations with boundary singularity. Applied Mathematics and Computation, 85 (2–3):149–164, September 1, 1997. CODEN AMHCBQ.

ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300396001324.

Agouzal:1998:EMA

[AD98] A. Agouzal and N. Debit.

Equilibrium method to approximate elliptic problems with non-standard boundary conditions. Applied Mathematics and Computation, 95 (2–3):165–171, September 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300397100789.

Adamatzky:1997:HCA

[Ada97] Andrew Adamatzky. How cellular automaton plays Minesweeper. Applied Mathematics and Computation, 85(2–3):127–137, September 1, 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
sciencedirect.com/science/
article/pii/S0096300396001178.

Adamatzky:1998:MSA

[Ada98] Andrew Adamatzky. Masterslave algebra: On the binary compositions of distributed beliefs. Applied Mathematics and Computation, 95 (2–3):173–180, September 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300397100996.

Adamatzky:1999:NBC

[Ada99] Andrew Adamatzky. Nonconstructible blocks in 1D cellular automata: minimal generators and natural systems. Applied Mathematics and Computation, 99(1):77–91, March 1, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300397101771.

Adomian:1996:BKM

[Ado96a] G. Adomian. The Burridge– Knopoff model. Applied Mathematics and Computation, 77(2–3):131–132, July 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300395001956.

Adomian:1996:CME

[Ado96b] G. Adomian. Coupled Maxwell equations for electromagnetic scattering. Applied Mathematics and Computation, 77(2– 3):133–135, July 1996. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic). URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/S0096300395001964.

Adomian:1996:KPE

[Ado96c] G. Adomian. The Kadomtsev– Petviashvili equation. Applied Mathematics and Computation, 76(1):95–97, April 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0096300395001867.

Adomian:1996:NRV

[Ado96d]  G. Adomian.  Nonlinear

random vibration. Applied Mathematics and Computation, 77(2–3):109–112, July 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300395001921.

Adomian:1997:ESN

[Ado97a] G. Adomian. Explicit solutions of nonlinear partial differential equations. Applied Mathematics and Computation, 88 (2–3):117–126, December 30, 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300396001415.

Adomian:1997:NPS

[Ado97b] G. Adomian. Non-perturbative solution of the Klein–Gordon– Zakharov equation. Applied Mathematics and Computation, 81(1):89–92, January 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0096300395003118.

Adomian:1997:KTE

[Ado97c] G. Adomian. On KdV type equations. Applied Mathematics and Computation, 88 (2–3):131–135, December 30, 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300396002834.

Adomian:1997:DRD

[Ado97d] G. Adomian. On the dynamics of a reaction-diffusion system. Applied Mathematics and Computation, 81(1):93–95, January 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0096300395003126.

Adomian:1997:OPR

[Ado97e] G. Adomian. Optical propagation in random media. Applied Mathematics and Computation, 88(2–3):127–129, December 30, 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300396002822.

Ahmad:1998:ENP

[AdO98] Shair Ahmad and Francisco Montes de Oca. Extinction in nonautonomous T -periodic competitive Lotka–Volterra system. Applied Mathematics and Computation, 90(2–3):155–166, March 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300397003962.

Ananda:1996:ABE

[ADS96] Malwane M. Ananda, Rohan J.

Dalpatadu, and Ashok K. Singh. Adaptive Bayes estimators for parameters of the Gompertz survival model. Applied Mathematics and Computation, 75(2–3):167–177, March 15, 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300396900594.

Antoine:1997:MCL

[AFM97] Jacques Antoine, Günther Fischer, and Marek Makowski. Multiple criteria land use analysis. Applied Mathematics and Computation, 83(2–3):195–215, May 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300396001907.

Al-Gahtani:1999:IBS

[AG99] H. J. Al-Gahtani. Integralbased solution for a class of second order boundary value problems. Applied Mathematics and Computation, 98(1):43–48, January 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300397101503.

Awartani:1999:NRG

[AH99a] Marwan Awartani and M. H. Hamdan. Non-reactive gasparticulate models of flow through porous media. Applied Mathematics and Computation, 100(1):93–102, April 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300398000137.

Awartani:1999:SAG

[AH99b] Marwan Awartani and M. H. Hamdan. Some admissible geometries in the study of steady plane flow of a dusty fluid through porous media. Applied Mathematics and Computation, 100(1):85–92, April 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300398000125.

Ahmed:1999:IGP

[Ahm99] S. M. Ahmed. Influence of gravity on the propagation of waves in granular medium. Applied Mathematics and Computation, 101(2–3):269–280, June 15, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300398100061.

Ali:1996:AFC

[AK96] Ismail Ali and Shyam Kalla. An application of fractional calculus to the solution of a general class of differintegral equations. Applied Mathematics and Computation, 77(2–3):137–152, July 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300395001972.

Ali:1999:TDM

[AKK99] I. Ali, S. L. Kalla, and H. G.

Khajah. A time dependent model for the transport of heavy pollutants from groundlevel aerial sources. Applied Mathematics and Computation, 105(1):91–99, October 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300398100917.

Alwash:1997:LCD

[Alw97] M. A. M. Alwash. Limit cycles in a disease transmission model. Applied Mathematics and Computation, 86(1):85–92, September 15, 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300396001749.

Allegretto:1999:NAS

[ALZ99] W. Allegretto, Y. Lin, and A. Zhou. Numerical analysis for systems with memory arising from semiconductor simulations. Applied Mathematics and Computation, 105(2–3): 101–119, November/December 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300398100814.

Agwu:1998:OCD

[AM98] Nwojo N. Agwu and Clyde F. Martin. Optimal control of dynamic systems: Application to spline approximations. Applied Mathematics and Computation, 97(2–3):99–138, December 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300397101011.

Aminzadeh:1997:ERE

[Ami97] M. S. Aminzadeh. Estimation of reliability for exponential stress-strength models with explanatory variables. Applied Mathematics and Computation, 84(2–3):269–274, July 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300396000902.

Aminzadeh:1999:ECS

[Ami99] Mostafa S. Aminzadeh. Estimation of P (Z < Y) for correlated stochastic time series models. Applied Mathematics and Computation, 104(2–3):179–189, September 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300398100723.

Anonymous:1995:EBa

[Ano95a] Anonymous. Editorial Board.

Applied Mathematics and Computation, 67(1–3):ii–iii, January/February 1995. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
sciencedirect.com/science/
article/pii/0096300395900055.

Anonymous:1995:EBb

[Ano95b] Anonymous. Editorial Board.

Applied Mathematics and Computation, 68(1):ii–iii, March 1, 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0096300395900543.

Anonymous:1995:EBc

[Ano95c] Anonymous. Editorial Board. Applied Mathematics and Computation, 69(1):ii–iii, April 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0096300395900098.

Anonymous:1995:EBd

[Ano95d] Anonymous. Editorial Board. Applied Mathematics and Computation, 70(1):ii–iii, June 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0096300395900438.

Anonymous:1995:EBe

[Ano95e] Anonymous. Editorial Board. Applied Mathematics and Computation, 71(1):ii–iii, August 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0096300395900535.

Anonymous:1995:EBf

[Ano95f] Anonymous. Editorial Board. Applied Mathematics and Computation, 72(1):ii–iii, September 15, 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0096300395900136.

Anonymous:1995:EBg

[Ano95g] Anonymous. Editorial Board.

Applied Mathematics and Computation, 73(1):??, November 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0096300395900217.

Anonymous:1995:SIV

[Ano95h] Anonymous. Subject index to volumes 67–73, 1995. Applied Mathematics and Computation, 73(2–3):303–309, December 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0096300395900454.

Anonymous:1995:Va

[Ano95i] Anonymous. Volume 68. Applied Mathematics and Computation, 68(2–3):??, March 15, 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0096300395900551.

Anonymous:1995:Vd

[Ano95j] Anonymous. Volume 69. Applied Mathematics and Computation, 69(2–3):i–ii, May 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0096300395900519.

Anonymous:1995:Vb

[Ano95k] Anonymous. Volume 70. Applied Mathematics and Computation, 70(2–3):i–ii, July 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0096300395900373.

Anonymous:1995:Ve

[Ano95l] Anonymous. Volume 71. Applied Mathematics and Computation, 71(2–3):??, September 1, 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        009630039590011X.

Anonymous:1995:Vc

[Ano95m] Anonymous. Volume 72. Applied Mathematics and Computation, 72(2–3):i–ii, October 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0096300395900039.

Anonymous:1995:Vf

[Ano95n] Anonymous. Volume 73. Applied Mathematics and Computation, 73(2–3):i–ii, December 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0096300395900470.

Anonymous:1995:VCX

[Ano95o]  Anonymous.  Volume Con-

tents. Applied Mathematics and Computation, 67 (1–3):??, January/February 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0096300395900071.

Anonymous:1996:EBa

[Ano96a] Anonymous. Editorial Board.

Applied Mathematics and Computation, 74(1):ii–iii, January 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300396900478.

Anonymous:1996:EBb

[Ano96b] Anonymous. Editorial Board.

Applied Mathematics and Computation, 75(1):??, March 1, 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300396900375.

Anonymous:1996:EBc

[Ano96c] Anonymous. Editorial Board. Applied Mathematics and Computation, 76(1):ii–iii, April 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300316900018.

Anonymous:1996:EBd

[Ano96d] Anonymous. Editorial Board. Applied Mathematics and Computation, 77(1):ii–iii, June 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0096300396900296.

Anonymous:1996:EBe

[Ano96e] Anonymous. Editorial Board. Applied Mathematics and Computation, 78(1):ii–iii, August 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300396900399.

Anonymous:1996:EBf

[Ano96f] Anonymous. Editorial Board. Applied Mathematics and Computation, 79(1):ii–iii, September 15, 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300396900430.

Anonymous:1996:EBg

[Ano96g] Anonymous. Editorial Board.

Applied Mathematics and Computation, 80(1):ii–iii, November 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0096300396900417.

Anonymous:1996:SIV

[Ano96h] Anonymous. Subject index to volumes 74–80, 1996. Applied Mathematics and Computation, 80(2–3):301–306, December 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0096300396900193.

Anonymous:1996:Vb

[Ano96i] Anonymous. Volume 74. Applied Mathematics and Computation, 74(2–3):??, February 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
009630039690003X.

Anonymous:1996:Ve

[Ano96j] Anonymous. Volume 75. Applied Mathematics and Computation, 75(2–3):i–ii, March 15, 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        0096300396900053.

Anonymous:1996:Vc

[Ano96k] Anonymous. Volume 76. Applied Mathematics and Computation, 76(2–3):i–ii, May 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300396900454.

Anonymous:1996:Vf

[Ano96l] Anonymous. Volume 77. Applied Mathematics and Computation, 77(2–3):??, July 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0096300396900119.

Anonymous:1996:Vd

[Ano96m] Anonymous. Volume 78. Applied Mathematics and Computation, 78(2–3):??, September 1, 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0096300396900090.

Anonymous:1996:Vg

[Ano96n] Anonymous. Volume 79. Applied Mathematics and Computation, 79(2–3):i–ii, October 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300396900351.

Anonymous:1996:Va

[Ano96o] Anonymous. Volume 80. Applied Mathematics and Computation, 80(2–3):i–ii, December 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0096300396900211.

Anonymous:1997:EBa

[Ano97a] Anonymous. Editorial Board.

Applied Mathematics and Computation, 81(1):ii–iii, January 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300397900230.

Anonymous:1997:EBb

[Ano97b] Anonymous. Editorial Board.

Applied Mathematics and Computation, 82(1):ii–iii, March 1, 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300397900795.

Anonymous:1997:EBc

[Ano97c] Anonymous. Editorial Board.

Applied Mathematics and Computation, 83(1):ii–iii, April 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300397900333.

Anonymous:1997:EBd

[Ano97d] Anonymous. Editorial Board. Applied Mathematics and Computation, 84(1):??, June 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/S0096300397900151.

Anonymous:1997:EBe

[Ano97e] Anonymous. Editorial Board. Applied Mathematics and Computation, 85(1):ii–iii, August 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0096300397900412.

Anonymous:1997:EBf

[Ano97f] Anonymous. Editorial Board. Applied Mathematics and Computation, 86(1):??, September 15, 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0096300397900291.

Anonymous:1997:EBg

[Ano97g] Anonymous. Editorial Board. Applied Mathematics and Computation, 87(1):??, November 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300397900217.

Anonymous:1997:EBh

[Ano97h] Anonymous. Editorial Board.

Applied Mathematics and Computation, 88(1):??, December 15, 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300397900758.

Anonymous:1997:EDF

[Ano97i] Anonymous. Erratum: “The dimensional family approach in (hyper)sphere packing: a typological study for new patterns, structures, and interdimensional functions”. Applied Mathematics and Computation, 88(1):??, December 15, 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300397004025.  See

[Mus97a, Mus98].

Anonymous:1997:PN

[Ano97j] Anonymous. Publisher’s note.

Applied Mathematics and Computation, 88(1):vii, December 15, 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300397900771.

Anonymous:1997:SIV

[Ano97k] Anonymous. Subject index to volumes 81–87, 1997. Applied Mathematics and Computation, 87(2–3):313–318, December 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300397003457.

Anonymous:1997:Va

[Ano97l] Anonymous. Volume 81. Applied Mathematics and Computation, 81(2–3):i–ii, February 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0096300397900035.

Anonymous:1997:Vb

[Ano97m] Anonymous. Volume 82. Applied Mathematics and Computation, 82(2–3):i–ii, March 15, 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0096300397900096.

Anonymous:1997:Ve

[Ano97n] Anonymous. Volume 82. Applied Mathematics and Computation, 82(2–3):i–ii, March 15, 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.

com/science/article/pii/
0096300397900096.

Anonymous:1997:Vf

[Ano97o] Anonymous. Volume 83. Applied Mathematics and Computation, 83(2–3):??, May 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0096300397900072.

Anonymous:1997:Vc

[Ano97p] Anonymous. Volume 84. Applied Mathematics and Computation, 84(2–3):??, July 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0096300397900114.

Anonymous:1997:Vg

[Ano97q] Anonymous. Volume 85. Applied Mathematics and Computation, 85(2–3):i–ii, September 1, 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0096300397900059.

Anonymous:1997:Vd

[Ano97r] Anonymous. Volume 86. Applied Mathematics and Computation, 86(2–3):??, October 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300397817894.

Anonymous:1997:Vh

[Ano97s] Anonymous. Volume 87. Applied Mathematics and Computation, 88(2–3):i–ii, December 30, 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0096300397900011.

Anonymous:1997:Vi

[Ano97t] Anonymous. Volume 88. Applied Mathematics and Computation, 88(2–3):i–ii, December 30, 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300397892922.

Anonymous:1997:VCX

[Ano97u] Anonymous. Volume contents. Applied Mathematics and Computation, 87(2–3):??, December 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300397817900.

Anonymous:1998:AIVd

[Ano98a] Anonymous. Author index to volume 91 (1998). Applied Mathematics and Computation, 91(2–3):297–298, May 1998. CODEN AMHCBQ.

ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300398100218.

Anonymous:1998:AIVe

[Ano98b] Anonymous. Author index to volume 92 (1998). Applied Mathematics and Computation, 92(2–3):299–300, June 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300398100322.

Anonymous:1998:AIVf

[Ano98c] Anonymous. Author index to volume 93 (1998). Applied Mathematics and Computation, 93(2–3):299–300, July 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300398100334.

Anonymous:1998:AIVa

[Ano98d] Anonymous. Author index to volume 94 (1998). Applied Mathematics and Computation, 94(2–3):293–294, August 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300398100437.

Anonymous:1998:AIVb

[Ano98e] Anonymous. Author index to volume 96 (1998). Applied Mathematics and Computation, 96(2–3):295–296, November 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300398101017.

Anonymous:1998:AIVc

[Ano98f] Anonymous. Author index to volume 97 (1998). Applied Mathematics and Computation, 97(2–3):297–298, December 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300398101029.

Anonymous:1998:EBa

[Ano98g] Anonymous. Editorial Board. Applied Mathematics and Computation, 89(1–3):ii–iii, January/February 1998. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic). URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300398900054.

Anonymous:1998:EBb

[Ano98h] Anonymous. Editorial Board. Applied Mathematics and Computation, 90(1):ii–iii, March 1, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300398900091.

Anonymous:1998:EBc

[Ano98i] Anonymous. Editorial Board.

Applied Mathematics and Computation, 91(1):??, April 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/S0096300398900078.

Anonymous:1998:EBd

[Ano98j] Anonymous. Editorial Board.

Applied Mathematics and Computation, 92(1):??, June 1, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S009630039890011X.

Anonymous:1998:EBe

[Ano98k] Anonymous. Editorial Board.

Applied Mathematics and Computation, 93(1):??, July 1, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0096300398900133.

Anonymous:1998:EBf

[Ano98l] Anonymous. Editorial Board.

Applied Mathematics and Computation, 95(1):??, September 1, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0096300398900017.

Anonymous:1998:EBg

[Ano98m] Anonymous. Editorial Board. Applied Mathematics and Computation, 96(1):??, October 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0096300398900170.

Anonymous:1998:EBh

[Ano98n] Anonymous. Editorial Board. Applied Mathematics and Computation, 97(1):??, December 1, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300398900157.

Anonymous:1998:Va

[Ano98o] Anonymous. Volume 89. Applied Mathematics and Computation, 89(1–3):??, January/February 1998. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
sciencedirect.com/science/
article/pii/S0096300398900194.

Anonymous:1998:Vb

[Ano98p] Anonymous. Volume 90. Applied Mathematics and Computation, 90(2–3):??, March 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300397802019.

Anonymous:1999:AIVb

[Ano99a] Anonymous. Author index to volume 100 (1999). Applied Mathematics and Computation, 100(2–3):297–298, May 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300399000144.

Anonymous:1999:AIVd

[Ano99b] Anonymous. Author index to volume 98 (1999). Applied Mathematics and Computation, 98(2–3):301–302, February 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300398101479.

Anonymous:1999:AIVa

[Ano99c] Anonymous. Author index to volume 99 (1999). Applied Mathematics and Computation, 99(2–3):295–296, March 15, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300398101480.

Anonymous:1999:AIVc

[Ano99d] Anonymous. Author index to volumes 88–100. Applied Mathematics and Computation, 100(2–3):299–313, May 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300399000156.

Anonymous:1999:Ca

[Ano99e] Anonymous. Conference. Applied Mathematics and Computation, 105(2–3):??, November/December 1999. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
sciencedirect.com/science/
article/pii/S0096300399001873.

Anonymous:1999:Cb

[Ano99f] Anonymous. Conference. Applied Mathematics and Computation, 106(2–3):??, December 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300399002210.

Anonymous:1999:Ib

[Ano99g]  Anonymous.  Index.  Ap-

plied Mathematics and Computation, 103(2–3):287, August 15, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300399000971.

Anonymous:1999:Ic

[Ano99h] Anonymous. Index. Applied Mathematics and Computation, 104(2–3):299–300, September 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S009630039900185X.

Anonymous:1999:Ia

[Ano99i] Anonymous. Index. Applied Mathematics and Computation, 105(2–3):307–308, November/December 1999. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
sciencedirect.com/science/
article/pii/S0096300399001861.

Anonymous:1999:Id

[Ano99j] Anonymous. Index. Applied Mathematics and Computation, 106(2–3):305–306, December 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300399002209.

Agarwal:1995:SDI

[AP95] Ravi P. Agarwal and Peter Y. H. Pang. Sharp discrete inequalities in n independent variables. Applied Mathematics and Computation, 72(2–3):97–112, October 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
009630039400130V.

Ahmad:1998:NSG

[AR98a] Falih Ahmad and Mohsen Razzaghi. A numerical solution to the Gel’fand–Levitan– Marchenko equation. Applied Mathematics and Computation, 89(1–3):31–39, January/February 1998. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
sciencedirect.com/science/
article/pii/S0096300397816463.

Ahmad:1998:SVD

[AR98b] Shair Ahmad and M. Rama Mohana Rao. Stability of Volterra diffusion equations with time delays. Applied Mathematics and Computation, 90(2–3):143–154, March 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300397003950.

Argyros:1996:RNMa

[Arg96a] Ioannis K. Argyros. Results on Newton methods: Part I. A unified approach for constructing perturbed Newtonlike methods in Banach space and their applications. Applied Mathematics and Computation, 74(2–3):119–141, February 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0096300395000909.

Argyros:1996:RNMb

[Arg96b] Ioannis K. Argyros. Results on Newton methods: Part II. Perturbed Newton-like methods in generalized Banach spaces. Applied Mathematics and Computation, 74(2–3):143–159, February 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0096300395001182.

Argyros:1997:MIP

[Arg97a] Ioannis K. Argyros. A mesh independence principle for inexact Newton-like methods and their discretizations under generalized Lipschitz conditions. Applied Mathematics and Computation, 87(1):15–48, November 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300396002196.

Argyros:1997:CTS

[Arg97b]  Ioannis K. Argyros.  On

the convergence of two-step methods generated by pointto-point operators. Applied Mathematics and Computation, 82(1):85–96, March 1, 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        S0096300396000252.

Argyros:1998:SCC

[Arg98] Ioannis K. Argyros. Sufficient conditions for constructing methods faster than Newton’s. Applied Mathematics and Computation, 93(2–3):169–181, July 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300397101072.

Argyros:1999:CRI

[Arg99a] Ioannis K. Argyros. Convergence rates for inexact Newtonlike methods at singular points and applications. Applied Mathematics and Computation, 102(2–3):185–201, July 15, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300398100152.

Argyros:1999:NMU

[Arg99b] Ioannis K. Argyros. On Newton’s method under mild differentiability conditions and applications. Applied Mathematics and Computation, 102(2–3):177–183, July 15, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300398100140.

Argyros:1999:CTS

[Arg99c]  Ioannis K. Argyros.  On

the convergence of two-step methods generated by pointto-point operators. Applied Mathematics and Computation, 102(2–3):165–176, July 15, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300398100164.

Argyros:1999:RNMa

[Arg99d] Ioannis K. Argyros. Results on Newton methods. Part 1: a unified approach for constructing perturbed Newtonlike methods in Banach space and their applications. Applied Mathematics and Computation, 102(2–3):203–222, July 15, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300398100176.

Argyros:1999:RNMb

[Arg99e] Ioannis K. Argyros. Results on Newton methods. Part II: Perturbed Newton-like methods in generalized Banach spaces. Applied Mathematics and Computation, 102(2–3):223–236, July 15, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300398100188.

Adomian:1997:NIA

[ARM97] G. Adomian, R. C. Rach, and R. E. Meyers. Numerical integration, analytic continuation, and decomposition. Applied Mathematics and Computation, 88(2–3):95–116, December 30, 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300396000525.

Ananda:1999:PAM

[AS99] Malwane M. A. Ananda and Ashok K. Singh. Performance assessment of multiple engineered barrier systems. Applied Mathematics and Computation, 102(1):25–33, July 1, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300398100115.

Asaithambi:1997:NMS

[Asa97a] N. S. Asaithambi. A numerical method for the solution of the Falkner–Skan Equation. Applied Mathematics and Computation, 81(2–3):259–264, February 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300395003258.

Asaithambi:1997:VTS

[Asa97b] N. S. Asaithambi. A variable time step Galerkin method for a one-dimensional Stefan problem. Applied Mathematics and Computation, 81(2–3): 189–200, February 1997. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
sciencedirect.com/science/
article/pii/0096300395003290.

Asaithambi:1998:FDM

[Asa98] Asai Asaithambi. A finitedifference method for the Falkner–Skan equation. Applied Mathematics and Computation, 92(2–3):135–141, June 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S009630039710042X.

Al-Saqabi:1998:ESF

[ASK98] Bader Al-Saqabi and Virginia S.

Kiryakova. Explicit solutions of fractional integral and differential equations involving Erdélyi–Kober operators. Applied Mathematics and Computation, 95(1):1–13, September 1, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300397100959.

Agarwal:1999:ATT

[AW99a] Ravi P. Agarwal and Fu-Hsiang Wong. An application of topological transversality to non-positive higher order difference equations. Applied Mathematics and Computation, 99(2–3):167–178, March 15, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300398000058.

Agarwal:1999:ESB

[AW99b] Ravi P. Agarwal and Fu-Hsiang Wong. Existence of solutions to (k, n − k − 2) boundary value problems. Applied Mathematics and Computation, 104(1):33–50, September 1, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300398100462.

Aykut:1998:BVP

[AY98] Arzu Aykut and Bünyamin Yildiz. On a boundary value problem for a differential equation with variant retarded argument. Applied Mathematics and Computation, 93(1):63–71, July 1, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300397101151.

Al-Zamel:1996:EHE

[AZK96] A. Al-Zamel and S. Kalla. Epstein–Hubbell elliptic-type integral and its generalizations. Applied Mathematics and Computation, 77(1):9–32, June 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0096300395001476.

Bai:1997:PHI

[Bai97] Zhong-Zhi Bai. Parallel hybrid iteration methods for block bordered linear systems. Applied Mathematics and Computation, 86(1):37–60, September 15, 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300396001725.

Bai:1998:APN

[Bai98] Zhong-Zhi Bai. Asynchronous parallel nonlinear multisplitting relaxation methods for large sparse nonlinear complementarity problems. Applied Mathematics and Computation, 92(1):85–100, June 1, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300397100200.

Bai:1999:AMA

[Bai99a] Zhong-Zhi Bai. Asynchronous multisplitting AOR methods for a class of systems of weakly nonlinear equations. Applied Mathematics and Computation, 98(1):49–59, January 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300397101540.

Bai:1999:CPH

[Bai99b] Zhong-Zhi Bai. A class of parallel hybrid two-stage iteration methods for block bordered linear systems. Applied Mathematics and Computation, 101(2–3):245–267, June 15, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300398100048.

Bardet:1995:DSP

[Bar95] J. P. Bardet. The damping of saturated poroelastic soils during steady-state vibrations. Applied Mathematics and Computation, 67 (1–3):3–31, January/February 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0096300394000526.

Bayram:1997:NMA

[Bay97] Mustafa Bayram. A novel method for analyzing enzyme kinetic systems. Applied Mathematics and Computation, 87(2–3):161–174, December 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.

com/science/article/pii/
S0096300396002275.

Bayram:1998:ACAb

[Bay98a] Mustafa Bayram. Application of computer algebra techniques to affinity binding equations. Applied Mathematics and Computation, 94(1):83–90, August 1, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300397100303.

Bayram:1998:ACAa

[Bay98b] Mustafa Bayram. Application of computer algebra techniques to enzyme kinetics. Applied Mathematics and Computation, 94(1):73–81, August 1, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300397100285.

Bhatti:1996:SOT

[BB96] M. I. Bhatti and A. M. Barry.

Some optimal tests for the equicorrelation coefficient in standard symmetric multivariate normal distribution. Applied Mathematics and Computation, 75(2–3):269–285, March 15, 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300396900715.

Barry:1997:OII

[BBB97] A. M. Barry, S. M. A. Burney, and M. I. Bhatti. Optimum influence of initial observations in regression models with AR(2) errors. Applied Mathematics and Computation, 82(1):57–65, March 1, 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300396000240.

Banerjee:1995:DCC

[BD95] Arun K. Banerjee and Van N. Do. Deployment control of a cable connecting a ship to an underwater vehicle. Applied Mathematics and Computation, 70(2–3):97–116, July 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        009630039400124M.

Bainov:1996:QOS

[BD96] D. D. Bainov and M. B. Dimitrova. Quickly oscillating solutions of operatordifferential equations. Applied Mathematics and Computation, 75(1):1–11, March 1, 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0096300395000658.

Bohm:1998:MBS

[BDJR98] Michael Böhm, Joseph Devinny, Fereidoun Jahani, and Gary Rosen. On a moving-boundary system modeling corrosion in sewer pipes. Applied Mathematics and Computation, 92(2–3):247–269, June 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S009630039710039X.

Bella:1996:MAW

[BDS96] Aimée Bella, Lucien Duckstein, and Ferenc Szidarovszky. A multicriterion analysis of the water allocation conflict in the Upper Rio Grande basin. Applied Mathematics and Computation, 77(2–3):245–265, July 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300395002189.

Becarie:1998:IPS

[Béc98]  Cécile Bécarie.  Intrinsi-

cally parallel solution of systems of linear partial differential equations. Applied Mathematics and Computation, 90(2–3):205–227, March 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300397003998.

Behforooz:1995:CKC

[Beh95] G. Hossein Behforooz. A comparison of the E(3) and nota-knot cubic splines. Applied Mathematics and Computation, 72(2–3):219–223, October 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0096300394001857.

Behforooz:1996:CRS

[Beh96a] G. Hossein Behforooz. Consistency relations of the spline functions derived from a Pascal-like triangle. Applied Mathematics and Computation, 74(2–3):293–297, February 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0096300395001050.

Behforooz:1996:USS

[Beh96b] G. Hossein Behforooz. The use of spline-on-spline for the approximation of Cauchy principal value integrals. Applied Mathematics and Computation, 80(1):23–32, November 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0096300395002774.

Beletsky:1995:SSP

[Bel95] V. V. Beletsky. Some stability problems in applied mechanics. Applied Mathematics and Computation, 70(2–3):117–141, July 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
009630039400123L.

Blount:1997:NDP

[BGY97] S. Blount, A. Galambosi, and S. Yakowitz. Nonlinear and dynamic programming for epidemic intervention. Applied Mathematics and Computation, 86(2–3):123–136, October 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300396001774.

Bogaert:1998:GRP

[BI98] J. Bogaert and I. Impens.

Generating random percolation clusters. Applied Mathematics and Computation, 91(2–3):197–208, May 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300397100182.

Billard:1996:SDG

[Bil96] Edward A. Billard. Stability of dynamic groups in distributed computing systems. Applied Mathematics and Computation, 74(2–3):223–248, February 1996. CODEN AMHCBQ.

ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0096300395000941.

Beddhu:1998:CSU

[BJTW98] Murali Beddhu, Min-Yee Jiang, Lafayette K. Taylor, and David L. Whitfield. Computation of steady and unsteady flows with a free surface around the Wigley hull. Applied Mathematics and Computation, 89 (1–3):67–84, January/February 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300397816487.

Beutler:1999:SRS

[BK99] R. Beutler and B. G. Konopelchenko. Surfaces of revolution via the Schrödinger equation: Construction, integrable dynamics and visualization. Applied Mathematics and Computation, 101(1):13–43, June 1, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300398000368.

Bainov:1995:PBV

[BKM95] Drumi Bainov, Zdzisław Kamont, and Emil Minchev. Periodic boundary value problem for impulsive hyperbolic partial differential equations of first order. Applied Mathematics and Computation, 68(2–3):95–104, March 15, 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
009630039400083G.

Bajpai:1998:BEE

[BKS98] R. P. Bajpai, S. Kumar, and V. A. Sivadasan. Biophoton emission in the evolution of a squeezed state of frequency stable damped oscillator. Applied Mathematics and Computation, 93(2–3):277–288, July 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300397101175.

Blank:1996:SRC

[Bla96]  Luise Blank.  Stability re-

sults for collocation methods for Volterra integral equations. Applied Mathematics and Computation, 79(2–3):267–288, October 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0096300395002707.

Balendran:1995:IIC

[BNN95] B. Balendran and Sia Nemat-Nasser. Integration of inelastic constitutive equations for constant velocity gradient with large rotation. Applied Mathematics and Computation, 67(1– 3):161–195, January/February 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        009630039400058C.

Boyd:1995:LAG

[Boy95]  John P. Boyd.  A lag-

averaged generalization of Euler’s method for accelerating series. Applied Mathematics and Computation, 72(2–3): 143–166, October 1995. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic). URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/009630039400180C.

Boyd:1997:CLU

[Boy97a]  John P. Boyd.  Construc-

tion of Lighthill’s unitary functions: the imbricate series of unity. Applied Mathematics and Computation, 86(1):1–10, September 15, 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300396001452.

Boyd:1997:PCW

[Boy97b]  John P. Boyd.  Peakons

and coshoidal waves: Traveling wave solutions of the Camassa–Holm equation. Applied Mathematics and Computation, 81(2–3):173–187, February 1997. CODEN AMHCBQ.

ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0096300395003266.

Barrio:1999:CCM

[BPE99] Roberto Barrio, Manuel Palacios, and Antonio Elipe. Chebyshev collocation methods for fast orbit determination. Applied Mathematics and Computation, 99(2–3):195–207, March 15, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300398000034.

Baronio:1996:MSI

[BPZ96] A. Baronio, E. Loli Piccolomini, and F. Zama. A method for solving the indirect approximation problem. Applied Mathematics and Computation, 77(2–3):97–107, July 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300395001549.

Baxley:1998:CUC

[BR98] J. V. Baxley and S. B. Robinson. Coexistence in the unstirred chemostat. Applied Mathematics and Computation, 89(1–3):41–65, January/February 1998. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/S0096300397816475.

Barrett:1999:ETC

[BR99a] C. L. Barrett and C. M. Reidys. Elements of a theory of computer simulation I: Sequential CA over random graphs. Applied Mathematics and Computation, 98(2–3):241–259, February 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300397101667.

Bonham:1999:ISA

[BR99b] Charles D. Bonham and Robin M. Reich. Influence of spatial autocorrelation on a fixed-effect model used to evaluate treatment of oil spills. Applied Mathematics and Computation, 106(2–3):149–162, December 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300398101091.

Bai:1997:CCP

[BS97] Zhongzhi Bai and Yangfeng Su. On the convergence of a class of parallel decompositiontype relaxation methods. Applied Mathematics and Computation, 81(1):1–21, January 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.

com/science/article/pii/
0096300395002979.

Buetow:1998:MAF

[BS98] Gerald W. Buetow, Jr. and James S. Sochacki. A more accurate finite difference approach to the pricing of contingent claims. Applied Mathematics and Computation, 91(2–3):111–126, May 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300397100297.

Brugnano:1997:BBV

[BT97] L. Brugnano and D. Trigiante.

Block boundary value methods for linear Hamiltonian systems. Applied Mathematics and Computation, 81(1):49–68, January 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0096300395003088.

Bratsos:1998:FPF

[BT98] A. G. Bratsos and E. H.

Twizell. A family of parametric finite-difference methods for the solution of the sine-Gordon equation. Applied Mathematics and Computation, 93(2–3):117–137, July 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.

com/science/article/pii/
S0096300397101102.

Bursal:1996:IBP

[Bur96]  Faruk H. Bursal.  On in-

terpolating between probability distributions. Applied Mathematics and Computation, 77(2–3):213–244, July 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300395002162.

Bayram:1999:PEE

[BY99] Mustafa Bayram and Bunyamin Yildiz. Parameter estimation of an enzyme kinetic system using computer algebra techniques. Applied Mathematics and Computation, 99(1):93–98, March 1, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300397101783.

Cabada:1994:MMF

[Cab94] Alberto Cabada. The monotone method for first-order problems with linear and nonlinear boundary conditions. Applied Mathematics and Computation, 63(2–3):163–186, July 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0096300394901937. See erra-

tum [Cab95].

Cabada:1995:ERM

[Cab95] Alberto Cabada. Erratum:

“Remarks on the monotone method for first-order problems with linear and nonlinear boundary conditions”, Applied Math. Comput. 63(2– 3), 163–186, July 1994. Applied Mathematics and Computation, 69(2–3):353–354, May 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
009630039400147V. See [Cab94].

Carmody:1997:CHQ

[Car97] Kevin Carmody. Circular and hyperbolic quaternions, octonions, and sedenions — further results. Applied Mathematics and Computation, 84(1):27–47, June 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300396000513.

Cho:1999:NAN

[CBYK99] Chung-Ki Cho, Guo Ben-Yu, and YongHoon Kwon. A new approach for numerical identification of conductivity. Applied Mathematics and Computation, 100(2–3):265–283, May 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300398000277.

Chan:1999:ECS

[CC99a] C. Y. Chan and W. Y. Chan. Existence of classical solutions for degenerate semilinear parabolic problems. Applied Mathematics and Computation, 101(2–3):125–149, June 15, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300398100024.

Choe:1999:SMS

[CC99b] Hi Jun Choe and Jeong Ho Chu. A spectral method for Stokes equations. Applied Mathematics and Computation, 104(2–3):131–151, September 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300398100620.

Cengiz:1997:SEH

[Cen97]  Ahmet Cengiz.  A study

on the Epstein–Hubbell generalized elliptic-type integral using residue theory. Applied Mathematics and Computation, 83(1):19–26, April 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300396000410.

Carl:1996:SMI

[CG96] S. Carl and C. Grossmann.

Smoothing and monotone iterations for elliptic differential inclusions. Applied Mathematics and Computation, 74(1):15–35, January 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0096300395000704.

Chang:1998:PES

[CG98] Carolina Chang and Henryk Gzyl. Parameter estimation in superposition of decaying exponentials. Applied Mathematics and Computation, 96(2–3):101–116, November 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300397101084.

Chen:1996:ADT

[CH96] Chaó-Kuang Chen and Shing-Huei Ho. Application of differential transformation to eigenvalue problems. Applied Mathematics and Computation, 79(2–3):173–188, October 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0096300395002537.

Chen:1999:SPD

[CH99] Cha’o Kuang Chen and Shing Huei Ho. Solving partial differential equations by twodimensional differential transform method. Applied Mathematics and Computation, 106(2–3):171–179, December 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300398101157.

Chang:1995:PLA

[Cha95a] Ching Lung Chang. Piecewise linear approach to the Stokes equations in 3-D. Applied Mathematics and Computation, 72(1):61–75, September 15, 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0096300394001765.

Chang:1995:DBE

[Cha95b] Mou-Hsiung Chang. On duality between estimation and control for linear stochastic functional evolution equations in Hilbert spaces. Applied Mathematics and Computation, 68(1):51–70, March 1, 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        009630039400089M.

Chang:1996:LSF

[Cha96a] Ching Lung Chang. Leastsquares finite element for second order boundary value problem with optimal rates of convergence. Applied Mathematics and Computation, 76(2–3):267–284, May 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0096300395001646.

Chang:1996:LDN

[Cha96b] Mou-Hsiung Chang. Large deviation for Navier–Stokes equations with small stochastic perturbation. Applied Mathematics and Computation, 76 (1):65–93, April 1996. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
sciencedirect.com/science/
article/pii/0096300395001506.

Chen:1995:RCD

[Che95] Y. H. Chen. Robust control design for a class of mismatched uncertain nonlinear systems. Applied Mathematics and Computation, 70(2–3):155–167, July 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
009630039400105D.

Chen:1996:IMC

[Che96a] Yong-Lin Chen. Iterative methods for computing the generalized inverses A_{T,S}^{(2)} of a matrix A. Applied Mathematics and Computation, 75(2–3):207–222, March 15, 1996. CODEN AMHCBQ.

ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0096300396900636.

Chernousko:1996:CES

[Che96b] F. L. Chernousko. Control of elastic systems by bounded distributed forces. Applied Mathematics and Computation, 78 (2–3):103–110, September 1, 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        009630039600001X.

Chen:1997:IMS

[Che97]  Yong-Lin Chen.  Iterative

methods for solving restricted linear equations. Applied Mathematics and Computation, 86(2–3):171–184, October 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300396001804.

Chukwu:1996:ULC

[Chu96]  E. N. Chukwu.  Universal

laws for the control of global economic growth with nonlinear hereditary dynamics. Applied Mathematics and Computation, 78(1):19–81, August 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.

com/science/article/pii/
0096300395001557.

Chukwu:1997:CND

[Chu97]  E. N. Chukwu.  Control

of nonlinear delay differential equations in W _{2} ^{(1)} with economic applications. Applied Mathematics and Computation, 85(1):17–59, August 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300396001087.

Chukwu:1998:CNE

[Chu98] Ethelbert N. Chukwu. On the controllability of nonlinear economic systems with delay: the Italian example. Applied Mathematics and Computation, 95 (2–3):245–274, September 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300397100790.

Cicci:1999:FPT

[Cic99] David A. Cicci. Filter performance in target tracking using space-based observers. Applied Mathematics and Computation, 99(2–3):275–293, March 15, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300398000071.

Chan:1995:TEM

[CK95] C. Y. Chan and P. C. Kong. A thermal explosion model. Applied Mathematics and Computation, 71 (2–3):201–210, September 1, 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        009630039400154V.

Chen:1999:GNR

[CK99] X. B. Chen and M. M. Kostreva. A generalization of the norm-relaxed method of feasible directions. Applied Mathematics and Computation, 102(2–3):257–272, July 15, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300398100255.

Chen:1998:DTT

[CL98a] Chieh-Li Chen and Yung-Chin Liu. Differential transformation technique for steady nonlinear heat conduction problems. Applied Mathematics and Computation, 95(2– 3):155–164, September 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300397100960.

Chen:1998:PPN

[CL98b] Jinru Chen and Likang Li. Preconditioning projection nonconforming element method for the lowest-order Raviart– Thomas mixed triangular element method. Applied Mathematics and Computation, 93(1):31–49, July 1, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300397101126.

Cao:1999:CRP

[CL99] Zhi-Hao Cao and Zhong-Yun Liu. Convergence of relaxed parallel multisplitting methods with different weighting schemes. Applied Mathematics and Computation, 106(2–3):181–196, December 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300398101200.

Chen:1997:EBE

[CLC97] Cha’o-Kuang Chen, Jin-Mu Lin, and Chieh-Li Chen. Error bounds estimate of weighted residuals method using genetic algorithms. Applied Mathematics and Computation, 81(2–3):207–219, February 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        S0096300396000173.

Cordero:1996:GQP

[CMM96] M. Cordero, C. Martin, and J. Miller. Gaussian quadrature for products of exponential functions. Applied Mathematics and Computation, 79(2–3):189–202, October 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0096300395002545.

Cabada:1997:RCI

[CN97] Alberto Cabada and Juan J. Nieto. Rapid convergence of the iterative technique for first order initial value problems. Applied Mathematics and Computation, 87(2–3):217–226, December 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300396002858.

Cicci:1996:FSE

[CO96] David A. Cicci and David L. Oakley. Filter sensitivity in exoatmospheric target vehicle tracking. Applied Mathematics and Computation, 79(2–3):249–265, October 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.

com/science/article/pii/
0096300395002693.

Coskun:1999:CSF

[Cos99]  Erhan Coskun.  Computa-

tional simulation of flux trapping and vortex pinning in Type-II superconductors. Applied Mathematics and Computation, 106(1):31–49, December 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300398100899.

Crombez:1995:PPM

[Cro95] G. Crombez. A parallel projection method based on sequential most remote set in convex feasibility problems. Applied Mathematics and Computation, 72(2–3):113–124, October 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0096300394001776.

Choi:1999:SHO

[CRS99] Junesang Choi, Arjun K.

Rathie, and H. M. Srivastava. Some hypergeometric and other evaluations of ζ(2) and allied series. Applied Mathematics and Computation, 104(2–3):101–108, September 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.

com/science/article/pii/
S0096300398100826.

Chen:1995:ORG

[CS95] Ming-Po Chen and H. M. Srivastava. Orthogonality relations and generating functions for Jacobi polynomials and related hypergeometric functions. Applied Mathematics and Computation, 68(2–3):153–188, March 15, 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
009630039400092I.

Chou:1996:MFE

[CS96] Chi-Chin Chou and William A. Sethares. Multiplicationfree evaluation of polynomials via a Stochastic Bernstein Representation. Applied Mathematics and Computation, 79(1):1–25, September 15, 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0096300395001565.

Chen:1997:FCO

[CS97] Ming-Po Chen and H. M. Srivastava. Fractional calculus operators and their applications involving power functions and summation of series. Applied Mathematics and Computation, 81(2–3):287–304, February 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S009630039500310X.

Chang:1998:MQE

[CS98] Feng Cheng Chang and Ching-Tzong Su. More on quick evaluation of determinants. Applied Mathematics and Computation, 93(1):97–99, July 1, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300397100443.

Chiue:1999:CCP

[CS99] Wen-Sz Chiue and Bih-Sheue Shieh. On connectivity of the Cartesian product of two graphs. Applied Mathematics and Computation, 102(2–3):129–137, July 15, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300398100413.

Chen:1998:SOF

[CSY98] Ming-Po Chen, H. M. Srivastava, and Ching-Shu Yu. Some operators of fractional calculus and their applications involving a novel class of analytic functions. Applied Mathematics and Computation, 91(2–3):285–296, May 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        S0096300397100340.

Choi:1998:IIF

[CSZ98] Junesang Choi, H. M. Srivastava, and Nan-Yue Zhang. Integrals involving a function associated with the Euler– Maclaurin summation formula. Applied Mathematics and Computation, 93(2–3):101–116, July 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300397100704.

Chen:1996:NFM

[CT96] Cha’o-Kuang Chen and Tzu-Li Tien. A new forecasting method for time continuous model of dynamic system. Applied Mathematics and Computation, 80(2–3):225–244, December 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0096300395002960.

Chen:1997:NFM

[CT97] Cha’o-Kuang Chen and Tzu-Li Tien. A new forecasting method of discrete dynamic system. Applied Mathematics and Computation, 86(1):61–84, September 15, 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300396001737.

Charafi:1997:NAR

[CW97] A. Charafi and L. C. Wrobel. A new h-adaptive refinement scheme for the boundary element method using local reanalysis. Applied Mathematics and Computation, 82(2–3):239–271, March 15, 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300396000288.

Chan:1998:IEG

[CY98] C. Y. Chan and S. I. Yuen.

Impulsive effects on global existence of solutions for degenerate semilinear parabolic equations. Applied Mathematics and Computation, 90(2–3):97–116, March 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300396003530.

Chang:1999:EOH

[CY99] Mou-Hsiung Chang and Roger K.

Youree. The European option with hereditary price structures: Basic theory. Applied Mathematics and Computation, 102(2–3):279–296, July 15, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300398100358.

Cha:1999:PES

[CyGK99] Ki Hyun Cha, Ben yu Guo, and Yong Hoon Kwon. Parameter estimation by spectral approximation. Applied Mathematics and Computation, 104(1):1–14, September 1, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300398100553.

Chan:1999:BSS

[CZ99] C. Y. Chan and J. K. Zhu. Blow-up of solutions of semilinear Euler–Poisson– Darboux equations with nonlocal boundary conditions. Applied Mathematics and Computation, 99(1):17–28, March 1, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300397101710.

Davis:1998:SCP

[Dav98]  Stephen F. Davis.  Shock

capturing with Padé methods. Applied Mathematics and Computation, 89(1– 3):85–98, January/February 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300397816499.

deBarros:1995:SDL

[dBL95] F. C. P. de Barros and J. E. Luco. Stresses and displacements in a layered half-space for a moving line load. Applied Mathematics and Computation, 67(1– 3):103–134, January/February 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
009630039400056A.

Degot:1996:MPC

[Dég96] Jérôme Dégot. Massively parallel computation of manyvariable polynomial interpolation. Applied Mathematics and Computation, 80(2–3):105– 125, December 1996. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
sciencedirect.com/science/
article/pii/0096300395002839.

DeKlerk:1995:AMN

[DEV95] J. H. De Klerk, D. Eyre, and L. M. Venter. L_{p}- approximation method for the numerical solution of singular integral equations. Applied Mathematics and Computation, 72(2–3):285–300, October 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
009630039500072P.

Dolezal:1996:NSD

[DF96] Jaroslav Doležal and Jiří Fidler. Numerical solution of dynamic optimization problems using parametrization and Op^{ti} A software. Applied Mathematics and Computation, 78 (2–3):111–121, September 1, 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0096300396000021.

Duchting:1996:CSA

[DGU96] Werner Düchting, Thomas Ginsberg, and Waldemar Ulmer. Computer simulation applied to radiation therapy in cancer research. Applied Mathematics and Computation, 74(2–3):191–207, February 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0096300395000925.

Daripa:1999:NSI

[DH99a] Prabir Daripa and Wei Hua. A numerical study of an illposed Boussinesq equation arising in water waves and nonlinear lattices: Filtering and regularization techniques. Applied Mathematics and Computation, 101(2–3):159–207, June 15, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.

com/science/article/pii/
S009630039810070X.

Dror:1999:SRU

[DH99b] Moshe Dror and Bruce C.

Hartman. Stopping rules for utility functions and the St. Petersburg gamble. Applied Mathematics and Computation, 98(2–3):279–291, February 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300397101692.

DiBella:1998:PJM

[Di 98] Beatrice Di Bella. On a paper by J. M. Soriano. Applied Mathematics and Computation, 92(2–3):297–298, June 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300397100510.

Dietzenbacher:1996:AFB

[Die96] Erik Dietzenbacher. An algorithm for finding blocktriangular forms. Applied Mathematics and Computation, 76(2–3):161–171, May 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0096300395001573.

Ding:1998:MEM

[Din98] J. Ding. A maximum entropy method for solving Frobenius– Perron operator equations. Applied Mathematics and Computation, 93(2–3):155–168, July 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300397100613.

Deeba:1996:DMA

[DK96] E. Y. Deeba and S. A. Khuri. The decomposition method applied to Chandrasekhar Hequation. Applied Mathematics and Computation, 77(1):67–78, June 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0096300395001883.

Dunyak:1998:AIS

[DML98] James Dunyak, Clyde Martin, and Richard Lampe. Analysis of the influence of social structure on a measles epidemic. Applied Mathematics and Computation, 92(2–3):283–296, June 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300397100546.

Djidjeli:1998:PIS

[DPTT98] K. Djidjeli, W. G. Price, P. Temarel, and E. H. Twizell. Partially implicit schemes for the numerical solutions of some non-linear differential equations. Applied Mathematics and Computation, 96(2–3):177–207, November 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300397101333.

deRomero:1998:SAF

[dRS98] Susana Salinas de Romero and H. M. Srivastava. Some applications of fractional calculus involving summation of infinite series. Applied Mathematics and Computation, 90(2–3):129–142, March 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300397003949.

Dziok:1999:CAF

[DS99] J. Dziok and H. M. Srivastava.

Classes of analytic functions associated with the generalized hypergeometric function. Applied Mathematics and Computation, 103(1):1–13, August 1, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300398100425.

deSampaio:1997:TRA

[dSYS97] R. J. B. de Sampaio, Jin-Yun Yuan, and Wen-Yu Sun. Trust region algorithm for nonsmooth optimization. Applied Mathematics and Computation, 85 (2–3):109–116, September 1, 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300396001129.

Duan:1999:RTD

[Dua99a]  Jinqiao Duan.  A remark

on the three dimensional baroclinic quasi-geostrophic dynamics. Applied Mathematics and Computation, 106(2–3):285–288, December 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300398101303.

Duan:1999:TPQ

[Dua99b] Jinqiao Duan. Time-periodic quasigeostrophic motion under dissipation and forcing. Applied Mathematics and Computation, 102(2–3):121–127, July 15, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300398100346.

Dang-Vu:1995:HBS

[DVD95] H. Dang-Vu and C. Delcarte. Hopf bifurcation and strange attractors in Chebyshev spectral solutions of the Burgers equation. Applied Mathematics and Computation, 73(2–3):99–113, December 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0096300394002428.

Ding:1999:FEM

[DZ99] Jiu Ding and Aihui Zhou.

A finite element method for the Frobenius–Perron operator equation. Applied Mathematics and Computation, 102(2–3):155–164, July 15, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300398100395.

El-Bassiouny:1999:RTD

[EB99] A. F. El-Bassiouny. Response of a three-degree-of-freedom system with cubic non-linearities to harmonic excitation. Applied Mathematics and Computation, 104(1):65–84, September 1, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300398100516.

Eskandari:1995:UMC

[EFS95] A. Eskandari, P. Ffolliott, and F. Szidarovszky. Uncertainty and method choice in discrete multiobjective programming problems. Applied Mathematics and Computation, 69(2–3):335–351, May 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        009630039400145T.

Ezquerro:1997:CPI

[EGH97] J. A. Ezquerro, J. M. Gutiérrez, and M. A. Hernández. A construction procedure of iterative methods with cubical convergence. Applied Mathematics and Computation, 85 (2–3):181–199, September 1, 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300396001348.

Ezquerro:1998:CPI

[EGH98] J. A. Ezquerro, J. M. Gutiérrez, and M. A. Hernández. A construction procedure of iterative methods with cubical convergence II: Another convergence approach. Applied Mathematics and Computation, 92(1):59–68, June 1, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300397100315.

Eisenfeld:1998:IIB

[Eis98] J. Eisenfeld. Identification, interval bounds, and linear algebraic estimation for a class of nonlinear problems. Applied Mathematics and Computation, 89(1– 3):17–29, January/February 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300397816451.

Egerstedt:1998:CTM

[EM98] Magnus Egerstedt and Clyde Martin. A control theoretic model of the combined planar motion of the human head and eye. Applied Mathematics and Computation, 90(1):61–95, March 1, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300396003517.

El-Misiery:1996:ACJ

[EMEH96] A. E. M. El-Misiery and El-Sayed M. El-Horbaty. An algorithm for calculating Jones polynomials. Applied Mathematics and Computation, 74(2– 3):249–259, February 1996. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
sciencedirect.com/science/
article/pii/0096300395000968.

El-Naggar:1995:RNH

[ENAAA95] A. M. El-Naggar, A. M. Abd-Alla, and S. M. Ahmed. On the rotation of a non-homogeneous composite infinite cylinder of orthotropic material. Applied Mathematics and Computation, 69(2–3):147–157, May 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        009630039400074E.

El-Naggar:1996:WPL

[ENS96] A. M. El-Naggar and M. M. Saliem. Wave propagation in layered media under initial stresses. Applied Mathematics and Computation, 74(2–3):95–117, February 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        009630039500064X.

El-Owaidy:1999:SNI

[EORA99] H. El-Owaidy, A. Ragab, and A. Abdeldaim. On some new integral inequalities of Growall–Bellman type. Applied Mathematics and Computation, 106(2–3):289–303, December 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300398101315.

El-Sayed:1999:FOE

[ES99] Ahmed M. A. El-Sayed. Fractionalorder evolutionary integral equations. Applied Mathematics and Computation, 98(2–3): 139–146, February 1999. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
sciencedirect.com/science/
article/pii/S0096300397101655.

El-Sayed:1995:MFD

[ESI95] A. M. A. El-Sayed and A. G.

Ibrahim. Multivalued fractional differential equations. Applied Mathematics and Computation, 68(1):15–25, March 1, 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
009630039400080N.

Eom:1998:NSL

[ESL98] Tae-Dok Eom, Masanori Sugisaka, and Ju-Jang Lee. New skill learning paradigm using various kinds of neurons. Applied Mathematics and Computation, 91(1):9–22, April 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S009630039710011X.

El-Sayed:1999:BPA

[ESR99] A. M. A. El-Sayed and S. Z.

Rida. Bell polynomials of arbitrary (fractional) orders+. Applied Mathematics and Computation, 106(1):51–62, December 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        S0096300398100887.

Fabien:1996:NSC

[Fab96] Brian C. Fabien. Numerical solution of constrained optimal control problems with parameters. Applied Mathematics and Computation, 80(1): 43–62, November 1996. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic). URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300395002804.

Fernandez-Berdaguer:1996:IPE

[FBSS96] Elena M. Fernandez-Berdaguer, Juan E. Santos, and Dongwoo Sheen. An iterative procedure for estimation of variable coefficients in a hyperbolic system. Applied Mathematics and Computation, 76(2–3):213–250, May 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0096300395001611.

Fridrich:1995:RBO

[FG95] Jiri Fridrich and James F. Geer. Reconstruction of blurred orbits under finite resolution. Applied Mathematics and Computation, 71(2–3):227–245, September 1, 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        009630039400156X.

Ford:1995:APF

[FH95] R. A. Ford and M. H. Hamdan.

Analysis of the polar form of the von Mises transformation. Applied Mathematics and Computation, 72(2–3):205–217, October 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0096300394001846.

Ford:1998:CPF

[FH98] R. A. Ford and M. H.

Hamdan. Coupled parallel flow through composite porous layers. Applied Mathematics and Computation, 97 (2–3):261–271, December 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300397101412.

Ford:1996:ECT

[FHG96a] R. A. Ford, M. H. Hamdan, and L. E. Garey. Effects of Coordinate transformation and differencing schemes on matrix structure and convergence history. Applied Mathematics and Computation, 77(1):53–66, June 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0096300395001913.

Fritz:1996:ESE

[FHG96b] Christophe Fritz, Michel Hassenforder, and Gérard L. Gissinger. Evaluating systems of equations: Toward a formal approach. Applied Mathematics and Computation, 75(2–3):223–237, March 15, 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
009630039690065X.

Fang:1997:HPA

[FHK97] Liping Fang, Keith W. Hipel, and D. Marc Kilgour. How penalty affects enforcement of environmental regulations. Applied Mathematics and Computation, 83(2–3):281–301, May 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300396001890.

Finkelstein:1995:EFD

[Fin95] Mark Finkelstein. Estimating the frequency distribution of the numbers bet on the California lottery. Applied Mathematics and Computation, 69(2–3):195–207, May 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
009630039400126O.

Fiorito:1995:PPM

[Fio95] Giovanni Fiorito. On properties of periodically monotone sequences. Applied Mathematics and Computation, 72(2–3):259–275, October 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0096300394001879.

Fu:1999:UBS

[FL99] Xilin Fu and Xinzhi Liu. Uniform boundedness and stability criteria in terms of two measures for impulsive integrodifferential equations. Applied Mathematics and Computation, 102(2–3):237–255, July 15, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300398100371.

Freedman:1996:BPT

[FM96] H. I. Freedman and P. Moson. Bifurcations in persistence theory. Applied Mathematics and Computation, 79(2–3):125–136, October 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0096300395002480.

Foroutan-pour:1999:AIB

[FpDS99] K. Foroutan-pour, P. Dutilleul, and D. L. Smith. Advances in the implementation of the box-counting method of fractal dimension estimation. Applied Mathematics and Computation, 105(2–3): 195–210, November/December 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300398100966.

Fei:1995:NSN

[FPGV95] Zhang Fei, Víctor M. Pérez-García, and Luis Vázquez. Numerical simulation of nonlinear Schrödinger systems: a new conservative scheme. Applied Mathematics and Computation, 71(2–3):165–177, September 1, 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
009630039400152T.

Fridrich:1997:DTD

[Fri97] Jiri Fridrich. Discrete-time dynamical systems under observational uncertainty. Applied Mathematics and Computation, 82(2–3):181–205, March 15, 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S009630039600029X.

Franke:1998:SPD

[FS98] C. Franke and R. Schaback.

Solving partial differential equations by collocation using radial basis functions. Applied Mathematics and Computation, 93(1):73–82, July 1, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300397101047.

Frontini:1997:ECS

[FT97] Marco Frontini and Aldo Tagliani. Entropy-convergence in Stieltjes and Hamburger moment problem. Applied Mathematics and Computation, 88(1):39–51, December 15, 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300396003050.

Frontini:1998:MEL

[FT98] Marco Frontini and Aldo Tagliani. Maximum entropy and lacunary Stieltjes moment problem. Applied Mathematics and Computation, 97 (2–3):183–196, December 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300397101485.

Feng:1998:APP

[FW98] Wei Feng and Fan Wang.

Asymptotic periodicity and permanence in a competitiondiffusion system with discrete delays. Applied Mathematics and Computation, 89(1– 3):99–110, January/February 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300397816505.

Gao:1998:NCQ

[Gao98] Zhi-Hao Gao. A note on convergence of quasi-minimal residual smoothing. Applied Mathematics and Computation, 93(2–3):289–297, July 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300397100741.

Garner:1998:ERC

[Gar98] J. B. Garner. Existence results for a class of diffusion-reaction problems. Applied Mathematics and Computation, 89(1– 3):111–118, January/February 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300397816517.

Golberg:1998:OCR

[GB98] Michael Golberg and Harold Bowman. Optimal convergence rates for some discrete projection methods. Applied Mathematics and Computation, 96(2–3):237–271, November 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S009630039710131X.

Gulliver:1999:AGC

[GBS99] T. Aaron Gulliver, Vijay K.

Bhargava, and Jack M. Stein. Q-ary Gray codes and weight distributions. Applied Mathematics and Computation, 103 (1):97–109, August 1, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/S0096300398100504.

Guttalu:1996:SSP

[GF96] Ramesh S. Guttalu and Henryk Flashner. Stability study of a periodic system by a periodto-period mapping. Applied Mathematics and Computation, 78(2–3):123–135, September 1, 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0096300396000033.

Garey:1996:HOM

[GG96] L. E. Garey and C. J. Gladwin. Higher-order methods for eigenvalues of second-order ordinary differential equations with boundary conditions. Applied Mathematics and Computation, 76(2–3):293–300, May 1996. CODEN AMHCBQ.

ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0096300395001875.

Garey:1997:USM

[GGS97] L. E. Garey, C. J. Gladwin, and R. E. Shaw. Unconditionally stable methods for second-order Fredholm integrodifferential equations. Applied Mathematics and Computation, 81(2–3):275–286, February 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300395003282.

Ghorbanzadeh:1996:DRC

[Gho96] Dariush Ghorbanzadeh. Detection of random change point in one-parameter exponential families. Applied Mathematics and Computation, 77(2–3):167–177, July 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300395002103.

Gu:1996:OIP

[GKT96] Mu Gu, Robert E. Kalaba, and Grant A. Taylor. Obtaining initial parameter estimates for chaotic dynamical systems using linear associative memories. Applied Mathematics and Computation, 76(2–3):143–159, May 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0096300395001530.

Guoqiang:1995:AEE

[GL95] Han Guoqiang and Zhang Liqing. Asymptotic error expansion of a collocationtype method for Hammerstein equations. Applied Mathematics and Computation, 72(1):1–19, September 15, 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0096300394001606.

Gunzburger:1999:AAO

[GL99] Max D. Gunzburger and Hyung-Chun Lee. Analysis and approximation of optimal control problems for first-order elliptic systems in three dimensions. Applied Mathematics and Computation, 100(1):49–70, April 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300398000174.

Gopalsamy:1998:GHB

[GLL98] K. Gopalsamy, Issic K. C.

Leung, and Pingzhou Liu. Global Hopf-bifurcation in a neural netlet. Applied Mathematics and Computation, 94 (2–3):171–192, August 15, 1998. CODEN AMHCBQ.

ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S009630039710087X.

Gouesbet:1999:CEH

[GMGCL99] G. Gouesbet, S. Meunier-Guttin-Cluzel, and C. Letellier. Computer evaluation of Homfly polynomials by using Gauss codes, with a skein-template algorithm. Applied Mathematics and Computation, 105(2–3): 271–289, November/December 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300398101066.

Guariso:1996:PSC

[GMS96] Giorgio Guariso, Vittorio Maniezzo, and Paola Salomoni. Parallel simulation of a cellular pollution model. Applied Mathematics and Computation, 79(1):27–41, September 15, 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0096300395002146.

Golberg:1995:NSR

[Gol95] Michael Golberg. A note on the sparse representation of discrete integral operators. Applied Mathematics and Computation, 71(2–3):97–118, September 1, 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
009630039400146U.

Golberg:1996:RDN

[Gol96] Michael Golberg. Recent developments in the numerical evaluation of particular solutions in the boundary element method. Applied Mathematics and Computation, 75(1):91–101, March 1, 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0096300395001239.

Golberg:1999:NDM

[Gol99] Michael A. Golberg. A note on the decomposition method for operator equations. Applied Mathematics and Computation, 106(2–3):215–220, December 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300398101248.

Goyon:1997:PNM

[GP97] Olivier Goyon and Pascal Poullet. Preconditioned Newton methods using incremental unknowns methods for the resolution of a steady-state Navier– Stokes-like problem. Applied Mathematics and Computation, 87(2–3):289–311, December 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300396003049.

Graef:1998:SPM

[GQS98] J. R. Graef, C. Qian, and P. W. Spikes. Stability in a population model. Applied Mathematics and Computation, 89(1– 3):119–132, January/February 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300397816529.

Galligani:1997:TSA

[GR97] E. Galligani and V. Ruggiero. The Two-Stage Arithmetic Mean Method. Applied Mathematics and Computation, 85(2–3):245–264, September 1, 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300396001397.

Grace:1995:OMN

[Gra95]  S. R. Grace.  Oscillations

of mixed neutral functional differential equations. Applied Mathematics and Computation, 68(1):1–13, March 1, 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        009630039400075F.

Gragnani:1997:BAT

[Gra97] Alessandra Gragnani. Bifurcation analysis of two predatorprey models. Applied Mathematics and Computation, 85 (2–3):97–108, September 1, 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300396001117.

Gremaud:1997:NSP

[Gre97] Pierre-Alain Gremaud. Numerical simulation of pattern formation in grain flows. Applied Mathematics and Computation, 84(2–3):145–162, July 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300396000847.

Grimm:1998:OSV

[Gri98]  Stefan Grimm.  On or-

bit sum values of elements of finite order. Applied Mathematics and Computation, 97(1):17–22, December 1, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300397101539.

Goel:1995:SCU

[GS95] Narendra S. Goel and Bo Shen.

Symbolic computation using Lsystems II: Extensions. Applied Mathematics and Computation, 69(2–3):227–240, May 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
009630039400129R.

Gould:1997:SCI

[GS97] H. W. Gould and H. M. Srivastava. Some combinatorial identities associated with the Vandermonde convolution. Applied Mathematics and Computation, 84(2–3):97–102, July 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300396000549.

Garey:1999:PAS

[GS99] L. E. Garey and R. E. Shaw. A parallel algorithm for solving Toeplitz linear systems. Applied Mathematics and Computation, 100(2–3):241–247, May 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300398000289.

Goodman:1995:LDE

[GSC95] T. N. T. Goodman, H. B. Said, and L. H. T. Chang. Local derivative estimation for scattered data interpolation. Applied Mathematics and Computation, 68(1):41–50, March 1, 1995. CODEN AMHCBQ.

ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
009630039400086J.

Gunji:1996:BSD

[GSI96] Yukio-Pegio Gunji, Hisato Sadaoka, and Keisuke Ito. Bootstrapping system defined by inconsistent relation between Boolean and non-Boolean algebra. Applied Mathematics and Computation, 79(1):43–97, September 15, 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0096300395002219.

Graef:1998:SID

[GSZS98] John Graef, Rathasingham Shivaji, Jiangping Zhu, and Bharat Soni. Special issue on differential equations and computational simulations II. Applied Mathematics and Computation, 89 (1–3):1–3, January/February 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300397816438.

Gu:1999:CFE

[Gu99] Haiming Gu. Characteristic finite element methods for nonlinear Sobolev equations. Applied Mathematics and Computation, 102(1):51–62, July 1, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S009630039810019X.

Guo:1999:BVP

[Guo99] Dajun Guo. Boundary value problems for impulsive integrodifferential equations on unbounded domains in a Banach space. Applied Mathematics and Computation, 99 (1):1–15, March 1, 1999. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic). URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/S0096300397101746.

Gupta:1998:GMP

[Gup98] Chaitan P. Gupta. A generalized multi-point boundary value problem for second order ordinary differential equations. Applied Mathematics and Computation, 89(1– 3):133–146, January/February 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300397816530.

Gao:1999:NIM

[GXQ99] Weiguo Gao, Jungong Xue, and Yanyun Qu. A new implementation of EN method. Applied Mathematics and Computation, 98(2–3):199–208, February 1999. CODEN AMHCBQ.

ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300397101576.

Gzyl:1995:MRF

[Gzy95] Henryk Gzyl. Maxentropic reconstruction of Fourier and Laplace transforms. Applied Mathematics and Computation, 73(2–3):181–189, December 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0096300394002487.

Gzyl:1997:MRF

[Gzy97] Henryk Gzyl. Maxentropic reconstruction of Fourier and Laplace transforms under nonlinear constraints. Applied Mathematics and Computation, 85(2–3):117–126, September 1, 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300396001166.

Haber:1995:PCN

[Hab95] R. Haber. Predictive control of nonlinear dynamic processes. Applied Mathematics and Computation, 70(2–3):169–184, July 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
009630039400122K.

Hady:1995:MCB

[Had95] F. M. Hady. Mixed convection boundary-layer flow of non-Newtonian fluids on a horizontal plate. Applied Mathematics and Computation, 68(2–3):105–112, March 15, 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        009630039400084H.

Haidar:1997:RPI

[Hai97] Nassar H. S. Haidar. Recursive pseudo-inversion of the Laplace transform on the real line. Applied Mathematics and Computation, 84(2–3):213–220, July 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300396000872.

Hale:1998:DNA

[Hal98]  Jack K. Hale.  Dynam-

ics of numerical approximations. Applied Mathematics and Computation, 89 (1–3):5–15, January/February 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S009630039781644X.

Hamdan:1997:NCS

[Ham97] M. H. Hamdan. Natural coordinate system approach to coupled n-phase fluid flow in curved domains. Applied Mathematics and Computation, 85 (2–3):297–304, September 1, 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300396001440.

Hamdan:1998:AAE

[Ham98a] M. H. Hamdan. An alternative approach to exact solutions of a special class of Navier–Stokes flows. Applied Mathematics and Computation, 93(1):83–90, July 1, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300397100479.

Hamdan:1998:NCU

[Ham98b] M. H. Hamdan. A note on computational uncertainty. Applied Mathematics and Computation, 94(2–3):285–291, August 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300397100881.

Han:1998:PCM

[Han98]  Qiaoming Han.  Projec-

tion and contraction methods for semidefinite programming. Applied Mathematics and Computation, 95(2– 3):275–289, September 15, 1998. CODEN AMHCBQ.

ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300397101138.

Hartfiel:1995:IMS

[Har95]  D. J. Hartfiel.  Indetermi-

nate Markov systems. Applied Mathematics and Computation, 72(1):51–59, September 15, 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0096300394001754.

Hartfiel:1997:SBQ

[Har97] D. J. Hartfiel. System behavior in quotient systems. Applied Mathematics and Computation, 81(1):31–48, January 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0096300395003002.

Honma:1998:AEH

[HAST98] Noriyasu Honma, Kenichi Abe, Mitsuo Sato, and Hiroshi Takeda. Adaptive evolution of holon networks by an autonomous decentralized method. Applied Mathematics and Computation, 91(1): 43–61, April 1998. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic). URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/S009630039710008X.

Hraba:1996:UMM

[HD96] Tomáš Hraba and Jaroslav Doležal. Use of mathematical model for effectivity estimation of HIV infection therapies. Applied Mathematics and Computation, 78 (2–3):153–161, September 1, 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0096300396000057.

He:1996:DDM

[He96] Qiming He. Domain decomposition method for nonlinear generalized Schrödingertype systems: Semi-discrete problem. Applied Mathematics and Computation, 77(1):33–52, June 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
009630039500159X.

He:1998:LCG

[He98]  Jiaxing He.  On the lin-

ear combination of Grünwald polynomial operator. Applied Mathematics and Computation, 96(2–3):117–126, November 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300397101345.

Hamdan:1995:SPF

[HF95] M. H. Hamdan and R. A. Ford. Single-phase flow through porous channels part II: Flow models, critical length, and viscous separation. Applied Mathematics and Computation, 69(2–3):241–254, May 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
009630039400132N.

Hamdan:1998:NSG

[HFR98] M. H. Hamdan, R. A. Ford, and R. Abu Rabia. Numerical simulation of gasparticulate flow through curvilinear porous channels. Applied Mathematics and Computation, 94(2–3):267–284, August 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300397100947.

He:1999:PBV

[HG99] Zhimin He and Weigao Ge. Periodic boundary value problem for first order impulsive delay differential equations. Applied Mathematics and Computation, 104(1):51–63, September 1, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300398100590.

Hu:1996:SSC

[HH96a] Guang-Di Hu and Guang-Da Hu. Some simple criteria for stability of neutral delaydifferential systems. Applied Mathematics and Computation, 80(2–3):257–271, December 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0096300395003010.

Hu:1996:SDD

[HH96b] Guang-Di Hu and Guang-Da Hu. Stability of discretedelay systems: Boundary criteria. Applied Mathematics and Computation, 80(2–3):95– 104, December 1996. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
sciencedirect.com/science/
article/pii/0096300395002820.

Hu:1997:SND

[HH97] Guang-Da Hu and Guang-Di Hu. Stability of neutral delay-differential systems: Boundary criteria. Applied Mathematics and Computation, 87(2–3):247–259, December 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300396003001.

Hickernell:1999:RBF

[HH99] Fred J. Hickernell and Y. C.

Hon. Radial basis function approximations as smoothing splines. Applied Mathematics and Computation, 102(1): 1–24, July 1, 1999. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
sciencedirect.com/science/
article/pii/S0096300398100127.

Hartung:1998:PIC

[HHT98] Ferenc Hartung, Terry L. Herdman, and Janos Turi. Parameter identification in classes of hereditary systems of neutral type. Applied Mathematics and Computation, 89(1– 3):147–160, January/February 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300397816542.

Huang:1997:ENS

[HHZ97] Hongci Huang, Weimin Han, and Jinshi Zhou. Extrapolation of numerical solutions for elliptic problems on corner domains. Applied Mathematics and Computation, 83(1):53–67, April 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S009630039600046X.

Hohl:1997:MPS

[HI97] Jean-Christophe Hohl and ICM. Massively parallel search for linear factors in polynomials with many variables. Applied Mathematics and Computation, 85(2–3):227–243, September 1, 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300396001385.

Heymann:1996:FDP

[HK96] Valerii I. Heymann and Arkadii V.

Kryazhimskii. On finitedimensional parametrizations of attainability sets. Applied Mathematics and Computation, 78(2–3):137–151, September 1, 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0096300396000045.

Hipel:1997:DSS

[HKFP97] Keith W. Hipel, D. Marc Kilgour, Liping Fang, and Xiaoyong (John) Peng. The decision support system GMCR in environmental conflict management. Applied Mathematics and Computation, 83(2–3):117–152, May 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300396001701.

Hutchinson:1998:HIC

[HKL98] John Hutchinson, Mark J.

Kaiser, and Hamid M. Lankarani. The Head Injury Criterion (HIC) functional. Applied Mathematics and Computation, 96(1):1–16, October 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300397101060.

Hon:1997:MMN

[HLXZ97] Y. C. Hon, M. W. Lu, W. M. Xue, and Y. M. Zhu. Multiquadric method for the numerical solution of a biphasic mixture model. Applied Mathematics and Computation, 88 (2–3):153–175, December 30, 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300396003098.

Hon:1998:ENS

[HM98] Y. C. Hon and X. Z. Mao. An efficient numerical scheme for Burgers’ equation. Applied Mathematics and Computation, 95(1):37–50, September 1, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300397100601.

Hossain:1999:CNG

[HM99] Md. Akram Hossain and A. S. Miah. Crank–Nicolson– Galerkin model for transport in groundwater: Refined criteria for accuracy. Applied Mathematics and Computation, 105(2–3):173–181, November/December 1999. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
sciencedirect.com/science/
article/pii/S0096300398101005.

Hossain:1999:MAD

[Hos99] Md. Akram Hossain. Modeling advective–dispersive transport with reaction: an accurate explicit finite difference model. Applied Mathematics and Computation, 102(2–3):101–108, July 15, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S009630039810036X.

Hartfield:1998:CFI

[HRA98] Roy Hartfield, Steve Rose, and John Abbitt. Computational fluid imaging for iodine fluorescence in compressible flows. Applied Mathematics and Computation, 95(1):63–73, September 1, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S009630039710090X.

Hamdan:1996:DGF

[HS96] M. H. Hamdan and K. D.

Sawalha. Dusty gas flow through porous media. Applied Mathematics and Computation, 75(1):59–73, March 1, 1996. CODEN AMHCBQ.

ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0096300395001042.

Hernandez:1998:CMC

[HS98] M. A. Hernández and M. A. Salanova. Chebyshev method and convexity. Applied Mathematics and Computation, 95(1):51–62, September 1, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300397100911.

Hernandez:1999:ICC

[HS99] M. A. Hernández and M. A. Salanova. Indices of convexity and concavity. Application to Halley method. Applied Mathematics and Computation, 103(1):27–49, August 1, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300398100474.

Hicks:1997:CSS

[HSA97] D. L. Hicks, J. W. Swegle, and S. W. Attaway. Conservative smoothing stabilizes discrete-numerical instabilities in SPH material dynamics computations. Applied Mathematics and Computation, 85 (2–3):209–226, September 1, 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300396001361.

Hamann:1998:DTS

[HT98] Bernd Hamann and Po-Yu Tsai. Decomposing trimmed surfaces using the Voronoĭ diagram and a scan line algorithm. Applied Mathematics and Computation, 89(1– 3):327–344, January/February 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300397816669.

Huang:1999:EES

[Hua99]  Zhengda Huang.  On the

error estimates of several Newton-like methods. Applied Mathematics and Computation, 106(1):1–16, December 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300398100851.

Han:1998:CDN

[HW98] Danfu Han and Xinghua Wang.

Convergence on a deformed Newton method. Applied Mathematics and Computation, 94(1):65–72, August 1, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.

com/science/article/pii/
S0096300397100662.

Hossain:1998:MCT

[HY98] Md. Akram Hossain and David R. Yonge. Modeling contaminant transport in groundwater: an optimized finite element method. Applied Mathematics and Computation, 96(1):89–100, October 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S009630039710128X.

Hossain:1999:SAD

[HY99a] Md. Akram Hossain and D. R. Yonge. Simulating advectivedispersive transport in groundwater: an accurate finite difference model. Applied Mathematics and Computation, 105(2–3): 221–230, November/December 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300398101054.

Hossain:1999:ATG

[HY99b] Md. Akram Hossain and David R. Yonge. Accuracy of the Taylor–Galerkin model for contaminant transport in groundwater. Applied Mathematics and Computation, 102(2–3):109–119, July 15, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300398100383.

Hossain:1999:GMT

[HY99c] Md. Akram Hossain and David R. Yonge. On Galerkin models for transport in groundwater. Applied Mathematics and Computation, 100(2– 3):249–263, May 1999. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
sciencedirect.com/science/
article/pii/S0096300398000253.

Heilman:1997:TFI

[HYL97] Philip Heilman, Diana S.

Yakowitz, and Leonard J. Lane. Targeting farms to improve water quality. Applied Mathematics and Computation, 83(2–3):173–194, May 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300396001683.

Huang:1995:DOF

[HZ95] Y. Huang and N. G. Zamani. Design optimization via finite elements: an elementary account. Applied Mathematics and Computation, 69(2–3):275–298, May 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
009630039400135Q.

Huang:1996:OBD

[HZ96] Y. Huang and N. G. Zamani. Optimum beam design using finite elements. Applied Mathematics and Computation, 80(2–3):155–179, December 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0096300395002855.

Huang:1998:FPR

[HZ98] Zhengda Huang and Shiming Zheng. On a family of parallel root-finding methods for generalized polynomials. Applied Mathematics and Computation, 91(2–3):221–231, May 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300397100224.

Han:1997:WCG

[HZL97] B. Han, M. L. Zhang, and J. Q. Liu. A widely convergent generalized pulse-spectrum technique for the coefficient inverse problem of differential equations. Applied Mathematics and Computation, 81(2–3):97–112, February 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0096300395003045.

Ibrahim:1996:DSV

[Ibr96] A-G. M. Ibrahim. On the differentiability of set-valued functions defined on a Banach space and mean value theorem. Applied Mathematics and Computation, 74(1):79–94, January 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
009630039500095X.

Iwakoshi:1998:FCS

[IFU98] Yasushi Iwakoshi, Takeshi Furuhashi, and Yoshiki Uchikawa. A Fuzzy Classifier System for evolutionary learning of robot behaviors. Applied Mathematics and Computation, 91 (1):73–81, April 1998. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
sciencedirect.com/science/
article/pii/S0096300397100066.

Iannelli:1997:SMN

[IKP97] Mimmo Iannelli, Mi-Young Kim, and Eun-Jae Park. Splitting methods for the numerical approximation of some models of age-structured population dynamics and epidemiology. Applied Mathematics and Computation, 87(1):69–93, November 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300396002226.

Itiki:1996:ICP

[IKU96] Cinthia Itiki, Robert Kalaba, and Firdaus Udwadia. Inequality constraints in the process of jumping. Applied Mathematics and Computation, 78 (2–3):163–173, September 1, 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0096300396000069.

Ivanenko:1995:NMU

[IL95] V. I. Ivanenko and V. A. Labkovsky. On the natural measure of uncertainty and information. Applied Mathematics and Computation, 70(2–3):185–190, July 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        009630039400121J.

Inohara:1999:CMR

[Ino99] Takehiro Inohara. On conditions for a meeting not to reach a recurrent argument. Applied Mathematics and Computation, 101(2–3):281–298, June 15, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S009630039810005X.

Ibidapo-Obe:1995:ACB

[IOAB95] O. Ibidapo-Obe, A. B. Alonge, and Adedeji B. Badiru. On active controls for a biped mechanism. Applied Mathematics and Computation, 69(2–3):159–183, May 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
009630039400090Q.

Inohara:1997:IDC

[ITN97] Takehiro Inohara, Shingo Takahashi, and Bunpei Nakano. Impossibility of deception in a conflict among subjects with interdependent preference. Applied Mathematics and Computation, 81(2–3):221–244, February 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300396000185.

Inohara:1998:CSI

[ITN98a] Takehiro Inohara, Shingo Takahashi, and Bunpei Nakano. Complete stability and inside commonality of perceptions. Applied Mathematics and Computation, 90(1):11–25, March 1, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300396003347.

Inohara:1998:CMR

[ITN98b] Takehiro Inohara, Shingo Takahashi, and Bunpei Nakano. On conditions for a meeting not to reach a deadlock. Applied Mathematics and Computation, 90(1):1–9, March 1, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300396003335.

Jang:1999:SFC

[JA99] Sophia R.-J. Jang and Linda J. S. Allen. A simple food chain with a growth inhibiting nutrient. Applied Mathematics and Computation, 104(2–3):277–298, September 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300398100760.

Jacobson:1995:AQF

[Jac95] Sheldon H. Jacobson. Analyzing the M/M/1 queue in frequency domain experiments. Applied Mathematics and Computation, 69(2–3):185–194, May 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        009630039400125N.

Janssens:1995:EMS

[Jan95] F. Janssens. Exact mode shapes for a continuous tether. Applied Mathematics and Computation, 70(2–3):191–201, July 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
009630039400120S.

Jayasuriya:1996:MAU

[Jay96] Kumara Jayasuriya. Multiprecision arithmetic using fast Hartley transforms. Applied Mathematics and Computation, 75(2–3):239–251, March 15, 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0096300396900673.

Jimenez:1998:ESA

[JBV98] S. Jiménez, S. Bulgakov, and L. Vázquez. Efficient shooting algorithms for solving the nonlinear one-dimensional scalar Helmholtz equation. Applied Mathematics and Computation, 95(2–3):101–114, September 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300397101266.

Jang:1997:ARS

[JC97a] Ming-Jyi Jang and Chieh-Li Chen. Analysis of the response of a strongly nonlinear damped system using a differential transformation technique. Applied Mathematics and Computation, 88 (2–3):137–151, December 30, 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300396003086.

Jonckheere:1997:BFT

[JC97b] Edmond A. Jonckheere and Chung-Kuang Chu. Bounded flatness in Q-triangulated regular N-simplexes. Applied Mathematics and Computation, 88 (2–3):177–198, December 30, 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300396003104.

Jonckheere:1998:SAC

[JCC98] Edmond A. Jonckheere, Chung-Kuang Chu, and Chih-Yung Cheng. Simplicial algorithms for computing stationary probabilities of stochastic matrices. Applied Mathematics and Computation, 93(2–3):207–217, July 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300397100972.

Jenkins:1996:NAU

[JFW96] Rhonald M. Jenkins, Winfred A. Foster, Jr., and Lora S. Wirth. Numerical analysis of unsteady multiple jet plume interactions. Applied Mathematics and Computation, 79(2–3):239–247, October 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0096300395002685.

Jiaqi:1997:NBV

[Jia97] Mo Jiaqi. The nonlocal boundary value problems of nonlinear elliptic systems in unbounded domains. Applied Mathematics and Computation, 86(2–3):115–121, October 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300396001762.

Jiaxing:1999:LCB

[Jia99]  He Jiaxing.  On a linear

combination of S. N. Bernstein trigonometric interpolation polynomial. Applied Mathematics and Computation, 106(2–3):197–203, December 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300398101224.

Jin:1995:FAB

[Jin95] Xiao-Qing Jin. A fast algorithm for block Toeplitz systems with tensor structure. Applied Mathematics and Computation, 73(2–3):115–124, December 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.

com/science/article/pii/
0096300394002436.

Jin:1997:NCC

[Jin97] Xiao-Qing Jin. A note on construction of circulant preconditioners from kernels. Applied Mathematics and Computation, 83(1):3–12, April 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300396003323.

Janicki:1998:WOS

[JK98] Ryszard Janicki and Waldemar W. Koczkodaj. A weak order solution to a group ranking and consistency-driven pairwise comparisons. Applied Mathematics and Computation, 94(2–3):227–241, August 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300397100807.

Jeong:1997:EMA

[JL97] Il-Kwon Jeong and Ju-Jang Lee. Evolving multi-agents using a self-organizing genetic algorithm. Applied Mathematics and Computation, 88 (2–3):293–303, December 30, 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300396003372.

Jukic:1998:DTN

[JMS98] D. Jukić, T. Marosšević, and R. Scitovski. Discrete total l_{p}-norm approximation problem for the exponential function. Applied Mathematics and Computation, 94 (2–3):137–143, August 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300397100686.

Juang:1995:CRI

[JN95] Jonq Juang and Paul Nelson. Convergence rates of iterative solutions of algebraic matrix Riccati equations. Applied Mathematics and Computation, 72(2–3):125–142, October 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0096300394001798.

Jodar:1995:ECS

[JNF95] L. Jódar, E. Navarro, and M. V.

Ferrer. Erratum: a correction on the stability of implicit higher order difference systems. Applied Mathematics and Computation, 71(1):91–96, August 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
009630039500060U. See [NFJ94].

Jodar:1996:CSI

[JNF96] L. Jódar, E. Navarro, and M. V. Ferrer. A correction on the stability of implicit higher order difference systems. Applied Mathematics and Computation, 74(2–3):299–304, February 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0096300395002170.

Jodar:1995:SOD

[JPF95] Lucas Jódar, Enrique Ponsoda, and M. Legua Fernández. Second-order difference systems in unbounded domains: Existence conditions and construction of exact and approximate bounded solutions. Applied Mathematics and Computation, 71(2–3):271–287, September 1, 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        009630039400158Z.

Jayakumar:1995:CMS

[JR95] J. Jayakumar and N. Ramanujam. A computational method for solving quasilinear singular perturbation problems. Applied Mathematics and Computation, 71(1):1–14, August 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.

com/science/article/pii/
009630039400077H.

Jukic:1996:EOP

[JS96] Dragan Jukić and Rudolf Scitovski. The existence of optimal parameters of the generalized logistic function. Applied Mathematics and Computation, 77(2–3):281–294, July 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300395002510.

Jin:1998:ANC

[JS98a] Xiao-Qing Jin and Vai-Kuong Sin. Addendum to “a note on construction of circulant preconditioners from kernels”. Applied Mathematics and Computation, 95(1):91–99, September 1, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300397100753.

Jinzenji:1998:ASS

[JS98b] T. Jinzenji and M. Sasaki. An advanced simulation scheme for electric railway power systems based on artificial life approach. Applied Mathematics and Computation, 91(1):23–32, April 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300397100121.

Jiang:1998:QMS

[JV98] Jiahui Jiang and A. S. Vatsala. The quasilinearization method in the system of reaction diffusion equations. Applied Mathematics and Computation, 97 (2–3):223–235, December 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300397101382.

Jayawardene:1997:ULO

[JY97] Sumedha Jayawardene and Shie-Shien Yang. On using linear ordered rank statistics for detecting early differences between two distributions. Applied Mathematics and Computation, 84(2–3):103–113, July 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300396000811.

Kadalbajoo:1995:BEM

[KA95] Mohan K. Kadalbajoo and K. Arumugam. Boundary element method for elliptic differential equations with a small parameter. Applied Mathematics and Computation, 71 (2–3):151–163, September 1, 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        009630039400151S.

Kumar:1996:AHM

[KA96a] Kuldeep Kumar and M. A. Alsaleh. Application of Hankel matrices in polynomial regression. Applied Mathematics and Computation, 77(2–3):205–211, July 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300395002154.

Kumar:1996:CSE

[KA96b] Kuldeep Kumar and M. A. Alsaleh. A comparative study for the estimation of parameters in nonlinear models. Applied Mathematics and Computation, 77(2–3):179–183, July 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300395002111.

Kananthai:1997:SDD

[Kan97] A. Kananthai. On the solutions of the n-dimensional diamond operator. Applied Mathematics and Computation, 88(1):27–37, December 15, 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300396003037.

Kananthai:1999:FTD

[Kan99]  Amnuay Kananthai.  On

the Fourier transform of the Diamond Kernel of Marcel Riesz. Applied Mathematics and Computation, 101(2–3):151–158, June 15, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300398100012.

Karsai:1998:ABV

[Kar98]  János Karsai.  Asymptotic

behavior and its visualization of the solutions of intermittently and impulsively damped nonlinear oscillator equations. Applied Mathematics and Computation, 89(1– 3):161–172, January/February 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300397816554.

Kazantsev:1999:LLE

[Kaz99] E. Kazantsev. Local Lyapunov exponents of the quasigeostrophic ocean dynamics. Applied Mathematics and Computation, 104(2–3):217– 257, September 1999. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
sciencedirect.com/science/
article/pii/S0096300398100784.

Khan:1998:PIE

[KB98] Shahjahan Khan and M. I. Bhatti. Predictive inference on equicorrelated linear regression models. Applied Mathematics and Computation, 95 (2–3):205–217, September 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S009630039710100X.

Kutluay:1999:VFD

[KBÖ99] S. Kutluay, A. R. Bahadir, and A. Özdeş. A variety of finite difference methods to the thermistor with a new modified electrical conductivity. Applied Mathematics and Computation, 106(2–3):205–213, December 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300398101236.

Katti:1996:CFD

[KBS96] C. P. Katti, S. N. Baboo, and S. Sivaloganathan. On the convergence of finite difference methods for a class of two-point boundary value problems with periodic boundary conditions. Applied Mathematics and Computation, 75(2–3):287–302, March 15, 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0096300396900739.

Keshavamurthy:1995:EWSa

[KD95a] Ramdass Keshavamurthy and Marijan Dravinski. Elastic wave scattering by an inclusion in a multilayered medium submerged in fluid. I. Plane strain model. Applied Mathematics and Computation, 67 (1–3):33–60, January/February 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0096300394000537.

Keshavamurthy:1995:EWSb

[KD95b] Ramdass Keshavamurthy and Marijan Dravinski. Elastic wave scattering by an inclusion in a multilayered medium submerged in fluid. II. Threedimensional model. Applied Mathematics and Computation, 67(1–3):265–292, January/February 1995. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic). URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300394000618.

Kozak:1995:SRS

[KH95] Š. Kozák and J. Hejdiš. Simple, robust, selftuning controller. Applied Mathematics and Computation, 70(2–3):203–214, July 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        009630039400119O.

Kamel:1996:ATF

[KH96] M. T. Kamel and M. H. Hamdan. Aspects of thin-film polar fluid lubrication. Applied Mathematics and Computation, 80(1):33–41, November 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0096300395002782.

Khan:1996:RBP

[Kha96] Shahjahan Khan. Regressionbased prediction for twostage survey data with correlated normal errors. Applied Mathematics and Computation, 79(2–3):105–124, October 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0096300395002456.

Khuri:1998:BCA

[Khu98a] S. A. Khuri. Biorthogonality condition for axisymmetric Stokes flow in a toroidal region. Applied Mathematics and Computation, 97 (2–3):255–259, December 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300397101497.

Khuri:1998:NAC

[Khu98b] S. A. Khuri. A new approach to the cubic Schrödinger equation: an application of the decomposition technique. Applied Mathematics and Computation, 97(2–3):251–254, December 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300397101473.

Khuri:1998:SIE

[Khu98c] S. A. Khuri. On the solution of an integral equation arising in water waves theory. Applied Mathematics and Computation, 95(2–3):291–295, September 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300397100248.

Kim:1996:ESS

[Kim96] Mi-Young Kim. Existence of steady state solutions to an epidemic model with screening and their asymptotic stability. Applied Mathematics and Computation, 74(1):37–58, January 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0096300395000712.

Kincanon:1997:AAR

[Kin97]  Eric Kincanon.  Ambigui-

ties and almost reflectionless potentials. Applied Mathematics and Computation, 83 (1):79–85, April 1997. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
sciencedirect.com/science/
article/pii/S0096300396000483.

Kjellstrom:1999:EB

[Kje99] Gregor Kjellström. The evolution in the brain. Applied Mathematics and Computation, 98(2–3):293–300, February 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300397101564.

Kagiwada:1995:FEF

[KK95] Harriet H. Kagiwada and Robert E. Kalaba. Fuzzy evidential filter for detection and tracking of dim objects. Applied Mathematics and Computation, 69(1):75–96, April 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
009630039400100I.

Kawamata:1995:SDT

[KKH95] Masayuki Kawamata, Eiichiro Kawakami, and Tatsuo Higuchi. Separable-denominator twodimensional adaptive filters with application to noise reduction in images. Applied Mathematics and Computation, 69(1):111–122, April 1995. CODEN AMHCBQ.

ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
009630039400102A.

Kwak:1999:MAC

[KKL99] Do Y. Kwak, Hyun J. Kwon, and Sungyun Lee. Multigrid algorithm for cell centered finite difference on triangular meshes. Applied Mathematics and Computation, 105(1):77–85, October 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300398100942.

Kappeller:1996:OEM

[KKPL96] M. Kappeller, M. Kiehl, M. Perzl, and M. Lenke. Optimized extrapolation methods for parallel solution of IVPs on different computer architectures. Applied Mathematics and Computation, 77(2–3):301–315, July 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300395002197.

Kang:1999:NIT

[KL99a] Kab Seok Kang and Sung Yun Lee. New intergrid transfer operator in multigrid method for P _{1}-nonconforming finite element method. Applied Mathematics and Computation, 100(2–3):139–149, May 1999. CODEN AMHCBQ.

ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300398000332.

Kim:1999:PGS

[KL99b] Taegil Kim and Chang-Ock Lee. A parallel Gauss– Seidel method using NR data flow ordering. Applied Mathematics and Computation, 99(2–3):209–220, March 15, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300398000083.

Kalachev:1995:OSS

[KM95] L. V. Kalachev and R. M. M.

Mattheij. On optimally scaled systems for second-order scalar singularly perturbed problems. Applied Mathematics and Computation, 68(1):71–93, March 1, 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
009630039400091H.

Kolowrocki:1995:ARF

[Koł95] Krzysztof Kołowrocki. Asymptotic reliability functions of some nonhomogeneous seriesparallel and parallel-series systems. Applied Mathematics and Computation, 73(2–3):133–151, December 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0096300394002452.

Koreas:1997:NCC

[Kor97] Diamantis P. Koreas. The NPcompleteness of chromatic index in triangle free graphs with maximum vertex of degree 3. Applied Mathematics and Computation, 83(1):13–17, April 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300396000215.

Kozera:1995:CIU

[Koz95] Ryszard Kozera. On complete integrals and uniqueness in shape from shading. Applied Mathematics and Computation, 73(1):1–37, November 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        009630039400127P.

Kreuzer:1996:CPR

[KP96] Edwin Kreuzer and Fernando C. Pinto. Controlling the position of a remotely operated underwater vehicle. Applied Mathematics and Computation, 78(2–3):175–185, September 1, 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.

com/science/article/pii/
0096300396000070.

Kim:1998:CFE

[KP98] Mi-Young Kim and Eun-Jae Park. Characteristic finite element methods for diffusion epidemic models with agestructured populations. Applied Mathematics and Computation, 97(1):55–70, December 1, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300397101643.

Kadalbajoo:1995:AGE

[KR95a] Mohan K. Kadalbajoo and A. Appaji Rao. The Alternating Group Explicit (AGE) method for singularly perturbed boundary value problems. Applied Mathematics and Computation, 68(2–3):125–142, March 15, 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
009630039400087K.

Kaye:1995:SBF

[KR95b] M. Kaye and D. Reuster. Subdomain basis functions from entire domain basis functions for a certain class of problems. Applied Mathematics and Computation, 72(2–3):167–173, October 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        0096300394001813.

Kuhn:1995:CVM

[KSZ^{+}95] A. Kuhn, W. Steiner, J. Zemann, D. Dinevski, and H. Troger. A comparison of various mathematical formulations and numerical solution methods for the large amplitude oscillations of a string pendulum. Applied Mathematics and Computation, 67(1– 3):227–264, January/February 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        009630039400060H.

Kamont:1996:DDI

[KTZK96] Z. Kamont, J. Turo, and B. Zubik-Kowal. Differential and difference inequalities generated by mixed problems for hyperbolic functional differential equations with impulses. Applied Mathematics and Computation, 80(2–3):127–154, December 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0096300395002847.

Kubica:1997:EAC

[Kub97] Krystian Kubica. The effect of amphiphilic counterions on the gel-fluid phase transition of the lipid bilayer. Applied Mathematics and Computation, 87(2–3):261–270, December 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300396003013.

Kumar:1995:ETD

[Kum95] Ravi Kumar. Effects of time delays on the stability of collocated and noncollocated point control of discrete dynamic structural systems. Applied Mathematics and Computation, 70(2–3):215–232, July 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
009630039400107F.

Kawata:1995:SAR

[KUO95] Y. Kawata, S. Ueno, and A. Ohtani. The surface albedo retrieval of mountainous forest area from satellite MSS data. Applied Mathematics and Computation, 69(1):41–59, April 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
009630039400098O.

Kurowicka:1999:DAA

[Kur99] Dorota Kurowicka. Domains of attraction of asymptotic reliability functions of some homogeneous series-parallel systems. Applied Mathematics and Computation, 98(1):61–74, January 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300397101321.

Kalaba:1995:CMR

[KUX95] Robert Kalaba, Firdaus Udwadia, and Rong Xu. Constrained motion revisited. Applied Mathematics and Computation, 70(1):67–76, June 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        009630039400143R.

Kalla:1998:SFO

[KVT98] Shyam L. Kalla, N. Virchenko, and V. Tsarenko. On some fractional order integral transforms generated by orthogonal polynomials. Applied Mathematics and Computation, 91(2–3):209–219, May 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300397100194.

Khuri:1996:SPD

[KW96] S. A. Khuri and A. M. Wazwaz. The solution of a partial differential equation arising in fluid flow theory. Applied Mathematics and Computation, 77(2–3):295–300, July 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S009630039500193X.

Khuri:1997:SPD

[KW97] S. A. Khuri and A. M. Wazwaz.

On the solution of a partial differential equation arising in Stokes flow. Applied Mathematics and Computation, 85 (2–3):139–147, September 1, 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300396001312.

Kwak:1996:CAM

[Kwa96] Do Y. Kwak. A p-cycle analysis of multigrid method. Applied Mathematics and Computation, 76(1):41–48, April 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0096300395001484.

Kwak:1997:PGM

[Kwa97] Do Y. Kwak. A preconditioned GMRES method. Applied Mathematics and Computation, 85(2–3):201–208, September 1, 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S009630039600135X.

Kang:1995:FSP

[KWS95] Long-Chyuan Kang, Cheng-Hsiu Wu, and Charles R. Steele. Fourier series for polygonal plate bending: a very large plate element. Applied Mathematics and Computation, 67(1– 3):197–225, January/February 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        009630039400059D.

Lam:1998:LDC

[Lam98] S. H. Lam. On Lagrangian dynamics and its control formulations. Applied Mathematics and Computation, 91(2–3):259–284, May 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300397100042.

Lameche:1999:ISA

[Lam99] Khira Lameche. Inverse for the shuffle for algebraic series. Applied Mathematics and Computation, 98(1):1–27, January 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300397100595.

Larsen:1997:SCS

[Lar97]  Jens Chr. Larsen.  Stable

chemical systems. Applied Mathematics and Computation, 81(2–3):245–257, February 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300396000203.

Lybeck:1996:SMD

[LB96] Nancy J. Lybeck and Kenneth L. Bowers. Sinc methods for domain decomposition. Applied Mathematics and Computation, 75(1):13–41, March 1, 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0096300395000992.

Li:1997:PIF

[LC97] Likang Li and Jinru Chen.

Preconditioning isoparametric finite element methods taking into account numerical integration. Applied Mathematics and Computation, 87(2–3):271–288, December 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300396003025.

Le:1998:SGB

[Le98] Vy Khoi Le. Some global bifurcation results for elastic plates. Applied Mathematics and Computation, 89(1– 3):185–197, January/February 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300397816578.

Lee:1998:CMF

[Lee98] Chang-Ock Lee. A conforming mixed finite element method for the pure traction problem of linear elasticity. Applied Mathematics and Computation, 93(1):11–29, July 1, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S009630039710073X.

Leitmann:1995:OAC

[Lei95] G. Leitmann. One approach to the control of uncertain dynamical systems. Applied Mathematics and Computation, 70(2–3):261–272, July 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        009630039400118N.

Lether:1997:CNM

[Let97]  Frank G. Lether.  Con-

strained near-minimax rational approximations to Dawson’s integral. Applied Mathematics and Computation, 88 (2–3):267–274, December 30, 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.

com/science/article/pii/
S009630039600330X.

Liu:1999:HON

[LF99] Xinzhi Liu and Xilin Fu.

High order nonlinear differential inequalities with distributed deviating arguments and applications. Applied Mathematics and Computation, 98(2–3):147–167, February 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300397101370.

Lee:1995:PEN

[LFS95] F. C. Lee, H. Flashner, and M. G. Safonov. Positivity embedding for noncolocated and nonsquare flexible systems. Applied Mathematics and Computation, 70(2–3):233–246, July 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
009630039400108G.

Liu:1997:MCP

[LG97] Pingzhou Liu and K. Gopalsamy. On a model of competition in periodic environments. Applied Mathematics and Computation, 82(2–3):207–238, March 15, 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        S0096300396000446.

Liu:1999:GSC

[LG99] Pingzhou Liu and K. Gopalsamy. Global stability and chaos in a population model with piecewise constant arguments. Applied Mathematics and Computation, 101(1): 63–88, June 1, 1999. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic). URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/S009630039800037X.

Li:1998:SAR

[Li98] Xin Li. On simultaneous approximations by radial basis function neural networks. Applied Mathematics and Computation, 95(1):75–89, September 1, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300397100893.

Liao:1997:SEA

[Lia97]  Aiping Liao.  Some effi-

cient algorithms for unconstrained discrete-time optimal control problems. Applied Mathematics and Computation, 87(2–3):175–198, December 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300396002287.

Liaw:1998:ACM

[Lia98] Der-Cherng Liaw. Application of center manifold reduction to nonlinear system stabilization. Applied Mathematics and Computation, 91(2–3):243–258, May 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300397100212.

Lin:1996:ECO

[Lin96] Cheng-Chang Lin. An existence condition for open set nonlinear complementarity problems with an application to a closed economic system. Applied Mathematics and Computation, 76(2–3):285–292, May 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0096300395001859.

Lin:1998:FEM

[Lin98] Tao Lin. A C^{0} finite element method for an inverse problem. Applied Mathematics and Computation, 90(2–3):253–284, March 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300397004013.

Liu:1995:ICO

[Liu95] Xinzhi Liu. Impulsive control and optimization. Applied Mathematics and Computation, 73(1):77–98, November 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
009630039400204H.

Lubarda:1995:CSR

[LK95] V. A. Lubarda and D. Krajcinovic. Constitutive structure of rate theory of damage in brittle elastic solids. Applied Mathematics and Computation, 67(1– 3):81–101, January/February 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0096300394000559.

Layton:1995:TLP

[LL95] W. Layton and W. Lenferink. Two-level Picard and modified Picard methods for the Navier–Stokes equations. Applied Mathematics and Computation, 69(2–3):263–274, May 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
009630039400134P.

Liao:1996:RIS

[LL96] Xiao-Xin Liao and Jia Li. Robust interval stability, persistence, and partial stability on Lotka–Volterra systems with time-delay. Applied Mathematics and Computation, 75(2–3):103–115, March 15, 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300396900491.

Lin:1997:MNC

[LLL97] R. Lin, G. Leng, and H. P.

Lee. A method for the numerical computation of Hopf bifurcation. Applied Mathematics and Computation, 86(2–3): 137–156, October 1997. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
sciencedirect.com/science/
article/pii/S0096300396001786.

Luo:1996:RTC

[LLX96] Fa-Long Luo, Yan-Da Li, and Hong-Qinq Xu. Real-time computation of the eigenvectors of a class of positive definite matrices. Applied Mathematics and Computation, 77(2–3):267–280, July 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300395002200.

Lee:1995:VIE

[LM95] Jungki Lee and Ajit K. Mal.

A volume integral equation technique for multiple scattering problems in elastodynamics. Applied Mathematics and Computation, 67(1– 3):135–159, January/February 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        009630039400057B.

Lopez:1998:NMI

[LM98] David J. López and Pablo Martín. A numerical method for the integration of perturbed linear problems. Applied Mathematics and Computation, 96(1):65–73, October 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300397101217.

Lakshmikantham:1997:SLP

[LMSS97] V. Lakshmikantham, A. K. Maulloo, S. K. Sen, and S. Sivasundaram. Solving linear programming problems exactly. Applied Mathematics and Computation, 81(1):69–87, January 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0096300395003096.

Lee:1999:RPS

[LO99] K. H. Lee and E. H. Ong. A reduction principle for singular perturbation problems. Applied Mathematics and Computation, 101(1):45–62, June 1, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300398000423.

Linzhang:1997:MSF

[LP97] Lu Linzhang and C. E. M.

Pearce. On the matrix-sign-function method for solving algebraic Riccati equations. Applied Mathematics and Computation, 86(2–3):157–170, October 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300396001798.

Leitmann:1995:CSB

[LR95] G. Leitmann and E. Reithmeier. A control scheme based on ER-materials for vibration attenuation of dynamical systems. Applied Mathematics and Computation, 70(2–3):247–259, July 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
009630039400109H.

Ladde:1995:SAH

[LS95] G. S. Ladde and S. Sathananthan. Stability analysis of hereditary iterative processes. Applied Mathematics and Computation, 73(1):39–54, November 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
009630039400202F.

Lin:1997:TK

[LS97] Li Jian Lin and H. M. Srivastava. On a theorem of Kobori. Applied Mathematics and Computation, 85 (2–3):287–296, September 1, 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300396001439.

Liu:1998:GCT

[LS98a] Jinglian J. Liu and Bharat K. Soni. 2D groundwater contaminant transport modeling by using the finite volume method on an unstructured grid system. Applied Mathematics and Computation, 89(1– 3):199–211, January/February 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S009630039781658X.

Lu:1998:NSR

[LS98b] Xiaowu Lu and Rudolf Schmid. A numerical study of the Riemann solutions for gasdynamic combustion. Applied Mathematics and Computation, 91(2–3):143–160, May 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300397100157.

Li:1999:ERS

[LS99] Weiye Li and Ferenc Szidarovszky. An elementary result in the stability theory of time-invariant nonlinear discrete dynamical systems. Applied Mathematics and Computation, 102(1):35–49, July 1, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300398100139.

Lehtokangas:1996:NAP

[LSKH96] Mikko Lehtokangas, Jukka Saarinen, Kimmo Kaski, and Pentti Huuhtanen. A network of autoregressive processing units for time series modeling. Applied Mathematics and Computation, 75(2–3):151–165, March 15, 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0096300396900570.

Lin:1998:DNT

[LSY98] C. Lin, F. Szidarovszky, and J. Yen. Dynamic negotiation with time-varying Pareto frontier. Applied Mathematics and Computation, 91(2–3):99–109, May 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300397100017.

Lu:1998:CIM

[Lu98] Xin Lu. Combined iterative methods for numerical solutions of parabolic problems with time delays. Applied Mathematics and Computation, 89(1– 3):213–224, January/February 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300397816591.

Lui:1999:MBS

[Lui99] S. H. Lui. Multiple bifurcation from “simple” eigenvalues. Applied Mathematics and Computation, 100(2–3):111–130, May 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300398000162.

Luo:1997:RTC

[LUL97] Fa-Long Luo, Rolf Unbehauen, and Yan-Da Li. Real-time computation of singular vectors. Applied Mathematics and Computation, 86(2–3):197– 214, October 1997. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic). URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/S0096300396001828.

Layton:1998:NTL

[LY98] W. Layton and X. Ye. Nonconforming two-level discretization of stream function form of the Navier–Stokes equations. Applied Mathematics and Computation, 89(1– 3):173–183, January/February 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300397816566.

Lin:1997:CHI

[LYF97] Zhenghua Lin, Bo Yu, and Guochen Feng. A combined homotopy interior point method for convex nonlinear programming. Applied Mathematics and Computation, 84(2–3):193–211, July 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300396000860.

Logan:1996:TPT

[LZ96] J. David Logan and Vitaly Zlotnik. Time-periodic transport in heterogeneous porous media. Applied Mathematics and Computation, 75(2–3):119–138, March 15, 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0096300396900533.

Marchuk:1995:ROC

[MA95] G. I. Marchuk and V. I.

Agoshkov. Reflection operators and contemporary applications to radiative transfer. Applied Mathematics and Computation, 69(1):3–21, April 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        009630039400149X.

Martin:1995:UMS

[MAS95] Clyde F. Martin, Linda J. S. Allen, and Mark Stamp. Urn model simulations of a sexually transmitted disease epidemic. Applied Mathematics and Computation, 71 (2–3):179–199, September 1, 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        009630039400153U.

Moss:1995:DIV

[MB95] Darrell G. Moss and Haym Benaroya. A discrete inverse vibration problem with parameter uncertainty. Applied Mathematics and Computation, 69(2–3):313–333, May 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        009630039400140Y.

Masri:1995:ASO

[MBCvdV95] S. F. Masri, G. A. Bekey, T. K. Caughey, and E. van de Velde. Adaptive stochastic optimization using multiprocessors. Applied Mathematics and Computation, 72(2–3):225–257, October 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0096300394001868.

Matyasovszky:1995:IGC

[MBG95] Istvan Matyasovszky, Istvan Bogardi, and Jacques Ganoulis. Impact of global climate change on temperature and precipitation in Greece. Applied Mathematics and Computation, 71 (2–3):119–150, September 1, 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
009630039400148W.

Monaco:1997:NSB

[MBS97] R. Monaco, M. Pandolfi Bianchi, and A. J. Soares. Numerical simulations of a Boltzmann model for reacting gases. Applied Mathematics and Computation, 85(1):61–85, August 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300396001130.

Mueller:1998:CMM

[MBS98] K. M. Mueller, S. A. Burns, and M. A. Savageau. A comparison of the monomial method and the S-system method for solving systems of algebraic equations. Applied Mathematics and Computation, 90(2–3):167–180, March 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300397003974.

McGough:1998:NCG

[McG98] Jeff S. McGough. Numerical continuation and the Gelfand problem. Applied Mathematics and Computation, 89(1– 3):225–239, January/February 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300397816608.

Malik:1998:VAC

[MD98] Moinuddin Malik and Hao Huy Dang. Vibration analysis of continuous systems by differential transformation. Applied Mathematics and Computation, 96(1):17–26, October 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300397100765.

Martin:1999:INI

[MF99] Pablo Martin and José-Miguel Farto. Improved numerical integration of perturbed oscillators via average. Applied Mathematics and Computation, 99(2–3):129–139, March 15, 1999. CODEN AMHCBQ.

ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300397101813.

Moskowitz:1997:PDB

[MFGW97] Simon Moskowitz, Emmanuel Fernández-Gaucherand, and Michael Whalen. Passagedetector-based traffic queue estimation in Intelligent Transportation Systems: a computational study of competing algorithms. Applied Mathematics and Computation, 86(2–3):93–113, October 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300396001750.

Milito:1995:ADQ

[Mil95] Rodolfo A. Milito. On arrival driven queueing models: Admission control, traffic policing, abandonments, and correlated arrivals. Applied Mathematics and Computation, 70(1):77–94, June 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
009630039400144S.

Miller:1998:OSP

[Mil98] Robert E. Miller. Optimal sensor placement via Gaussian quadrature. Applied Mathematics and Computation, 97(1):71–97, December 1, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300397101205.

Mitra:1997:GFV

[Mit97] Arun K. Mitra. On a geometry of the flat violin boundary. Applied Mathematics and Computation, 81(2–3):201–206, February 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0096300395003304.

Mitra:1999:VA

[Mit99] Arun K. Mitra. The violin arcs. Applied Mathematics and Computation, 98(1):91–102, January 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300397101394.

Morlet:1999:CSO

[MLB99] Anne C. Morlet, Nancy J. Lybeck, and Kenneth L. Bowers. Convergence of the sinc overlapping domain decomposition method. Applied Mathematics and Computation, 98(2–3):209–227, February 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300397101680.

McDonough:1998:DFP

[MMC98] J. M. McDonough, S. Mukerji, and S. Chung. A datafitting procedure for chaotic time series. Applied Mathematics and Computation, 95 (2–3):219–243, September 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300397100625.

Maier:1996:MLD

[MMMK96] Stefan Maier, Mario E. Magaña, Ronald Mohler, and Stephen Kaattari. Modified limiting dilution assay for cell systems. Applied Mathematics and Computation, 78 (2–3):187–195, September 1, 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0096300396000082.

Mastrangelo:1999:NGD

[MMT99] Michèle Mastrangelo, Victor Mastrangelo, and Jean-Marie Teuler. Non-Gaussian distributions. Applied Mathematics and Computation, 101(2–3):99–124, June 15, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300398100036.

Matsuzuka:1999:PTS

[MNK99] Isamu Matsuzuka, Koji Nagasawa, and Akihiro Kitahama. A proposal for two-sided Laplace transforms and its application to electronic circuits. Applied Mathematics and Computation, 100(1):1–11, April 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300398000198.

Mohr:1999:FEM

[Moh99a]  G. A. Mohr.  Finite el-

ement modelling of distribution problems. Applied Mathematics and Computation, 105(1):69–76, October 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300398100978.

Mohr:1999:NPI

[Moh99b] G. A. Mohr. Numerical procedures for input–output analysis. Applied Mathematics and Computation, 101(1):89–98, June 1, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300398000356.

Mohr:1999:TSM

[Moh99c] G. A. Mohr. Time stepping of macroeconomic models. Applied Mathematics and Computation, 102(2–3):273–278, July 15, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300398100449.

Miralles:1997:FGM

[MOP^{+}97] Juan Peña Miralles, Pedro José Jiménez Olivo, Damián Ginestar Peiró, Gumersindo Verdú Martín, and JoséLuis Muñoz-Cobo González. A fast Galerkin method to obtain the periodic solutions of a nonlinear oscillator. Applied Mathematics and Computation, 86(2–3):261–282, October 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300396001932.

Morhac:1995:ASH

[Mor95] M. Morháč. An algorithm to solve Hilbert systems of linear equations precisely. Applied Mathematics and Computation, 73(2–3):209–229, December 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0096300395000615.

Morchalo:1998:SAN

[Mor98] Jarosław Morchało. Stability analysis of nonlinear difference systems. Applied Mathematics and Computation, 95(1):27–36, September 1, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300397100923.

Moustafa:1995:CGM

[Mou95] Magdi. S. Moustafa. The conjugate gradient method for queueing networks. Applied Mathematics and Computation, 71 (2–3):289–294, September 1, 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0096300394001592.

Moustafa:1996:OAP

[Mou96]  Magdi S. Moustafa.  Op-

timal assignment policy of a single server attended by two queues. Applied Mathematics and Computation, 80(2–3):245–255, December 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0096300395002987.

Michalland:1997:BOD

[MPD97] B. Michalland, E. Parent, and L. Duckstein. Biobjective dynamic programming for trading off hydropower and irrigation. Applied Mathematics and Computation, 88(1):53–76, December 15, 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300396003062.

Murty:1999:TPB

[MR99] K. N. Murty and C. V.

Rao. Two-point boundary value problems associated with first order nonlinear difference system — Existence and uniqueness. Applied Mathematics and Computation, 100(2–3):177–188, May 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300398000216.

Morera:1995:ANI

[MRJ95] J. L. Morera, G. Rubio, and L. Jódar. Accurate numerical integration of stiff differential Riccati equations. Applied Mathematics and Computation, 72(2–3):183–203, October 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0096300394001835.

Miller:1995:FRC

[MS95a] A. R. Miller and H. M. Srivastava. Further reducible cases of certain Kampé de Fériet functions associated with incomplete integrals of cylindrical functions. Applied Mathematics and Computation, 68(2–3):199–216, March 15, 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        009630039400094K.

Morrice:1995:AMS

[MS95b] Douglas J. Morrice and Mark Song. An approximation method for a single server queue with cyclically indexed arrival and service rates. Applied Mathematics and Computation, 71 (2–3):247–269, September 1, 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        009630039400157Y.

Moghadam:1997:MFA

[MS97] M. Mohseni Moghadam and R. Sabeti. Markov finite approximation of Frobenius– Perron operator with doublystochastic tridiagonal action matrix. Applied Mathematics and Computation, 85 (2–3):265–280, September 1, 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300396001403.

Malakooti:1999:GPD

[MS99] Behnam Malakooti and Sriram Subramanian. Generalized polynomial decomposable multiple attribute utility functions for ranking and rating multiple criteria discrete alternatives. Applied Mathematics and Computation, 106(1):69–102, December 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300398101133.

Montaldi:1996:HIC

[MSZ96] E. Montaldi, H. M. Srivastava, and G. Zucchelli. Hypergeometric identities of Cayley and Orr types: Some alternative approaches. Applied Mathematics and Computation, 76(1):1–39, April 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
009630039500145X.

Muraskin:1996:ESN

[Mur96] Murray Muraskin. Equations satisfied by a nine parameter subsystem obtained from mathematical aesthetic principles. Applied Mathematics and Computation, 80(1):63–72, November 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0096300395002812.

Muses:1997:DFA

[Mus97a]  C. Musès.  The dimen-

sional family approach in (hyper)sphere packing: a typological study of new patterns, structures, and interdimensional functions. Applied Mathematics and Computation, 88(1):1–26, December 15, 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300397000040. See errata

[Ano97i, Mus98].

Muses:1997:SPB

[Mus97b] C. Musès. A sphere-packing breakthrough via dimensional families. Applied Mathematics and Computation, 83 (1):1–2, April 1997. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
sciencedirect.com/science/
article/pii/S0096300396003311.

Muses:1998:EDF

[Mus98] C. Musès. Erratum: “The dimensional family approach in (hyper)sphere packing: a typological study for new patterns, structures, and interdimensional functions”: Appl. Math. Comput. 88 (1997). Applied Mathematics and Computation, 97(2–3):295, December 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/

S0096300398000307.  See

[Mus97a, Ano97i].

Mollers:1998:GAI

[MvL98] Th. Möllers and O. van Laak.

On the global approximation and interpolation of locally described real valued functions. Applied Mathematics and Computation, 93(1):1–10, July 1, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300397100522.

Martin:1996:QSG

[MW96] C. F. Martin and D. I. Wallace. Quadrature schemes for a general family of functions. Applied Mathematics and Computation, 77(1):1–8, June 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0096300395000690.

Mckay:1998:SSL

[MW98] Steven M. Mckay and Cecilia L. Weingartner. Solution of the Sharpe–Lotka population model via a modified method of characteristics. Applied Mathematics and Computation, 91(2–3):161–177, May 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300397100169.

Malakooti:1998:APF

[MZ98] Behnam Malakooti and YingQing Zhou. Approximating polynomial functions by Feedforward Artificial Neural Networks: Capacity analysis and design. Applied Mathematics and Computation, 90(1):27–51, March 1, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300396003384.

Mohler:1996:NCA

[MZV96] Ronald R. Mohler, Radoslaw R. Zakrzewski, and Rajkumar Vedam. Nonlinear control algorithms and power system application. Applied Mathematics and Computation, 78 (2–3):197–207, September 1, 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0096300396000094.

Narayaninsamy:1999:FID

[Nar99a] Tony Narayaninsamy. Fractional iterates for n-dimensional maps. Applied Mathematics and Computation, 98(2–3):261– 278, February 1999. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic). URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/S0096300397101825.

Narayaninsamy:1999:BB

[Nar99b]  Tony Narayaninsamy.  On

basin boundaries. Applied Mathematics and Computation, 99(2–3):261–274, March 15, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300397101849.

Navarro:1994:CFG

[NFJ94] E. Navarro, M. V. Ferrer, and L. Jódar. Closed form general solution of nonhomogeneous implicit higher order difference systems. Applied Mathematics and Computation, 60(2–3):113–123, February 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0096300394900981. See erra-

tum [JNF95].

He:1997:CSF

[nHyG97] Song nian He and Ben yu Guo.

Chebyshev spectral-finite element method for threedimensional unsteady Navier– Stokes equation. Applied Mathematics and Computation, 82(2–3):97–127, March 15, 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300396000239.

Nedialkov:1999:VSI

[NJC99] N. S. Nedialkov, K. R. Jackson, and G. F. Corliss. Validated solutions of initial value problems for ordinary differential equations. Applied Mathematics and Computation, 105(1):21–68, October 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300398100838.

Ng:1995:CPW

[NLC95] Michael K. Ng, Fu-Rong Lin, and Raymond H. Chan. Construction of preconditioners for Wiener–Hopf equations by operator splitting. Applied Mathematics and Computation, 72(1):77–96, September 15, 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0096300394001787.

Natesan:1998:CMS

[NR98] S. Natesan and N. Ramanujam. A computational method for solving singularly perturbed turning point problems exhibiting twin boundary layers. Applied Mathematics and Computation, 93(2–3):259–275, July 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.

com/science/article/pii/
S009630039710056X.

Natesan:1999:BMS

[NR99a] S. Natesan and N. Ramanujam.

A “Booster method” for singular perturbation problems arising in chemical reactor theory. Applied Mathematics and Computation, 100(1):27–48, April 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300398000149.

Natesan:1999:INS

[NR99b] S. Natesan and N. Ramanujam. Improvement of numerical solution of selfadjoint singular perturbation problems by incorporation of asymptotic approximations. Applied Mathematics and Computation, 98(2–3):119–137, February 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300397101679.

Navarro:1996:FPS

[NS96] Ramón Navarro and José Sarabia. Focal points and secondorder linear differential equations. Applied Mathematics and Computation, 79(2–3): 203–206, October 1996. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0096300395002669.

Nelson:1997:IET

[NS97] Paul Nelson and Daniel L. Seth. Integrodifferential equations for the two-dimensional transition kernels of invariant imbedding. Applied Mathematics and Computation, 82(1):67–83, March 1, 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0096300395003312.

Obeyesekere:1997:MWU

[OA97] Mandri Obeyesekere and Ronald M. Anderson. A model for water uptake in plants. Applied Mathematics and Computation, 84(2–3):163–191, July 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300396000859.

Ozelkan:1998:MOF

[ÖGDB98] Ertunga C. Özelkan, Ágnes Galambosi, Lucien Duckstein, and András Bárdossy. A multi-objective fuzzy classification of large scale atmospheric circulation patterns for precipitation modeling. Applied Mathematics and Computation, 91(2–3):127–142, May 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300397100029.

Ohmori:1998:NST

[Ohm98] Katsushi Ohmori. Numerical solution of two-fluid flows using finite element method. Applied Mathematics and Computation, 92(2–3):125–133, June 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300397100364.

ORegan:1999:PAP

[OM99] D. O’Regan and M. Meehan. Periodic and almost periodic solutions of integral equations. Applied Mathematics and Computation, 105(2–3): 121–136, November/December 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300398100954.

OReilly:1995:RDS

[OMN95] Oliver M. O’Reilly, Naresh K.

Malhotra, and N. Sri Namachchivaya. Reversible dynamical systems: Dissipationinduced destabilization and follower forces. Applied Mathematics and Computation, 70(2–3):273–282, July 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        009630039400117M.

Onah:1997:DMD

[Ona97] E. S. Onah. On direct methods for the discretization of a heat-conduction equation using spline functions. Applied Mathematics and Computation, 85(1):87–96, August 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300396001142.

Okamoto:1998:NNO

[ONOT98] Masahiro Okamoto, Taisuke Nonaka, Shuichiro Ochiai, and Daisuke Tominaga. Nonlinear numerical optimization with use of a hybrid Genetic Algorithm incorporating the Modified Powell method. Applied Mathematics and Computation, 91(1):63–72, April 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300397100078.

Onah:1999:STD

[OU99] S. E. Onah and O. O. Ugbebor. Solution of a two-dimensional stochastic investment problem. Applied Mathematics and Computation, 98(1):75–82, January 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.

com/science/article/pii/
S0096300397101552.

Papageorgiou:1995:MCN

[Pap95] Nikolaos S. Papageorgiou. Minimax control of nonlinear evolution equations. Applied Mathematics and Computation, 68(2–3):217–236, March 15, 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
009630039400095L.

Papamarkos:1996:MFA

[Pap96] Nikos Papamarkos. Multirational function approximation via linear programming. Applied Mathematics and Computation, 75(1):75–89, March 1, 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0096300395001190.

Pardo:1995:SAU

[Par95]  J. A. Pardo.  Some ap-

plications of the useful mutual information. Applied Mathematics and Computation, 72(1):33–50, September 15, 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
009630039400162W.

Pavani:1995:NAR

[Pav95]  R. Pavani.  A numerical

approximation of the rotation number. Applied Mathematics and Computation, 73(2–3):191–201, December 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0096300394002495.

Peiguang:1995:IMB

[Pei95]  Wang Peiguang.  Iterative

methods for the boundary value problem of a fourth-order differential-difference equation. Applied Mathematics and Computation, 73(2–3):257–270, December 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0096300395000666.

Peiguang:1996:BVP

[Pei96] Wang Peiguang. Boundary value problem of a thirdorder mixed type differentialdifference equation. Applied Mathematics and Computation, 80(2–3):273–286, December 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0096300395003037.

Peiguang:1999:OCN

[Pei99] Wang Peiguang. Oscillatory criteria of nonlinear hyperbolic equations with continuous deviating arguments. Applied Mathematics and Computation, 106(2–3):163–169, December 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300398101108.

Pervozvanski:1996:SAL

[Per96] A. Pervozvanski. On some algorithms of learning control. Applied Mathematics and Computation, 78 (2–3):209–216, September 1, 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0096300396000100.

Papageorgiou:1999:UER

[PF99] G. Papageorgiou and I. Th.

Famelis. On using explicit Runge–Kutta–Nyström methods for the treatment of retarded differential equations with periodic solutions. Applied Mathematics and Computation, 102(1):63–76, July 1, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300398100206.

Peng:1999:COB

[PGHX99] Mingshu Peng, Weigao Ge, Lihong Huang, and Qianli Xu.

A correction on the oscillatory behavior of solutions of certain second order nonlinear differential equations. Applied Mathematics and Computation, 104(2–3):207–215, September 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300398100735.

Pottmann:1995:RBF

[PJ95] Martin Pottmann and H. Peter Jörgl. Radial basis function networks for internal model control. Applied Mathematics and Computation, 70(2–3):283–298, July 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        009630039400116L.

Primak:1998:MCV

[PK98] M. E. Primak and B. L. Kheyfets. Methods of centers for variational inequalities and linear programming. Applied Mathematics and Computation, 96(2–3):273–293, November 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300397101357.

Pitanguy:1996:DMA

[PLPW96] I. Pitanguy, F. Leta, D. Pamplona, and H. I. Weber. Defining and measuring aging parameters. Applied Mathematics and Computation, 78 (2–3):217–227, September 1, 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0096300396000161.

Prathapar:1997:SOH

[PMMA97] S. A. Prathapar, W. S. Meyer, J. C. Madden, and E. Alociljá. SWAGMAN Options: a hierarchical multicriteria framework to identify profitable land uses that minimize water table rise and salinization. Applied Mathematics and Computation, 83(2–3):217–240, May 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300396001865.

Parhi:1997:GPA

[PN97a] S. Parhi and G. Nath. The growth of perturbations associated with the linear stability of inviscid fluid flows. Applied Mathematics and Computation, 81(1):23–30, January 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0096300395002995.

Parton:1997:NPM

[PN97b] Kevin A. Parton and Ayut Nissapa. A nonlinear programming model for analyzing aquaculture policy decision making in southern Thailand. Applied Mathematics and Computation, 83(2–3):241–260, May 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300396001877.

Primak:1996:PCP

[PS96] Mordukh E. Primak and Daniel B. Szyld. A projection cutting plane algorithm for convex programming problems. Applied Mathematics and Computation, 74(2–3):261–271, February 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0096300395000976.

Primak:1997:SVI

[PZ97] M. E. Primak and A. Zeleke. Set variational inequalities. Applied Mathematics and Computation, 85(2–3):165–180, September 1, 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic). URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/S0096300396001336.

Piccolomini:1999:CGR

[PZ99] Elena Loli Piccolomini and Fabiana Zama. The conjugate gradient regularization method in Computed Tomography problems. Applied Mathematics and Computation, 102(1):87–99, July 1, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300398100073.

Ping:1999:ANN

[PZZy99] He Ping, Chen Zheng, and Lu Zhi-yong. Artificial neural networks model of share analysis specialist system. Applied Mathematics and Computation, 106(2–3):277–283, December 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300398101297.

Qu:1998:IEB

[QA98] Ruibin Qu and Ravi P. Agarwal. Improved error bounds for freezing solutions of linear boundary value problems. Applied Mathematics and Computation, 94(2–3):97–112, August 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300397100674.

Qiang:1998:CCM

[QC98] J. Qiang and W. H. Choe. A combination of characteristics and Monte Carlo methods for etch profile simulation. Applied Mathematics and Computation, 91(2–3):179–195, May 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300397100170.

Rajendran:1997:SUT

[Raj97]  K. Rajendran.  Simulation

of unsteady turbulent flows with the effect of fluctuating external velocity by using quasi-steady models. Applied Mathematics and Computation, 87(2–3):95–110, December 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300396002214.

Razzaghi:1995:IND

[RAL95] M. Razzaghi, A. Arabshahi, and S. D. Lin. Identification of nonlinear differential equations via Fourier series operational matrix for repeated integration. Applied Mathematics and Computation, 68(2–3):189–198, March 15, 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        009630039400093J.

Razzaghi:1997:ALD

[RAL97] Mohsen Razzaghi, Abdollah Arabshahi, and Shin-Feng D. Lin. Analysis of linear distributed parameter systems via double Fourier series. Applied Mathematics and Computation, 87(2–3):205–215, December 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300396002846.

Ramos:1997:LMRa

[Ram97a]  J. I. Ramos.  Lineariza-

tion methods for reactiondiffusion equations: 1-D problems. Applied Mathematics and Computation, 88 (2–3):199–224, December 30, 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300396003281.

Ramos:1997:LMRb

[Ram97b] J. I. Ramos. Linearization methods for reaction-diffusion equations: Multidimensional problems. Applied Mathematics and Computation, 88 (2–3):225–254, December 30, 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S009630039600327X.

Ramos:1998:ICL

[Ram98a] J. I. Ramos. Implicit, compact, linearized θ-methods with factorization for multidimensional reaction-diffusion equations.

Applied Mathematics and Computation, 94(1):17–43, August 1, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300397101035.

Ramos:1998:MIL

[Ram98b] J. I. Ramos. Maps of implicit, linearized θ-methods for the logistic differential equation. Applied Mathematics and Computation, 94(1):1–15, August 1, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300397101187.

Ramos:1998:PTL

[Ram98c] J. I. Ramos. A piecewise timelinearized method for the logistic differential equation. Applied Mathematics and Computation, 93(2–3):139–148, July 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300397100492.

Ramos:1998:UBC

[Ram98d] J. I. Ramos. Upstream boundary conditions for flows in porous channels. Applied Mathematics and Computation, 93(2–3):149–154, July 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300397100637.

Ramos:1999:AAC

[Ram99a] J. I. Ramos. Asymptotic analysis of compound liquid jets at low Reynolds numbers. Applied Mathematics and Computation, 100(2–3):223–240, May 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300398000228.

Ramos:1999:LFT

[Ram99b] J. I. Ramos. Linearized factorization techniques for multidimensional reaction — diffusion equations. Applied Mathematics and Computation, 100(2–3):201–222, May 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S009630039800023X.

Ramos:1999:LMO

[Ram99c]  J. I. Ramos.  Linearized

methods for ordinary differential equations. Applied Mathematics and Computation, 104(2–3):109–129, September 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300398100565.

Ramos:1999:DME

[Ram99d]  J. I. Ramos.  On diffu-

sive methods and exponentially fitted techniques. Applied Mathematics and Computation, 103(1):69–96, August 1, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300398100401.

Ruxton:1995:JTS

[RB95] David J. W. Ruxton and David J. Bell. Junction times in singular optimal control. Applied Mathematics and Computation, 70(2–3):143–154, July 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        009630039400115K.

Ravindran:1996:MBC

[RBPK96] Varadarajan Ravindran, Badri N. Badriyha, Massoud Pirbazari, and Sung-Hyun Kim. Modeling of bioactive carbon adsorbers: a hybrid weighted residual-finite difference numerical technique. Applied Mathematics and Computation, 76(2–3):99–131, May 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0096300395001514.

Rasulov:1999:ENM

[RC99] Mahir Rasulov and Erhan Coskun. An efficient numerical method for solving the Korteweg–de Vries equation in a class of discontinuous functions. Applied Mathematics and Computation, 102(2–3):139–154, July 15, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S009630039810022X.

Regan:1996:HWL

[Reg96] Marian P. Regan. Half-wave linear rectification of a frequency modulated sinusoid. Applied Mathematics and Computation, 79(2–3):137–162, October 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0096300395002499.

Rivera-Gallego:1998:GAC

[RG98]  Wilson Rivera-Gallego.  A

genetic algorithm for circulant Euclidean distance matrices. Applied Mathematics and Computation, 97 (2–3):197–208, December 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300397101436.

Ramos:1997:NFD

[RGL97a] J. I. Ramos and C. M. García-López. Nonstandard finite difference equations for ODEs and 1-D PDEs based on piecewise linearization. Applied Mathematics and Computation, 86(1):11–36, September 15, 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300396001464.

Ramos:1997:PLM

[RGL97b] J. I. Ramos and C. M. García-López. Piecewiselinearized methods for initialvalue problems. Applied Mathematics and Computation, 82(2–3):273–302, March 15, 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300396000409.

Ramos:1998:IBC

[RGL98] J. I. Ramos and C. M. García-López. Intermediate boundary conditions in operatorsplitting techniques and linearization methods. Applied Mathematics and Computation, 94(2–3):113–136, August 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300397100716.

Rinaldi:1998:LDC

[Rin98]  Sergio Rinaldi.  Love dy-

namics: the case of linear couples. Applied Mathematics and Computation, 95 (2–3):181–192, September 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300397100819.

Reithmeier:1996:RCC

[RL96] E. Reithmeier and G. Leitmann. Robust constrained control for vibration suppression of mismatched systems. Applied Mathematics and Computation, 78(2–3):245–257, September 1, 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0096300396000124.

Rolle:1998:MPP

[Rol98] Jean-Daniel Rolle. A model for perturbed production or measurement processes involving compound normal distributions. Applied Mathematics and Computation, 96(1): 75–88, October 1998. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
sciencedirect.com/science/
article/pii/S0096300397100820.

Ridgley:1997:MDS

[RPT97] Mark A. Ridgley, David C.

Penn, and Liem Tran. Multicriterion decision support for a conflict over stream diversion and land-water reallocation in Hawaii. Applied Mathematics and Computation, 83(2–3):153–172, May 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300396001695.

Riganti:1998:SPT

[RR98] R. Riganti and A. Rossani. Singular perturbation techniques in the study of a dynamical system arising from the kinetic theory of atoms and photons. Applied Mathematics and Computation, 96(1):47–63, October 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300397101242.

Ratschek:1999:ECS

[RR99] Helmut Ratschek and Jon Rokne. Exact computation of the sign of a finite sum. Applied Mathematics and Computation, 99(2–3):99–127, March 15, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300398000101.

Rodellar:1995:ACU

[RRB95] J. Rodellar, E. P. Ryan, and A. H. Barbat. Adaptive control of uncertain coupled mechanical systems with application to base-isolated structures. Applied Mathematics and Computation, 70(2–3):299–314, July 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
009630039400110P.

Ryan:1996:USC

[Rya96] E. P. Ryan. A universal servomechanism for a class of nonlinear planar systems. Applied Mathematics and Computation, 78(2–3):229–243, September 1, 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0096300396000112.

Shawagfeh:1996:NPA

[SA96] N. T. Shawagfeh and G. Adomian. Non-perturbative analytical solution of the general Lotka–Volterra threespecies system. Applied Mathematics and Computation, 76(2–3):251–266, May 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
009630039500162X.

Solomonovich:1997:DEM

[SAF^{+}97] M. Solomonovich, L. P. Apedaile, H. I. Freedman, A. H. Gebremedihen, S. G. M. Schilizzi, and L. Belostotski. A dynamical economic model of sustainable agriculture and the ecosphere. Applied Mathematics and Computation, 84(2–3):221–246, July 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300396000884.

Saied:1996:SSF

[Sai96] Effat A. Saied. On the similarity solutions for the Free Kramer equation. I. Applied Mathematics and Computation, 74(1):59–63, January 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0096300395000887.

Saied:1999:NCS

[Sai99] Effat A. Saied. The nonclassical solution of the inhomogeneous non-linear diffusion equation. Applied Mathematics and Computation, 98(2–3):103–108, February 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300397101588.

Soubeh:1996:CFR

[SAL96] Mohammed Soubeh and Mhd. Jamal Al-Laban. Context-free relations and their characteristics. Applied Mathematics and Computation, 79(2–3): 163–172, October 1996. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
sciencedirect.com/science/
article/pii/0096300395002529.

Sasagawa:1995:CCF

[Sas95a]  T. Sasagawa.  Construc-

tions of C^{∞}-class functions having specified asymptotic properties. Applied Mathematics and Computation, 73(2–3):125–132, December 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0096300394002444.

Sasagawa:1995:IPA

[Sas95b] T. Sasagawa. An infinite product arising from an elementary geometric problem and the estimate of its value. Applied Mathematics and Computation, 73(2–3):271–279, December 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0096300395000674.

Suleiman:1996:VCO

[SBH96] Mohamed Bin Suleiman, Saiman Baok, and George Hall. Varying the componentwise order of the multistep methods in solving ODEs and its absolute stability. Applied Mathematics and Computation, 74(2–3):209–222, February 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0096300395000933.

Su:1996:QED

[SC96] Ching-Tzong Su and Feng Cheng Chang. Quick evaluation of determinants. Applied Mathematics and Computation, 75(2–3):117–118, March 15, 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S009630039690051X.

Scitovski:1997:API

[Sci97]  Rudolf Scitovski.  Analy-

sis of a parameter identification problem. Applied Mathematics and Computation, 82(1):39–55, March 1, 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300396000227.

Sacco:1996:PAL

[SEG96] Riccardo Sacco, Danilo Erricolo, and Emilio Gatti. A perturbation approach for low frequency noise in junction field effect transistors. Applied Mathematics and Computation, 74(2–3):161–190, February 1996. CODEN AMHCBQ.

ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0096300395000917.

Selvakumar:1995:CPS

[Sel95] K. Selvakumar. A computational procedure for solving a chemical flow reactor problem using shooting method. Applied Mathematics and Computation, 68(1):27–40, March 1, 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
009630039400082F.

Sever:1999:CMH

[Sev99a]  Ali Sever.  A computa-

tional method for hyperbolic inverse problem. Applied Mathematics and Computation, 101(1):5–12, June 1, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S009630039800040X.

Sever:1999:UIB

[Sev99b]  Ali Sever.  On unique-

ness for an inverse boundary value problem in electrical prospection. Applied Mathematics and Computation, 106(2–3):103–115, December 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        S0096300398100772.

Shary:1997:CSS

[Sha97] Sergey P. Shary. Controllable solution set to interval static systems. Applied Mathematics and Computation, 86(2–3):185–196, October 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300396001816.

Shen:1996:NSD

[She96] J. H. Shen. The nonoscillatory solutions of delay differential equations with impulses. Applied Mathematics and Computation, 77(2–3):153–165, July 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300395001980.

Suleiman:1996:POD

[SIA96] Mohamed B. Suleiman, Fudziah Bt. Ismail, and Kamel Ariffin B. M. Atan. Partitioning ordinary differential equations using Runge–Kutta methods. Applied Mathematics and Computation, 79(2–3):289–309, October 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0096300395002472.

Simos:1997:MRK

[Sim97] T. E. Simos. Modified Runge– Kutta methods for the numerical solution of ODEs with oscillating solutions. Applied Mathematics and Computation, 84(2–3):131–143, July 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300396000835.

Simos:1999:EEF

[Sim99a] T. E. Simos. Explicit exponentially fitted methods for the numerical solution of the Schrödinger equation. Applied Mathematics and Computation, 98(2–3):185–198, February 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300397101631.

Simos:1999:NFD

[Sim99b]  T. E. Simos.  A new fi-

nite difference scheme with minimal phase-lag for the numerical solution of the Schrödinger equation. Applied Mathematics and Computation, 106(2–3):245–264, December 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300398101273.

Scitovski:1996:MSP

[SJ96] Rudolf Scitovski and Dragan Jukić. A method for solving the parameter identification problem for ordinary differential equations of the second order. Applied Mathematics and Computation, 74(2–3):273–291, February 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0096300395000984.

Smokty:1995:AAS

[SK95] O. I. Smokty and K. Y. Kondratyev. The analytical approximation of the spatial frequency filtration of radiation fields for the space spectrophotometry problems. Applied Mathematics and Computation, 69(1):23–39, April 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
009630039400097N.

Sivaloganathan:1996:NSD

[SK96] S. Sivaloganathan and A. Karageorghis. Numerical solution of the 3-D buoyancy driven cavity problem by a pseudospectral method. Applied Mathematics and Computation, 75(1):43–58, March 1, 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0096300395001026.

Sudarsan:1998:EAN

[SK98a] R. Sudarsan and S. Sathiya Keerthi. An efficient approach for the numerical simulation of multibody systems. Applied Mathematics and Computation, 92(2–3):195–218, June 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300397100418.

Sudarsan:1998:NAS

[SK98b] R. Sudarsan and S. Sathiya Keerthi. Numerical approaches for solution of differential equations on manifolds. Applied Mathematics and Computation, 92(2–3):153–193, June 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300397100650.

Sahin:1995:MRN

[ŞKB95] Ahmet Z. Şahin, Davut Kavranoğlu, and Maamar Bettayeb. Model reduction in numerical heat transfer problems. Applied Mathematics and Computation, 69(2–3):209–225, May 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.

com/science/article/pii/
009630039400128Q.

Sharan:1997:AMM

[SKG97] Maithili Sharan, E. J. Kansa, and Suman Gupta. Application of the multiquadric method for numerical solution of elliptic partial differential equations. Applied Mathematics and Computation, 84(2–3):275–302, July 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300396001099.

Szidarovszky:1996:AOB

[SL96] Ferenc Szidarovszky and Chia-Hung Lin. The alternating offer bargaining method under uncertainty. Applied Mathematics and Computation, 76(2–3):133–141, May 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0096300395001522.

Sun:1999:AQN

[SL99] Fangyu Sun and Xiangfang Li. On an accelerating quasi-Newton circular iteration. Applied Mathematics and Computation, 106(1):17–29, December 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300398100875.

Shen:1995:SDP

[SM95] Zhizhang Shen and Christian M. Marston. A study of a dice problem. Applied Mathematics and Computation, 73(2–3):231–247, December 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0096300395000623.

Szidarovszky:1995:PTC

[SMY95] Ferenc Szidarovszky, Sándor Molnár, and Jerome Yen. On the price trajectory control of a discrete dynamic producerconsumer market. Applied Mathematics and Computation, 73(2–3):249–256, December 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0096300395000631.

Sugisaka:1995:CTC

[SO95] Masanori Sugisaka and Takemasu Okada. Continuous-time Chandrasekhar smoothers in detection and tracking. Applied Mathematics and Computation, 69(1):123–136, April 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
009630039400103B.

Soliman:1998:SCD

[Sol98] A. A. Soliman. Stability criteria of differential systems via Liapunov’s second method. Applied Mathematics and Computation, 92(2–3):143–152, June 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300397100455.

Soliman:1999:SNV

[Sol99a] A. A. Soliman. On _{0}-stability for nonlinear Volterra integrodifferential systems. Applied Mathematics and Computation, 100(1):71–83, April 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300398000150.

Solis:1999:DLS

[Sol99b] Francisco J. Solis. Dimension and local structures of attracting manifolds of smooth dynamical systems. Applied Mathematics and Computation, 100(2–3):169–175, May 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300398000204.

Song:1996:CWR

[Son96]  Yongzhong Song.  Chaotic

waveform relaxation methods for dynamical systems. Applied Mathematics and Computation, 78(1):83–100, August 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0096300395002502.

Soriano:1994:EPC

[Sor94] J. M. Soriano. Extremum points of a convex function. Applied Mathematics and Computation, 66(2–3):261– 266, December 1994. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
sciencedirect.com/science/
article/pii/009630039490121X.

See erratum [Sor97a].

Soriano:1996:EZP

[Sor96] J. M. Soriano. On the existence of zero points. Applied Mathematics and Computation, 79(1):99–104, September 15, 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0096300395002464.

Soriano:1997:EEP

[Sor97a] J. M. Soriano. Erratum: “Extremum points of a convex function”. Applied Mathematics and Computation, 84(1):95, June 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300397002105.  See

[Sor94].

Soriano:1997:NZM

[Sor97b] J. M. Soriano. On the number of zeros of a mapping. Applied Mathematics and Computation, 88(2–3):287–291, December 30, 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300396003360.

Soriano:1999:EZB

[Sor99] J. M. Soriano. Existence of zeros for bounded perturbations of proper mappings. Applied Mathematics and Computation, 99(2–3):255–259, March 15, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300397101837.

Szidarovszky:1995:IMD

[SP95] Ferenc Szidarovszky and Olgierd Palusinski. An inverse method to determine parasitics of power interconnections in high speed electronics. Applied Mathematics and Computation, 73(2–3):203–208, December 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0096300394002509.

Spitaleri:1997:MMG

[SR97] Rosa Maria Spitaleri and Vania Regolo. Multiblock multigrid grid generation algorithms: Overcoming multigrid anisotropy. Applied Mathematics and Computation, 84(2–3):247–267, July 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300396000896.

Srivastava:1995:NEH

[Sri95] Rekha Srivastava. A note on the Epstein–Hubbell generalized elliptic-type integral. Applied Mathematics and Computation, 69(2–3):255–262, May 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
009630039400133O.

Srivastava:1998:FSR

[Sri98] H. M. Srivastava. Further series representations for ζ(2n + 1). Applied Mathematics and Computation, 97(1):1–15, December 1, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S009630039710145X.

Szafranski:1997:OTS

[SS97] Zdzisław Szafrański and Błażej Szmanda. Oscillation theorems for some nonlinear difference equations. Applied Mathematics and Computation, 83(1):43–52, April 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300396000458.

Shen:1998:GCA

[SS98a] Ling Shen and Ferenc Szidarovszky. A globally convergent algorithm for solving special utilization equations. Applied Mathematics and Computation, 90(1):53–60, March 1, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300396003396.

Szidarovszky:1998:SBP

[SS98b] Ferenc Szidarovszky and Ling Shen. A stochastic bargaining process and solution concept in the discrete case. Applied Mathematics and Computation, 92(2–3):219–227, June 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300397100327.

Steiner:1995:CMA

[SST95] W. Steiner, A. Steindl, and H. Troger. Center manifold approach to the control of a tethered satellite system. Applied Mathematics and Computation, 70(2–3):315–327, July 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
009630039400111G.

Salemi:1999:SCM

[SSW99] F. Salemi, V. Salone, and Wang Wendi. Stability of a competition model with two-stage structure. Applied Mathematics and Computation, 99(2–3):221–231, March 15, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300397101801.

Sun:1999:QNT

[SSY99] W. Sun, R. J. B. Sampaio, and J. Yuan. Quasi-Newton trust region algorithm for nonsmooth least squares problems. Applied Mathematics and Computation, 105(2–3): 183–194, November/December 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300398101030.

Steindl:1996:HCT

[ST96] A. Steindl and H. Troger. Heteroclinic cycles in the threedimensional postbifurcation motion of O(2)-symmetric fluid conveying tubes. Applied Mathematics and Computation, 78 (2–3):269–277, September 1, 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0096300396000148.

Stanimirovic:1999:LRG

[Sta99] Predrag S. Stanimirović. Limit representations of generalized inverses and related methods. Applied Mathematics and Computation, 103(1):51–68, August 1, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300398100486.

Sivaloganathan:1998:MPV

[STD98] S. Sivaloganathan, G. Tenti, and J. M. Drake. Mathematical pressure volume models of the cerebrospinal fluid. Applied Mathematics and Computation, 94(2–3):243–266, August 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300397100935.

Sugisaka:1995:ISI

[SU95] Masanori Sugisaka and Sueo Ueno. Introduction to the special issue on adaptive filtering and estimation. Applied Mathematics and Computation, 69(1):1–2, April 1995. CODEN AMHCBQ.

ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
009630039400096M.

Sugisaka:1995:ACF

[Sug95]  Masanori Sugisaka.  Adap-

tive Chandrasekhar filter for linear discrete-time stationary stochastic systems. Applied Mathematics and Computation, 69(1):137–145, April 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
009630039400104C.

Sugisaka:1998:ISI

[Sug98] Masanori Sugisaka. Introduction to the special issue on Artificial Life and Robotics. Applied Mathematics and Computation, 91(1):1–2, April 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300397100091.

Scitovski:1998:ASM

[SUJ98] R. Scitovski, Š. Ungar, and D. Jukić. Approximating surfaces by moving total least squares method. Applied Mathematics and Computation, 93(2–3):219–232, July 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        S0096300397100777.

Sun:1998:NMQ

[Sun98] Wenyu Sun. Newton’s method and quasi-Newton–SQP method for general LC^{1} constrained optimization. Applied Mathematics and Computation, 92 (1):69–84, June 1, 1998. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic). URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/S0096300397100534.

Saha:1996:STM

[SW96] Arindam Saha and Meghanad D. Wagh. Solutions of two minmax recurrences in parallel processing with variable recombination overhead. Applied Mathematics and Computation, 76(2–3):173–211, May 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0096300395001581.

Sugisaka:1998:ICS

[SWL98] Masanori Sugisaka, Xin Wang, and Ju-Jung Lee. Intelligent control strategy for a mobile vehicle. Applied Mathematics and Computation, 91 (1):91–98, April 1998. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic). URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/S0096300397100030.

Song:1998:CRN

[SY98] Jingyan Song and Yeung Yam.

Complex recurrent neural network for computing the inverse and pseudo-inverse of the complex matrix. Applied Mathematics and Computation, 93(2–3):195–205, July 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300397100649.

Satravaha:1997:ALT

[SZ97] Pornchai Satravaha and Songping Zhu. An application of the LTDRM to transient diffusion problems with nonlinear material properties and nonlinear boundary conditions. Applied Mathematics and Computation, 87(2–3):127–160, December 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300396002263.

Szidarovszky:1997:NAO

[Szi97] F. Szidarovszky. A note on the alternating offer method. Applied Mathematics and Computation, 82(2–3):129–135, March 15, 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300396000264.

Szidarovszky:1998:BOD

[Szi98]  Ferenc Szidarovszky.  Bar-

gaining with offer dependent break-down probabilities. Applied Mathematics and Computation, 90(2–3):117–127, March 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300396003542.

Szidarovszky:1999:NCN

[Szi99] Ferenc Szidarovszky. A new Characterization of the nonsymmetric Nash solution. Applied Mathematics and Computation, 106(1):63–68, December 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300398101078.

Tagliani:1998:ITS

[Tag98] Aldo Tagliani. Inverse twosided z transform and moment problem. Applied Mathematics and Computation, 95 (2–3):125–138, September 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300397100571.

Tagliani:1999:HMP

[Tag99] A. Tagliani. Hausdorff moment problem and maximum entropy: a unified approach. Applied Mathematics and Computation, 105(2–3): 291–305, November/December 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S009630039810084X.

Takac:1998:BIT

[Tak98] Peter Takáč. Bifurcations to invariant 2-tori for the complex Ginzburg–Landau equation. Applied Mathematics and Computation, 89(1– 3):241–257, January/February 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S009630039781661X.

Tabata:1999:NIP

[TET99] Minoru Tabata, Nobuoki Eshima, and Ichiro Takagi. The nonlinear integro-partial differential equation describing the logistic growth of human population with migration. Applied Mathematics and Computation, 98(2–3):169–183, February 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300397101722.

Tabata:1998:CPS

[TETH98] Minoru Tabata, Nobuoki Eshima, Ichiro Takagi, and Takashi Hiroyama. The Cauchy problem for the system of equations describing migration motivated by regional economic disparity. Applied Mathematics and Computation, 94(1):45–64, August 1, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300397101163.

Tian:1997:SVC

[TG97] Bo Tian and Yi-Tian Gao. Solutions of a variable-coefficient Kadomtsev–Petviashvili equation via computer algebra. Applied Mathematics and Computation, 84(2–3):125–130, July 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300396001154.

Tibken:1995:SCU

[TH95] B. Tibken and E. P. Hofer. Simulation of controlled uncertain nonlinear systems. Applied Mathematics and Computation, 70(2–3):329–338, July 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        009630039400112H.

Tibken:1996:IAT

[TH96] B. Tibken and E. P. Hofer. Interval analysis as a tool for sensitivity analysis of a hemopoietic model. Applied Mathematics and Computation, 78 (2–3):259–267, September 1, 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0096300396000136.

Tuckness:1995:CAN

[TJ95] Dan G. Tuckness and Bertrand Jost. A critical analysis of the numerical and analytical methods used in the construction of the lunar gravity potential model. Applied Mathematics and Computation, 71(1):69–90, August 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
009630039400141P.

Tang:1997:SMA

[TPGV97] Yi-Fa Tang, Víctor M. Pérez-García, and Luis Vázquez. Symplectic methods for the Ablowitz–Ladik model. Applied Mathematics and Computation, 82(1):17–38, March 1, 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300396000197.

Thandapani:1995:ONB

[TPL95] E. Thandapani, S. Pandian, and B. S. Lalli. Oscillatory and nonoscillatory behavior of second-order functional difference equations. Applied Mathematics and Computation, 70(1):53–66, June 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        009630039400142Q.

Tleimat:1998:NSU

[TPS98] Hazem Tleimat, R. B. Piercey, and B. K. Soni. Numerical simulation of ultrarelativistic, heavy-ion collisions. Applied Mathematics and Computation, 89(1– 3):275–293, January/February 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300397816633.

Tsitouras:1998:EHO

[TS98] Ch. Tsitouras and T. E. Simos. Explicit high order methods for the numerical integration of periodic initial-value problems. Applied Mathematics and Computation, 95(1):15–26, September 1, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300397100868.

Thornburg:1998:SGB

[TSK98] Hugh Thornburg, Bharat K. Soni, and Boyalakuntla Kishore.

A structured grid based solutionadaptive technique for complex separated flows. Applied Mathematics and Computation, 89(1– 3):259–273, January/February 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300397816621.

Tuckness:1995:FMQ

[Tuc95] Dan G. Tuckness. A fast method for quantitatively measuring stability in a threebody dynamical system. Applied Mathematics and Computation, 70(1):29–51, June 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
009630039400136R.

Tuckness:1996:AON

[Tuc96]  Dan G. Tuckness.  Analy-

sis of optical navigation error during Mars entry. Applied Mathematics and Computation, 80(1):1–22, November 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0096300395002758.

Tuckness:1997:PSS

[Tuc97]  Dan G. Tuckness.  Phase

space sensitivities of the Sun– Jupiter triangular libration point. Applied Mathematics and Computation, 88 (2–3):255–265, December 30, 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300396003293.

Toyoda:1999:AHS

[TyG99] Shin’ichi Toyoda and Pegio yukio Gunji. Appearance of hierarchical structure in hyper-dilation model: model of generalized measurement process. Applied Mathematics and Computation, 104(2–3):153–178, September 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300398100632.

Udwadia:1995:ISIa

[Udw95] Firdaus E. Udwadia. Introduction to the special issue on applied mechanics. Applied Mathematics and Computation, 67(1–3):1, January/February 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0096300394000515.

Ueno:1995:PAR

[Uen95] Sueo Ueno. Probabilistic approach for rarefied gas dynamics: Linearized Couette flow. Applied Mathematics and Computation, 69(1):61–73, April 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
009630039400099P.

Udwadia:1996:ISI

[UHL96] F. E. Udwadia, E. Hofer, and G. Leitmann. Introduction to the special issue on dynamics and control. Applied Mathematics and Computation, 78(2–3):101, September 1, 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0096300396900077.

Udwadia:1995:APE

[UK95] Firdaus E. Udwadia and Robert E. Kalaba. An alternate proof for the equation of motion for constrained mechanical systems. Applied Mathematics and Computation, 70(2–3):339–342, July 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
009630039400113I.

Unbehauen:1996:SNT

[Unb96] H. Unbehauen. Some new trends in identification and modeling of nonlinear dynamical systems. Applied Mathematics and Computation, 78 (2–3):279–297, September 1, 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        009630039600015X.

Udwadia:1997:DCN

[UR97] F. E. Udwadia and N. Raju. Dynamics of coupled nonlinear maps and its application to ecological modeling. Applied Mathematics and Computation, 82(2–3):137–179, March 15, 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300396000276.

Udwadia:1995:ISIb

[UTL^{+}95] F. E. Udwadia, H. Troger, G. Leitmann, H. P. Joergl, and H. Flashner. Introduction to the special issue on dynamics and control. Applied Mathematics and Computation, 70(2–3):95, July 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0096300395900357.

Vanmaele:1995:EFE

[Van95] M. Vanmaele. On an external finite element method for a second-order eigenvalue problem on a concave 2Ddomain with Dirichlet boundary conditions. Applied Mathematics and Computation, 71 (2–3):211–226, September 1, 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
009630039400155W.

VanHuylenbroeck:1997:MTT

[Van97] G. Van Huylenbroeck. Multicriteria tools for the tradeoff analysis in rural planning between economic and environmental objectives. Applied Mathematics and Computation, 83(2–3):261–280, May 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300396001889.

VanSiclen:1998:IML

[Van98] Clinton DeW. Van Siclen. Information measure of location precision. Applied Mathematics and Computation, 97 (2–3):287–293, December 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300397101424.

Vazquez:1998:UNA

[Váz98] C. Vázquez. An upwind numerical approach for an American and European option pricing model. Applied Mathematics and Computation, 97 (2–3):273–286, December 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        S0096300397101229.

Varner:1998:NSD

[VC98] T. N. Varner and S. Roy Choudhury. Non-standard difference schemes for singular perturbation problems revisited. Applied Mathematics and Computation, 92(2–3):101–123, June 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S009630039710025X.

VanKeer:1997:FEA

[VD97] R. Van Keer and H. De Schepper. Finite element approximation for 2nd order elliptic eigenvalue problems with nonlocal boundary or transition conditions. Applied Mathematics and Computation, 82(1):1–16, March 1, 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300396000914.

Vestroni:1995:SCT

[VLP95] Fabrizio Vestroni, Angelo Luongo, and Monica Pasca. Stability and control of transversal oscillations of a tethered satellite system. Applied Mathematics and Computation, 70(2–3):343–360, July 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
009630039400114J.

Villatoro:1999:MMEb

[VR99a] F. R. Villatoro and J. I.

Ramos. On the method of modified equations. I: Asymptotic analysis of the Euler forward difference method. Applied Mathematics and Computation, 103(2–3):111–139, August 15, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300398100310.

Villatoro:1999:MMEc

[VR99b] F. R. Villatoro and J. I. Ramos.

On the method of modified equations. II: Numerical techniques based on the equivalent equation for the Euler forward difference method. Applied Mathematics and Computation, 103(2–3):141–177, August 15, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300398100309.

Villatoro:1999:MMEd

[VR99c] F. R. Villatoro and J. I.

Ramos. On the method of modified equations. III. Numerical techniques based on the second equivalent equation for the Euler forward difference method. Applied Mathematics and Computation, 103(2–3):179–212, August 15, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300398100292.

Villatoro:1999:MMEe

[VR99d] F. R. Villatoro and J. I. Ramos. On the method of modified equations. IV. Numerical techniques based on the modified equation for the Euler forward difference method. Applied Mathematics and Computation, 103(2–3):213–240, August 15, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300398100280.

Villatoro:1999:MMEf

[VR99e] F. R. Villatoro and J. I. Ramos. On the method of modified equations. V: Asymptotic analysis of and direct-correction and asymptotic successive-correction techniques for the implicit midpoint method. Applied Mathematics and Computation, 103 (2–3):241–285, August 15, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300398100279.

Villatoro:1999:MMEa

[VR99f] F. R. Villatoro and J. I.

Ramos. On the method of modified equations. VI: Asymptotic analysis of and asymptotic successive-corrections techniques for two-point, boundary-value problems in ODE’s. Applied Mathematics and Computation, 105(2–3):137–171, November/December 1999. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
sciencedirect.com/science/
article/pii/S0096300398100991.

Vatsala:1996:GQF

[VS96] A. S. Vatsala and Donna Stutson. Generalized quasilinearization and first-order periodic boundary value problem. Applied Mathematics and Computation, 77(2–3):113–129, July 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300395001948.

Wong:1996:OAM

[WA96] Patricia J. Y. Wong and Ravi P.

Agarwal. The oscillation and asymptotically monotone solutions of second-order quasilinear differential equations. Applied Mathematics and Computation, 79(2–3):207–237, October 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0096300395002677.

Wong:1998:ABS

[WA98] Patricia J. Y. Wong and Ravi P. Agarwal. Asymptotic behaviour of solutions of higher order difference and partial difference equations with distributed deviating arguments. Applied Mathematics and Computation, 97 (2–3):139–164, December 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300397101254.

Wong:1999:ELB

[WA99] Patricia J. Y. Wong and Ravi P. Agarwal. Eigenvalues of Lidstone boundary value problems. Applied Mathematics and Computation, 104(1):15–31, September 1, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300398100450.

Wang:1995:QSS

[Wan95] Shao-Ming Wang. A quadratic spline structure over triangulations. Applied Mathematics and Computation, 68(2–3):143–152, March 15, 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        009630039400088L.

Wang:1996:NCK

[Wan96] Zeke Wang. Notes on computation of Kakutani fixed points. Applied Mathematics and Computation, 75(2–3):139– 150, March 15, 1996. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
sciencedirect.com/science/
article/pii/S0096300396900557.

Wang:1998:PGM

[Wan98] Yuan-Ming Wang. Petrov– Galerkin methods for systems of nonlinear reactiondiffusion equations. Applied Mathematics and Computation, 96(2–3):209–236, November 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300397101230.

Wang:1999:FOC

[Wan99a] Peiguang Wang. Forced oscillation of a class of delay hyperbolic equation boundary value problem. Applied Mathematics and Computation, 103(1):15–25, August 1, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300398100619.

Wang:1999:OCC

[Wan99b] Peiguang Wang. Oscillatory criteria for a class of delay hyperbolic equations boundary value problem (II). Applied Mathematics and Computation, 100(2–3):189–199, May 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300398000265.

Wazwaz:1995:DMA

[Waz95a] Abdul-Majid Wazwaz. The decomposition method for approximate solution of the Goursat problem. Applied Mathematics and Computation, 69(2–3):299–311, May 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
009630039400137S.

Wazwaz:1995:NAN

[Waz95b]  Abdul-Majid Wazwaz.  A

new approach to the nonlinear advection problem: an application of the decomposition method. Applied Mathematics and Computation, 72(2–3):175–181, October 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0096300394001824.

Wazwaz:1997:NCA

[Waz97a] Abdul-Majid Wazwaz. Necessary conditions for the appearance of noise terms in decomposition solution series. Applied Mathematics and Computation, 81(2–3):265–274, February 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300395003274.

Wazwaz:1997:SBL

[Waz97b]  Abdul-Majid Wazwaz.  A

study on a boundary-layer equation arising in an incompressible fluid. Applied Mathematics and Computation, 87(2–3):199–204, December 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300396002810.

Wazwaz:1997:SCI

[Waz97c] Abdul-Majid Wazwaz. A study on complex integrals involving absolute values. Applied Mathematics and Computation, 85(2–3):281–286, September 1, 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300396001427.

Wazwaz:1998:CBA

[Waz98a]  Abdul-Majid Wazwaz.  A

comparison between Adomian decomposition method and Taylor series method in the series solutions. Applied Mathematics and Computation, 97(1):37–44, December 1, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300397101278.

Wazwaz:1998:RTS

[Waz98b] Abdul-Majid Wazwaz. A reliable technique for solving the wave equation in an infinite one-dimensional medium. Applied Mathematics and Computation, 92(1):1–7, June 1, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300397100376.

Wazwaz:1999:AAP

[Waz99a] Abdul-Majid Wazwaz. Analytical approximations and Padé approximants for Volterra’s population model. Applied Mathematics and Computation, 100(1):13–25, April 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300398000186.

Wazwaz:1999:CSS

[Waz99b] Abdul-Majid Wazwaz. A comparative study on a singular perturbation problem with two singular boundary points. Applied Mathematics and Computation, 99(2–3):179–193, March 15, 1999. CODEN AMHCBQ.

ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S009630039800006X.

Wazwaz:1999:MDM

[Waz99c] Abdul-Majid Wazwaz. The modified decomposition method and Padé approximants for solving the Thomas–Fermi equation. Applied Mathematics and Computation, 105(1): 11–19, October 1999. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
sciencedirect.com/science/
article/pii/S0096300398100905.

Wazwaz:1999:RMA

[Waz99d] Abdul-Majid Wazwaz. A reliable modification of Adomian decomposition method. Applied Mathematics and Computation, 102(1):77–86, July 1, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300398100243.

Wu:1999:UFS

[WC99] Berlin Wu and Mei-Hui Chen.

Use of fuzzy statistical technique in change periods detection of nonlinear time series. Applied Mathematics and Computation, 99(2–3):241– 254, March 15, 1999. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/S0096300398000113.

Wong:1997:IBV

[WCG97] Yau Shu Wong, Qianshun Chang, and Lianger Gong. An initial-boundary value problem of a nonlinear Klein–Gordon equation. Applied Mathematics and Computation, 84(1):77–93, June 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300396000653.

Wei:1998:ISD

[Wei98] Yimin Wei. Index splitting for the Drazin inverse and the singular linear system. Applied Mathematics and Computation, 95(2–3):115–124, September 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300397100984.

Wei:1999:PBS

[Wei99a] Y. Wei. Perturbation bound of singular linear systems. Applied Mathematics and Computation, 105(2–3):211–220, November/December 1999. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic). URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/S0096300399001204.

Wei:1999:PGI

[Wei99b] Yimin Wei. On the perturbation of the group inverse and oblique projection. Applied Mathematics and Computation, 98(1):29–42, January 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300397101515.

White:1997:RRO

[Whi97] Luther W. White. Resolution of regularized output least squares estimation procedures. Applied Mathematics and Computation, 81(2–3):139–172, February 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
009630039500324X.

White:1997:EDE

[WJO97] Luther W. White, Ying-Jun Jin, and Daniel J. O’Meara. Estimation of discontinuous elliptic coefficients. Applied Mathematics and Computation, 81(2–3):113–138, February 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0096300395003231.

Watson:1995:AMC

[WK95] Brian C. Watson and Manohar P.

Kamat. Analysis of mistuned cyclic systems using mistune modes. Applied Mathematics and Computation, 67 (1–3):61–79, January/February 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0096300394000548.

Wazwaz:1996:RTS

[WK96a] A. M. Wazwaz and S. A. Khuri. A reliable technique for solving the weakly singular second-kind Volterra-type integral equations. Applied Mathematics and Computation, 80(2–3):287–299, December 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0096300395002790.

Wazwaz:1996:TMS

[WK96b] A. M. Wazwaz and S. A. Khuri. Two methods for solving integral equations. Applied Mathematics and Computation, 77(1):79–89, June 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0096300395001891.

Wazwaz:1998:NIS

[WK98] A. M. Wazwaz and S. A. Khuri. New ideas for solving sizestructured population models. Applied Mathematics and Computation, 93(1):91–96, July 1, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300397100728.

Wei:1997:LEC

[WL97] J. G. Wei and G. Leng. Lyapunov exponent and chaos of Duffing’s equation perturbed by white noise. Applied Mathematics and Computation, 88(1):77–93, December 15, 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300396003074.

Woodward:1997:STT

[WLB97] J. Arthur Woodward, Wan-Ching Liu, and Douglas G. Bonett. Shortest two-tailed confidence intervals. Applied Mathematics and Computation, 84(1):65–76, June 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S009630039600080X.

Wu:1999:CCI

[WLTS99] Tsu-Chen Wu, Shinq-Hwa Leu, Shih-Tong Tu, and H. M. Srivastava. A certain class of infinite sums associated with Digamma functions. Applied Mathematics and Computation, 105(1):1–9, October 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S009630039810098X.

Wolfe:1996:IMG

[Wol96] M. A. Wolfe. Interval methods for global optimization. Applied Mathematics and Computation, 75(2–3):179–206, March 15, 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0096300396900612.

Wolfe:1998:IEC

[Wol98] M. A. Wolfe. Interval enclosures for a certain class of multiple integrals. Applied Mathematics and Computation, 96(2–3):145–159, November 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300397101199.

Wong:1998:GDD

[Won98] Tzu-Tsung Wong. Generalized Dirichlet distribution in Bayesian analysis. Applied Mathematics and Computation, 97(2–3):165–181, December 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300397101400.

Woodward:1997:ECD

[WP97] J. Arthur Woodward and Christina G. S. Palmer. On the exact convolution of discrete random variables. Applied Mathematics and Computation, 83(1):69–77, April 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300396000471.

Waissi:1996:SAS

[WR96] Gary R. Waissi and Donald F.

Rossin. A sigmoid approximation of the standard normal integral. Applied Mathematics and Computation, 77(1):91–95, June 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0096300395001905.

Wang:1999:COF

[WSL99] Feng-Sheng Wang, Ching-Tzong Su, and Yung-Cheng Liu. Computation of optimal feedforward and feedback control by a modified reduced gradient method. Applied Mathematics and Computation, 104(1):85–100, September 1, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300398100541.

Wu:1997:CSR

[Wu97] Zongmin Wu. Compactly supported radial functions and the Strang–Fix condition. Applied Mathematics and Computation, 84(2–3):115–124, July 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300396001105.

Wang:1998:IDT

[WW98] Alan P. Wang and Cecilia Y. Wang. Identification of a dissipative transport system. Applied Mathematics and Computation, 91(1):33–41, April 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300397100133.

Wen:1999:OPT

[WW99a] Jyh-Horng Wen and Jee-Wey Wang. An occupancy problem with two groups of balls resulting from TDM radio communication application. Applied Mathematics and Computation, 105(1):87–90, October 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300398100863.

Wen:1999:RSO

[WW99b] Jyh-Horng Wen and Jee-Wey Wang. A recursive solution to an occupancy problem resulting from TDM radio communication application. Applied Mathematics and Computation, 101(1):1–3, June 1, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300398000393.

Wu:1995:PSO

[WZ95] J. Wu and X. Zou. Patterns of sustained oscillations in neural networks with delayed interactions. Applied Mathematics and Computation, 73(1):55–75, November 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
009630039400203G.

White:1998:SMF

[WZ98] Luther W. White and Musharraf Zaman. A simple model for fluid accumulation and flow in a porous medium. Applied Mathematics and Computation, 90(2–3):181–203, March 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300397003986.

Wanxie:1998:SAT

[WZZ98] Zhong Wanxie, Xinglai Zhuang, and Jianping Zhu. A selfadaptive time integration algorithm for solving partial differential equations. Applied Mathematics and Computation, 89(1–3):295–312, January/February 1998. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic). URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/S0096300397816645.

Xu:1998:CMM

[XL98] Xuejun Xu and Likang Li. A V -cycle multigrid method and additive multilevel preconditioners for the plate bending problem discretized by conforming finite elements. Applied Mathematics and Computation, 93(2–3):233–258, July 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300397100844.

Xu:1999:NCC

[XL99] Xuejun Xu and Likang Li. A note on convergence of V -cycle nonconforming multigrid methods. Applied Mathematics and Computation, 104(2–3):191–206, September 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300398100796.

Xin:1995:ASI

[XS95] Jingmin Xin and Akira Sano.

Adaptive system identification based on generalized wavelet decomposition. Applied Mathematics and Computation, 69(1):97–109, April 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0096300394001019.

Xiguan:1996:FPS

[XYY96] Lu Xiguan, Li Yong, and Su Yi.

Finding periodic solutions of ordinary differential equations via homotopy method. Applied Mathematics and Computation, 78(1):1–17, August 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
009630039400241X.

Zhu:1998:AAN

[xZ98] Wen xing Zhu. An approximate algorithm for nonlinear integer programming. Applied Mathematics and Computation, 93(2–3):183–193, July 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300397100832.

Xu:1997:GCF

[XZL97] Zong-Ben Xu, Jiang-She Zhang, and Yiu-Wing Leung. A general CDC formulation for specializing the cell exclusion algorithms of finding all zeros of vector functions. Applied Mathematics and Computation, 86(2–3):235–259, October 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300396001919.

Xu:1998:AAC

[XZL98] Zong-Ben Xu, Jiang-She Zhang, and Yiu-Wing Leung. An approximate algorithm for computing multidimensional convex hulls. Applied Mathematics and Computation, 94 (2–3):193–226, August 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300397100431.

Xu:1996:CEA

[XZW96] Zong-Ben Xu, Jiang-She Zhang, and Wei Wang. A cell exclusion algorithm for determining all the solutions of a nonlinear system of equations. Applied Mathematics and Computation, 80(2–3):181–208, December 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0096300395002863.

Yamamoto:1995:CSB

[Yam95] Tokufumi Yamamoto. Critical state between local and global interaction. Applied Mathematics and Computation, 73(2–3):153–180, December 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0096300394002479.

Yan:1997:OSO

[Yan97] Jurang Yan. Oscillations of second order neutral functional differential equations. Applied Mathematics and Computation, 83(1):27–41, April 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300396000434.

Yang:1998:CSS

[Yan98]  Suh-Yuh Yang.  On the

convergence and stability of the standard least squares finite element method for firstorder elliptic systems. Applied Mathematics and Computation, 93(1):51–62, July 1, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300397100509.

Yang:1999:CSL

[Yan99]  Z. J. Yang.  Cylindrical

symmetry-like solutions of Laplace equation ∇^{2}V (x_{j}) = 0. Applied Mathematics and Computation, 99(1):29–34, March 1, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S009630039710162X.

Yen:1999:NAI

[YB99] Jerome Yen and Tung X. Bui. The negotiable alternatives identifier for group negotiation support. Applied Mathematics and Computation, 104(2–3):259–276, September 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300398100693.

Yildiz:1997:MCF

[YBR97] Bünyamin Yildiz, Mustafa Bayram, and Mahir A. Rasulov. On a method of calculation of the first phase saturation during the process of displacement of oil by water from porous media. Applied Mathematics and Computation, 85(1):1–16, August 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300396000823.

Yevi:1998:PGP

[YCZ98] G. Yevi, P. Cinnella, and X. Zhuang. On parallelizing a groundwater pollution simulator. Applied Mathematics and Computation, 89(1– 3):313–325, January/February 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300397816657.

Ye:1998:DDLb

[Ye98a] X. Ye. Domain decomposition for a least-square finite element method for second order elliptic problem. Applied Mathematics and Computation, 91(2–3):233–242, May 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300397100145.

Ye:1998:DDLa

[Ye98b]  X. Ye.  Domain decom-

position for least-squares finite element methods for the Stokes equations. Applied Mathematics and Computation, 97(1):45–53, December 1, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S009630039710159X.

Ye:1999:TGD

[Ye99] X. Ye. Two grid discretizations with backtracking of the stream function form of the Navier–Stokes equations. Applied Mathematics and Computation, 100(2–3):131–138, May 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300398000241.

Yeh:1997:TRM

[Yeh97] Tyan Yeh. Trapezoidal rule for multiple integrals over hyperquadrilaterals. Applied Mathematics and Computation, 87(2–3):227–246, December 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S009630039600286X.

Yakowitz:1997:MOD

[YH97] Diana S. Yakowitz and Keith W. Hipel. Multiple objective decision making for “lokahi” (balance) in environmental management. Applied Mathematics and Computation, 83(2–3):97–115, May 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300396001713.

Yong:1995:PSF

[YHXX95] Li Yong, Wang Huaizhong, Lu Xianrui, and Lu Xiguan. Periodic solutions for functional differential equations with infinite lead and delay. Applied Mathematics and Computation, 70(1):1–28, June 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.

com/science/article/pii/
009630039400131M.

Hou:1999:CPF

[yHyG99] Jing yu Hou and Ben yu Guo.

Chebyshev pseudospectral-finite element method for the threedimensional unsteady Navier– Stokes equation. Applied Mathematics and Computation, 101(2–3):209–244, June 15, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300398000411.

Yildiz:1996:MIP

[Yil96] Bünyamin Yildiz. A method increasing preciseness of finite differences system and its application. Applied Mathematics and Computation, 75(2–3):253–267, March 15, 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0096300396900697.

Yuan:1999:CGA

[YJ99] Jin-Yun Yuan and Xiao-Qing Jin. Convergence of the generalized AOR method. Applied Mathematics and Computation, 99(1):35–46, March 1, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300397101758.

Yildiz:1997:TMS

[YK97a] Bünyamin Yildiz and Ahmet Kacar. Two methods for solution of a linear boundary value problem of order 2 n. Applied Mathematics and Computation, 86(2–3):215–233, October 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S009630039600183X.

Yon:1997:TLA

[YK97b] Yoon J. Yon and Do Y. Kwak. Two-level additive Schwarz preconditioners for P 1 nonconforming finite elements for nonsymmetric and indefinite problems. Applied Mathematics and Computation, 87(1):1–14, November 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300396001920.

Yu:1996:HMC

[YL96] Bo Yu and Zhenghua Lin. Homotopy method for a class of nonconvex Brouwer fixedpoint problems. Applied Mathematics and Computation, 74(1):65–77, January 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0096300395000895.

Yang:1998:UAW

[YL98a] Suh-Yuh Yang and Jinn-Liang Liu. A unified analysis of a weighted least squares method for first-order systems. Applied Mathematics and Computation, 92(1):9–27, June 1, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300397100467.

Yao:1998:TDA

[YL98b] Xin Yao and Yong Liu. Towards designing artificial neural networks by evolution. Applied Mathematics and Computation, 91(1):83–90, April 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300397100054.

Yavneh:1998:MSS

[YO98] Irad Yavneh and Elena Olvovsky.

Multigrid smoothing for symmetric nine-point stencils. Applied Mathematics and Computation, 92(2–3):229–246, June 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300397100352.

Yi:1998:SPC

[YS98a] Fahuai Yi and T. M. Shih.

Stefan problem with convection. Applied Mathematics and Computation, 95(2– 3):139–154, September 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300397100236.

Yildiz:1998:ICP

[YS98b] Bünyamin Yildiz and Ali Sever. On the inverse conductivity problem. Applied Mathematics and Computation, 94(1):91–96, August 1, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300397100339.

Yang:1997:ASH

[YST97] Xiaofeng Yang, Tielong Shen, and Katsutoshi Tamura. Approximate solution of Hamilton– Jacobi inequality by neural networks. Applied Mathematics and Computation, 84(1):49–64, June 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300396000537.

Yurekli:1999:NII

[Yür99] Osman Yürekli. New identities involving the Laplace and the L_{2}-transforms and their applications. Applied Mathematics and Computation, 99(2–3):141–151, March 15, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300398000022.

Zaghrout:1996:AMS

[ZA96] A. A. S. Zaghrout and S. H. Attalah. Analysis of a model of stage-structured population dynamics growth with time statedependent time delay. Applied Mathematics and Computation, 77(2–3):185–194, July 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S009630039500212X.

Zaghrout:1996:OGA

[ZAES96] A. Zaghrout, A. Ammar, and M. M. A. El-Sheikh. Oscillations and global attractivity in delay differential equations of population dynamics. Applied Mathematics and Computation, 77(2–3):195–204, July 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300395002138.

Zhang:1997:CSA

[ZB97] Jun Zhang and John A. Belward. Chebyshev series approximations for the Bessel function Y_{n}(z) of complex argument. Applied Mathematics and Computation, 88 (2–3):275–286, December 30, 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300396003359.

Zeng:1999:HAF

[Zen99] Pan Zeng. High-accuracy formula for discrete calculation of Fourier transforms. Applied Mathematics and Computation, 106(2–3):117–140, December 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300398100802.

Zhang:1995:SCT

[Zha95] B. G. Zhang. Sturm comparison theorem of difference equations. Applied Mathematics and Computation, 72(2–3):277–284, October 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        009630039400201E.

Zhang:1996:NWW

[Zha96a]  Hong Zhang.  A note on

windowing for the waveform relaxation method. Applied Mathematics and Computation, 76(1):49–63, April 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.

com/science/article/pii/
0096300395001492.

Zhang:1996:AFP

[Zha96b] Jun Zhang. Acceleration of fivepoint red-black Gauss–Seidel in multigrid for Poisson equation. Applied Mathematics and Computation, 80(1):73–93, November 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0096300395002766.

Zhang:1997:MRS

[Zha97a] Jun Zhang. Minimal residual smoothing in multi-level iterative method. Applied Mathematics and Computation, 84 (1):1–25, June 1997. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
sciencedirect.com/science/
article/pii/S0096300396000501.

Zhang:1997:RST

[Zha97b] Jun Zhang. Residual scaling techniques in multigrid, I: Equivalence proof. Applied Mathematics and Computation, 86(2–3):283–303, October 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300396001944.

Zhao:1997:EFF

[Zha97c] Yun-Bin Zhao. Exceptional families and finite-dimensional variational inequalities over polyhedral convex sets. Applied Mathematics and Computation, 87(2–3):111–126, December 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S009630039600224X.

Zhang:1998:RST

[Zha98a] Jun Zhang. Residual scaling techniques in multigrid, II: Practical applications. Applied Mathematics and Computation, 90(2–3):229–252, March 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300397004001.

Zhang:1998:TGA

[Zha98b] Jun Zhang. Two-grid analysis of minimal residual smoothing as a multigrid acceleration technique. Applied Mathematics and Computation, 96(1):27–45, October 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300397101059.

Zhang:1998:VCV

[Zha98c] Jun Zhang. VML: a class of virtual multi-level iterative methods for solving partial differential equations. Applied Mathematics and Computation, 92(1):29–48, June 1, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300397100388.

Zhao:1998:GCF

[Zha98d] Jinxi Zhao. The generalized Cholesky factorization method for saddle point problems. Applied Mathematics and Computation, 92(1):49–58, June 1, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300397100406.

Zhang:1999:ASP

[Zha99a] Jun Zhang. Acceleration and stabilization properties of minimal residual smoothing technique in multigrid. Applied Mathematics and Computation, 100(2–3):151–168, May 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300398000290.

Zhao:1999:OSC

[Zha99b] Yun-Bin Zhao. D-orientation sequences for continuous functions and nonlinear complementarity problems. Applied Mathematics and Computation, 106(2–3):221–235, December 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S009630039810125X.

Han:1999:FFF

[zHhT99] Zheng zhong Han and Yu hua Tang. Fuzzy forecast of flood disaster caused by solar proton flares. Applied Mathematics and Computation, 98(1):83–89, January 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300397101527.

Zaghrout:1998:CBT

[ZK98] A. A. S. Zaghrout and F. M. Kandil. Competition between three microbial populations for a single limiting resource in continuous culture. Applied Mathematics and Computation, 92(2–3):271–281, June 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300397100273.

Zhang:1998:ASP

[ZL98a] Wenfei Zhang and Xiaojiang Li. Analytic solution of poro-mechanical problems in a hollow axisymmetric domain. Applied Mathematics and Computation, 97 (2–3):209–221, December 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300397100583.

Zhu:1998:AMB

[ZL98b] Song-Ping Zhu and Huan-Wen Liu. On the application of multiquadric bases in conjunction with the LT-DRM method to solve nonlinear diffusion equations. Applied Mathematics and Computation, 96(2–3):161–175, November 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300397101308.

Zhu:1998:FWA

[ZL98c] Tonglin Zhu and Wei Lin.

Fast wavelet algorithm of the Poisson integral. Applied Mathematics and Computation, 96(2–3):127–144, November 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300397101114.

Zorzano:1999:NST

[ZMV99] M. P. Zorzano, H. Mais, and L. Vazquez. Numerical solution of two dimensional Fokker– Planck equations. Applied Mathematics and Computation, 98(2–3):109–117, February 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        S0096300397101618.

Zheng:1999:MMU

[ZPY99] Chen Zheng, He Ping, and Chen Yan. A mathematical model using artificial neural networks to forecast shares tendency. Applied Mathematics and Computation, 99(1):71–76, March 1, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300397101734.

Zaslavsky:1998:AMT

[ZS98] Leonid Y. Zaslavsky and Tamar Schlick. An adaptive multigrid technique for evaluating long-range forces in biomolecular simulations. Applied Mathematics and Computation, 97 (2–3):237–250, December 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300397101461.

Zheng:1999:SSI

[ZS99] Shiming Zheng and Fangyu Sun. Some simultaneous iterations for finding all zeros of a polynomial with high order convergence. Applied Mathematics and Computation, 99(2–3):233–240, March 15, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.

com/science/article/pii/
S0096300398000095.

Zhang:1998:LAT

[ZSL98] Y. G. Zhang, M. Sugisaka, and X. J. Li. Lifelike artificial trees based on Growth Iterated Function System. Applied Mathematics and Computation, 91(1):3–8, April 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300397100108.

Zhou:1995:NAO

[ZW95a] Jing Zhou and Luther W.

White. Numerical approximation of optimal control of fluid equations with buoyancy. Applied Mathematics and Computation, 71(1):47–68, August 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
009630039400139U.

Zhou:1995:OCF

[ZW95b] Jing Zhou and Luther W.

White. Optimal control of fluid equations with buoyancy. Applied Mathematics and Computation, 71(1):15–45, August 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
009630039400138T.

Zhenghua:1996:CHI

[ZYB96] Lin Zhenghua, Li Yong, and Yu Bo. A combined homotopy interior point method for general nonlinear programming problems. Applied Mathematics and Computation, 80(2–3):209–224, December 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0096300395002952.

Zuhe:1997:IME

[ZYW97] Shen Zuhe, Huane Zhen Yu, and M. A. Wolfe. An interval maximum entropy method for a discrete minimax problem. Applied Mathematics and Computation, 87(1):49–68, November 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300396002202.

Zeng:1998:SAS

[ZZ98] Jinping Zeng and Shuzi Zhou. Schwarz algorithm for the solution of variational inequalities with nonlinear source terms. Applied Mathematics and Computation, 97(1):23– 35, December 1, 1998. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic). URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/S0096300397101291.